### v4

# The syzygy distinguisher

Abstract
We present a new distinguisher for alternant and Goppa codes, whose complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover it does not suffer from the strong regime limitations of the previous distinguishers or structure recovery algorithms: in particular, it applies to the codes used in the Classic McEliece candidate for postquantum cryptography standardization. The invariants that allow us to distinguish are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code.
Since its introduction in 1978, this is the first time an analysis of the McEliece cryptosystem breaks the exponential barrier.
1 Introduction
In the McEliece cryptosystem [20], a private message is encoded as a codeword in a public binary Goppa code [13], with some noise added. Knowing the secret algebraic data that served to construct the public code, the legitimate recipient has an efficient decoding algorithm and can recover the message. However, to an attacker, the public code looks like a random code, and removing the noise is untractable.
Slightly more formally, a security proof for the system relies on two assumptions:
  1. 1.
    Goppa codes are computationally indistinguishable from generic linear codes (say, when described by generator matrices in reduced row echelon form).
  2. 2.
    Decoding a generic linear code is difficult.
Cryptanalytic attempts can be classified depending on whether they target assumption 1 or 2.
First, those aiming at assumption 1 themselves come in two flavours:
  • Distinguishers address the decisional version of the problem: given a generator matrix, decide whether it is that of a Goppa code or a generic code.
  • Key recovery attacks address the computational version: recover the Goppa structure of the code, or at least an equivalent one [12], if it exists.
Although certain arguments, such as the fact that the class of Goppa codes is very large, make assumption 1 plausible, it remains quite ad hoc from a theoretical point of view. Its sole virtue is that it passed the test of time. There seems to be something special with Goppa codes happening there: indeed, variants of the McEliece system were proposed, with Goppa codes replaced with other types of codes allowing more manageable parameters; however, most of these propositions were eventually broken, as the hidden structure of the codes could be recovered. Also, although the McEliece cryptosystem marked the birth of code-based cryptography, the idea of having an object (such as a Goppa code) constructed from data defined over an extension field and then masked by considering it over a small subfield, was then found in other branches of cryptography: for instance it is at the basis of the HFE cryptosystem [23] in multivariate cryptography. Such systems can often be attacked by algebraic methods (including, but not limited to, the use of Gröbner basis algorithms). This suggests that if a weakness in assumption 1 were to be uncovered one day, then algebraic methods should be a tool of choice. However, up to now, the best distinguishers and key recovery algorithms such as those in [11][7][6][1] only apply to alternant or Goppa codes with very degraded parameters. Against McEliece with cryptographically relevant parameters, either they have exponential complexity with large constants, which makes them useless, or worse: they simply cease to work.
Assumption 2, on the other hand, stands on a firm theoretical ground: the decoding problem for generic linear codes is known to be NP-hard [2]. As such, it is believed to resist the advent of quantum computers, which made the McEliece system a good candidate for postquantum cryptography. However, starting with Prange’s information set decoding algorithm [24], generic decoding methods saw continuous incremental improvements over time. Joint with technological progress in computational power, this eventually led to practical message recovery attacks against the McEliece system with its initially proposed parameter set. However, it appears that this weakness was only the result of a too optimistic choice of parameters. With the need for new standards for postquantum cryptography, an updated version named Classic McEliece was proposed, still relying on binary Goppa codes, but with more conservative parameter sets adapted to resist the best generic decoding attacks with some margin of safety. In the design rationale of this new system [4] one can find an impressive list of several dozens of papers on generic decoding algorithms, ranging over the last five decades. It is then observed that all these algorithms have complexity exponential in the error-correcting capability of the code. Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order! This gives a quite convincing feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover, for given finite parameters, it is more efficient than previous distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
Principles and organization
A natural strategy to build a distinguisher is to design code invariants — quantities that intrinsically depend only on the code, not on the choice of a generator matrix — that behave differently for the classes of codes we want to distinguish. The invariants we use here are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code. Generators of the dual of an alternant or Goppa code, after extension of scalars, satisfy quadratic relations of a special form: they can be expressed as $2\times 2$ minors of a matrix. As such, we can find relations between these quadratic relations, called syzygies. Then these syzygies also satisfy relations, and iterating this process we get higher syzygies up to some order that we can estimate. However, for generic codes, we do not expect this to happen in the same magnitude. This directly gives a distinguisher, at least in theory. In practice, computing these spaces of syzygies only involves basic linear algebra: they can be constructed iteratively, as the kernel of some generalized Macaulay matrices. This can be done efficiently, except for the fact that the dimension of the spaces involved grows exponentially. Our last ingredient is then shortening, which allows us to work with syzygies of a lesser order, and keep these dimensions more under control.
Let us quickly illustrate our result with two basic examples.
First, our distinguisher can be seen as a generalization of the so-called square code distinguisher, first presented in [11], reinterpreted in [19], and fully analyzed in [21]. Let ${\mathsf{C}}$ be a $[n,k]$-code, and $S_{2}$ the space of quadratic forms in $k$ indeterminates. Let
$\operatorname{ev}_{2}:S_{2}\longrightarrow\mathbb{F}^{n}$ (1)
be the evaluation map at the columns of a given generator matrix of ${\mathsf{C}}$. Then the image of $\operatorname{ev}_{2}$ is the square code ${\mathsf{C}}^{\langle 2\rangle}$, and its kernel is the space of quadratic relations $I_{2}({\mathsf{C}})$. The dimensions of these spaces are related, and can be expressed as a Betti number:
$\beta_{1,2}({\mathsf{C}})=\dim(I_{2}({\mathsf{C}}))=\binom{k+1}{2}-\dim({\mathsf{C}}^{\langle 2\rangle}).$ (2)
Now [11] gives a lower bound on this $\beta_{1,2}({\mathsf{C}})$ when ${\mathsf{C}}$ is the dual of an alternant or (binary) Goppa code. On the other hand, [3] shows $\beta_{1,2}({\mathsf{C}})=\left(\binom{k+1}{2}-n\right)^{+}$ with high probability when ${\mathsf{C}}$ is random. If this quantity is smaller than the said lower bound, we can distinguish.
Likewise we claim that Theorem 2.8 of [9] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes, where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher. For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$. Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$. By definition, $Q_{1},Q_{2},Q_{3}$ do not satisfy linear relations with coefficients in $\mathbb{F}$, however they can satisfy relations whose coefficients are forms of degree $1$. Such relations are also called degree $3$ syzygies. They live in the kernel of the degree $3$ Macaulay matrix
$\mathbf{M}_{3}:I_{2}({\mathsf{C}})\otimes S_{1}\longrightarrow S_{3}$ (3)
where $S_{1}$ (resp. $S_{3}$) is the space of homogeneous linear (resp. cubic) forms in $k=4$ variables. Now Theorem 2.8 of [9] shows that for a non-$\operatorname{{\mathsf{GRS}}}$ code the map $\mathbf{M}_{3}$ is injective. On the other hand, if ${\mathsf{C}}$ is $\operatorname{{\mathsf{GRS}}}$ with standard basis $\mathbf{y},\mathbf{y}\mathbf{x},\mathbf{y}\mathbf{x}^{2},\mathbf{y}\mathbf{x}^{3}$ then we can take
$Q_{1}=X_{1}X_{3}-X_{2}^{2},\quad Q_{2}=X_{1}X_{4}-X_{2}X_{3},\quad Q_{3}=X_{2}X_{4}-X_{3}^{2}$ (4)
and these satisfy the syzygies
[…]
Thus we can distinguish by computing $\beta_{2,3}({\mathsf{C}})=\dim\ker(\mathbf{M}_{3})$, which will yield $2$ for $\operatorname{{\mathsf{GRS}}}$ and 00 for non-$\operatorname{{\mathsf{GRS}}}$ $[7,4]$-codes.
We generalize these examples to higher Betti numbers, following the exact same pattern:
  • On one hand, we give lower bounds on the Betti numbers of algebraic codes (dual alternant, dual Goppa, and their shortened subcodes). This is done in section 3, using the Eagon-Northcott complex, a tool precisely crafted to detect long conjugate $\operatorname{{\mathsf{GRS}}}$ subcodes.
  • On the other hand, we estimate the Betti numbers of random codes in terms of raw code parameters (length, dimension, distance). This is done in section 4, partially relying on heuristics for which we do not have full proofs, but still providing experimental evidence, and sometimes also partial theoretical arguments.
Prior to that, section 2 explains how these invariants can be effectively computed. Last, section 5 combines everything and chooses parameters in order to optimize asymptotic complexity. Of special importance in this regard is Proposition 4, which shows that the property that the ideal of a code contains minors of a matrix of linear forms passes to its shortened subcodes.
Related (and unrelated) works
As already noted, our distinguisher can be seen as a generalization of the square distinguisher of [11]. Using an approach similar to ours, the work [6] also extends this square distinguisher by exploiting special properties of the space of quadratic relations, but in a different direction. Last, the key recovery attack in [1] combines shortening of the dual code with ideas from the square distinguisher, and a careful algebraic modeling in order to apply tools from Gröbner basis theory. All these results have limited range of applicability, but they introduced numerous techniques that influenced the present work.
From a geometric point of view, the Betti numbers and the syzygies we consider are those of a set of points in projective space (namely, defined by the columns of a parity check matrix of the code). As such, they have been already extensively studied. Of notable importance to us are the works [14][18][16][10], in relation with the so-called minimal resolution conjecture — regardless of it being false in general: we just request it being “true enough”. Initially, syzygies of sets of points were considered as a mere tool in the study of syzygies of curves. They were then studied for themselves, but the focus was mostly on points in generic position, over an algebraically closed field. Keeping applications to coding theory and cryptography in mind, we will have more interest in finite field effects.
Syzygies sometimes appear as a tool in cryptanalytic works, or in the study of Gröbner basis algorithms; however in general only the first module of syzygies is considered, not those of higher order. Likewise, a few works in coding theory (such as [26] or [15]) use homological properties of finite sets of points; but the applications differ from ours.
Last, note that our approach is apparently unrelated to the series of works initiated with [17]: while these authors also define Betti numbers for codes, these are constructed from the Stanley-Reisner ring of the code matroid, not the homogeneous coordinate ring. This leads to different theories, although seeking links between the two could be an interesting project.
Acknowledgements
The author learned a lot from the code-based cryptography seminar organized by Jean-Pierre Tillich at Inria Paris. More important perhaps, without his friendly pressure, probably this work would have stayed eternally in limbo. He should be thanked for all that. The very favourable environment towards research at ANSSI in general, and at LCR in particular, was also highly helpful. Henri Gilbert’s advices led to improvements in several aspects of the presentation. Anonymous members of LCR unsuccessfully suggested the title “The Dyztynguysher” for this work.
The author is part of the ANR-21-CE39-0009 BARRACUDA and ANR-21-CE39-0008 KLEPTOMANIAC research projects. Some of the computations were done on the lame servers in Télécom Paris.
Notation and conventions
We use row vector convention. We try to consistently use lowercase bold font for codewords and vectors: $\mathbf{c}$, $\mathbf{x}$, $\mathbf{y}$; uppercase bold for matrices: $\mathbf{G}$, $\mathbf{H}$, $\mathbf{M}$; sans-serif for codes: ${\mathsf{C}}$, $\operatorname{{\mathsf{GRS}}}$, $\operatorname{{\mathsf{Alt}}}$, $\operatorname{{\mathsf{Gop}}}$.
The book [9] will be our main source on syzygies. For codes, especially the link between powers of codes and the geometric view on coding theory, we will follow [25].
Given a field $\mathbb{F}$, we see $\mathbb{F}^{n}$ as the standard product algebra of dimension $n$. Thus $\mathbb{F}^{n}$ is not a mere vector space, it comes canonically equipped with componentwise multiplication: for $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$ in $\mathbb{F}^{n}$,
$\mathbf{x}\mathbf{y}=(x_{1}y_{1},\dots,x_{n}y_{n}).$ (6)
(Some authors call this the Schur product of $\mathbf{x}$ and $\mathbf{y}$; how the name of this great mathematician ended associated with this trivial operation is quite convoluted.)
In any algebra, we can define a trace bilinear form. In the case of $\mathbb{F}^{n}$, this trace bilinear form is the standard scalar product:
$\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}+\cdots+x_{n}y_{n}.$ (7)
A $k$-dimensional subspace ${\mathsf{C}}\subseteq\mathbb{F}^{n}$ is called a $[n,k]$-code (and a $[n,k]_{q}$-code in case $\mathbb{F}=\mathbb{F}_{q}$). The orthogonal space ${\mathsf{C}}^{\perp}$ is called the dual code of ${\mathsf{C}}$.
Componentwise multiplication extends to codes, taking the linear span: for ${\mathsf{C}},{\mathsf{C}}^{\prime}\subseteq\mathbb{F}^{n}$,
${\mathsf{C}}{\mathsf{C}}^{\prime}=\langle\mathbf{c}\mathbf{c^{\prime}}:\>\mathbf{c}\in{\mathsf{C}},\,\mathbf{c^{\prime}}\in{\mathsf{C}}^{\prime}\rangle_{\mathbb{F}}.$ (8)
Powers ${\mathsf{C}}^{\langle r\rangle}$ of a code are defined inductively: ${\mathsf{C}}^{\langle 0\rangle}=\mathbb{F}\cdot\mathbf{1}$ is the $1$-dimensional repetition code, and ${\mathsf{C}}^{\langle r+1\rangle}={\mathsf{C}}^{\langle r\rangle}{\mathsf{C}}$.
If ${\mathsf{C}}$ is a $[n,k]$-code, a generator matrix for ${\mathsf{C}}$ is a $k\times n$ matrix $\mathbf{G}$ whose rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ form a basis of ${\mathsf{C}}$. A parity check matrix $\mathbf{H}$ for ${\mathsf{C}}$ is a generator matrix for ${\mathsf{C}}^{\perp}$.
Thanks to the algebra structure, polynomials in one or several variables can be evaluated in $\mathbb{F}^{n}$. In particular, let
$S=\mathbb{F}[X_{1},\dots,X_{k}]$ (9)
be the algebra of polynomials in $k$ variables over $\mathbb{F}$, graded by total degree. Evaluation at the rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of $\mathbf{G}$ then gives linear map
$\operatorname{ev}_{\mathbf{G}}:S\longrightarrow\mathbb{F}^{n}.$ (10)
Observe that if $\mathbf{p}_{1},\dots,\mathbf{p}_{n}$ are the columns of $\mathbf{G}$, then for $f(X_{1},\dots,X_{k})\in S$ we have
$\begin{split}\operatorname{ev}_{\mathbf{G}}(f)&=f(\mathbf{c}_{1},\dots,\mathbf{c}_{k})\\ &=(f(\mathbf{p}_{1}),\dots,f(\mathbf{p}_{n}))\end{split}$ (11)
where in the first line we have one evaluation of $f$ at a $k$-tuple of vectors, while in the second line we have a vector of evaluations of $f$ at $k$-tuples of scalars.
A code is projective if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})\geq 3$, or equivalently if no two of the $\mathbf{p}_{i}$ are proportional. Any code can be “projectivized” by discarding (puncturing) coordinates, keeping only one $\mathbf{p}_{i}$ in each nonzero proportionality class.
Restricting to homogeneous polynomials of degree $r$, we have a surjective map $S_{r}\longrightarrow{\mathsf{C}}^{\langle r\rangle}$, whose kernel we denote $I_{r}({\mathsf{C}})$. We then define the homogeneous coordinate ring of ${\mathsf{C}}$:
${\mathsf{C}}^{\langle\cdot\rangle}=\bigoplus_{r\geq 0}{\mathsf{C}}^{\langle r\rangle}$ (12)
and its homogeneous ideal:
$I({\mathsf{C}})=\bigoplus_{r\geq 0}I_{r}({\mathsf{C}}).$ (13)
It turns out these are also the homogeneous coordinate ring and the homogeneous ideal of the finite set of points
$\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}\subseteq\mathbf{P}^{k-1}.$ (14)
The short exact sequence
$0\longrightarrow I({\mathsf{C}})\longrightarrow S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}\longrightarrow 0$ (15)
makes ${\mathsf{C}}^{\langle\cdot\rangle}$ a homogeneous quotient ring of $S$. In this work we will use coordinates, but identifying $S$ with the symmetric algebra of ${\mathsf{C}}$ would allow to make all this coordinatefree.
Given $\mathbf{x},\mathbf{y}\in\mathbb{F}^{n}$, all entries of $\mathbf{x}$ distinct, all entries of $\mathbf{y}$ nonzero, the generalized Reed-Solomon code of order $k$ with support vector $\mathbf{x}$ and multiplier $\mathbf{y}$ is
$\operatorname{{\mathsf{GRS}}}_{k}(\mathbf{x},\mathbf{y})=\langle\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}\rangle_{\mathbb{F}}=\{\mathbf{y}f(\mathbf{x}):\>f(X)\in\mathbb{F}[X]_{<k}\}\subseteq\mathbb{F}^{n}.$ (16)
It is a $[n,k]$-code if $k\leq n$.
Now let $\mathbb{F}_{q}\subseteq\mathbb{F}_{q^{m}}$ be an extension of finite fields. Given $\mathbf{x},\mathbf{y}\in(\mathbb{F}_{q^{m}})^{n}$ satisfying the same conditions as above, the alternant code of order $t$ and extension degree $m$ over $\mathbb{F}_{q}$, with support $\mathbf{x}$ and multiplier $\mathbf{y}$, is
$\begin{split}\operatorname{{\mathsf{Alt}}}_{t}(\mathbf{x},\mathbf{y})&=\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x},\mathbf{y})^{\perp}\cap(\mathbb{F}_{q})^{n}\\ &=\{\mathbf{c}\in(\mathbb{F}_{q})^{n}:\>c_{1}y_{1}x_{1}^{j}+\cdots+c_{n}y_{n}x_{n}^{j}=0\quad(0\leq j<t)\},\end{split}$ (17)
with parameters $[n,(\geq)n-mt]_{q}$.
Last, given a polynomial $g(X)\in\mathbb{F}_{q^{m}}[X]$ that does not vanish on any entry of $\mathbf{x}$, the $q$-ary Goppa code with support $\mathbf{x}$ and Goppa polynomial $g$ is
$\operatorname{{\mathsf{Gop}}}(\mathbf{x},g)=\operatorname{{\mathsf{Alt}}}_{\deg(g)}(\mathbf{x},g(\mathbf{x})^{-1}).$ (18)
We will work mostly in the class
$\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$ (19)
of dual $q$-ary alternant codes of extension degree $m$, length $n$, and order $t$. If $q$ is unspecified we take $q=2$. If $n$ is unspecified we take $n=q^{m}$. We say a code ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$ is proper if it has dimension
$k=mt.$ (20)
In this case, after extension of scalars, we have
${\mathsf{C}}_{\mathbb{F}_{q^{m}}}=\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x},\mathbf{y})\oplus\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x}^{q},\mathbf{y}^{q})\oplus\cdots\oplus\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x}^{q^{m-1}},\mathbf{y}^{q^{m-1}}).$ (21)
Likewise we define the corresponding class
$\operatorname{{\mathsf{Gop}}}^{\perp}_{q,m,n,t}$ (22)
of dual Goppa codes, and ${\mathsf{C}}\in\operatorname{{\mathsf{Gop}}}^{\perp}_{q,m,n,t}$ is said proper if it is when seen in $\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$. Also we define subclasses $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}_{q,m,n,t}\subseteq\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}_{q,m,n,t}\subseteq\operatorname{{\mathsf{Gop}}}^{\perp}_{q,m,n,t}$, in which the Goppa polynomial is irreducible or squarefree, respectively.
2 Minimal resolutions and graded Betti numbers
Generalities
We freely borrow results and terminology from [9], and then elaborate on the parts of the theory that we will need.
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the $k$-dimensional polynomial ring over $\mathbb{F}$, graded by total degree. If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$. Iterating this construction, we obtain a minimal resolution
$\cdots\longrightarrow F_{2}\longrightarrow F_{1}\longrightarrow F_{0}$ (23)
of $M_{0}$, where the graded free modules $F_{i}$, together with the iterated syzygy modules $M_{i}$, are constructed inductively:
  • $F_{i}$ is the free module on a minimal system of homogeneous generators of $M_{i}$
  • $M_{i+1}=\ker(\;F_{i}\longrightarrow M_{i}\;)$.
Thus the $F_{i}$ are of the form
$F_{i}=\bigoplus_{j\geq 0}S(-j)^{\beta_{i,j}}$ (24)
where $S(-j)$ is the free $S$-module of rank $1$ generated in degree $j$, so $S(-j)_{d}=S_{d-j}$; and $\beta_{i,j}$, the $(i,j)$-th graded Betti number of $M_{0}$, is the number of elements of degree $j$ in a minimal system of generators of $M_{i}$.
From now on let ${\mathsf{C}}$ be a $[n,k]$-code, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring. Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$. So we have an exact sequence
$0\longrightarrow F_{k-1}\longrightarrow\cdots\longrightarrow F_{2}\longrightarrow F_{1}\longrightarrow F_{0}=S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}\longrightarrow 0$ (25)
with $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$, $F_{0}=S$, $M_{1}=I({\mathsf{C}})$, hence $\beta_{0,0}({\mathsf{C}})=1$, $\beta_{0,j}({\mathsf{C}})=0$ for $j\neq 0$, and $\beta_{1,j}({\mathsf{C}})$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$.
It is customary to display the graded Betti numbers in the form of a Betti diagram, as follows:
$\begin{array}[]{c|cccccc}&0&1&\dots&i&\dots&k-1\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&&\vdots\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots&\beta_{k-1,r+k-1}\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots&\beta_{k-1,r+k}\\ \vdots&\vdots&\vdots&&\vdots&&\vdots\\ \end{array}$
with null entries marked as “--” for readability. In our situation, as $F_{0}=S$, the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
By the minimality property of the resolution, if for some $i$ and $D$ we have $\beta_{i,j}=0$ for all $j\leq D$, then also $\beta_{i+1,j}=0$ for all $j\leq D+1$. Applied recursively, this means all the upper-right quadrant of the Betti diagram defined by $\beta_{i,D}$ vanishes. In particular, as the evaluation map $S_{1}\longrightarrow C$ is an isomorphism, we see that $I({\mathsf{C}})$ is generated in degrees $\geq 2$, which implies:
Lemma 1
For any $i\geq 1$, the $i$-th syzygy module $M_{i}$ of ${\mathsf{C}}$ is generated in degrees $\geq i+1$. Thus we have $F_{i,j}=M_{i,j}=0$ for $j\leq i$, and
$F_{i,i+1}=M_{i,i+1}=\mathbb{F}^{\beta_{i,i+1}}.$ (26)
Hence the 00-th row of the Betti diagram is $(1,-,-,\dots)$.
Figures 3-3 offer a few such diagrams for contemplation.
$\begin{array}[]{c|cccc}&0&1&2&3\\ \hline\cr 0&1&-&-&-\\ 1&-&3&-&-\\ 2&-&1&6&3\\ \end{array}$
Figure 1: the $[7,4]_{2}$ Hamming code
$\begin{array}[]{c|cccccc}&0&1&2&3&4&5\\ \hline\cr 0&1&-&-&-&-&-\\ 1&-&10&16&-&-&-\\ 2&-&1&5&26&20&5\\ \end{array}$
Figure 2: the $[11,6]_{3}$ Golay code
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
Figure 3: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
Effective computation
There are several algorithms to compute minimal resolutions and graded Betti numbers in general. Many of them rely first on a Gröbner basis computation. However in this work we will only be interested in computing the first (nontrivial) row of the Betti diagram, or equivalently, the so-called linear strand of the resolution. This easier computation can be described in elementary terms. First, for any $r\geq 3$, consider the natural multiplication map
$\varphi_{r}:M_{r-2,r-1}\otimes S_{1}\longrightarrow M_{r-2,r}.$ (27)
Lemma 2
We have:
$\ker(\varphi_{r})=M_{r-1,r}$ $\simeq\mathbb{F}^{\beta_{r-1,r}}$ (28)
$\operatorname{coker}(\varphi_{r})$ $\simeq\mathbb{F}^{\beta_{r-2,r}}.$ (29)
Proof
We prove (29) first. Let $\mathcal{G}$ be a minimal system of homogeneous generators of $M_{r-2}$, and for each $j$ let $B_{r-2,j}\subseteq M_{r-2,j}$ be the linear subspace generated by the degree $j$ elements of $\mathcal{G}$. Then, as $M_{r-2}$ is generated in degrees $\geq r-1$, we have $B_{r-2,r-1}=M_{r-2,r-1}$, and $B_{r-2,r}$ is a supplementary subspace to $S_{1}\cdot B_{r-2,r-1}=\operatorname{im}(\varphi_{r})$ in $M_{r-2,r}$. This proves (29).
Now let $F_{r-2}$ be the free graded module on $\mathcal{G}$. Then $M_{r-1,r}$ is the kernel of the natural map $F_{r-2,r}\longrightarrow M_{r-2,r}$. However, under the decompositions $F_{r-2,r}=(B_{r-2,r-1}\otimes S_{1})\oplus B_{r-2,r}$ and $M_{r-2,r}=\operatorname{im}(\varphi_{r})\oplus B_{r-2,r}$, this natural map decomposes as $\varphi_{r}\oplus\operatorname{id}_{B_{r-2,r}}$. This proves (28).
This readily gives a coarse upper bound on the $\beta_{r-1,r}$:
Lemma 3
We have $\beta_{1,2}\leq\frac{k(k-1)}{2}$, and $\beta_{r-1,r}\leq(k-1)\beta_{r-2,r-1}$ for $r\geq 3$. Hence $\beta_{r-1,r}({\mathsf{C}})\leq\frac{k}{2}(k-1)^{r-1}$ for any $r\geq 2$.
Proof
We have $\dim({\mathsf{C}}^{\langle 2\rangle})\geq\dim(C)=k$ hence $\beta_{1,2}=\binom{k+1}{2}-\dim({\mathsf{C}}^{\langle 2\rangle})\leq\frac{k(k-1)}{2}$. Now let $r\geq 3$. As $M_{r-2}$ is a submodule of the free module $F_{r-3}$, multiplication by $X_{1}$ is injective on $M_{r-2}$. Hence $M_{r-2,r-1}\simeq X_{1}M_{r-2,r-1}\subseteq\operatorname{im}(\varphi_{r})$ from which it follows $\beta_{r-1,r}=k\beta_{r-2,r-1}-\dim\operatorname{im}(\varphi_{r})\leq(k-1)\beta_{r-2,r-1}$.
Proposition 1
The $M_{r-1,r}$ can be computed iteratively as follows. For $r=3$:
$M_{2,3}=\ker(\;I_{2}({\mathsf{C}})\otimes S_{1}\longrightarrow S_{3}\;)$ (30)
where $I_{2}({\mathsf{C}})\otimes S_{1}\longrightarrow S_{3}$ is the natural multiplication map. Then for $r\geq 4$:
$M_{r-1,r}=\ker(\;M_{r-2,r-1}\otimes S_{1}\longrightarrow M_{r-3,r-2}\otimes S_{2}\;)$ (31)
where the map $\psi_{r}:M_{r-2,r-1}\otimes S_{1}\longrightarrow M_{r-3,r-2}\otimes S_{2}$ is obtained first by tensoring the inclusion $M_{r-2,r-1}\subseteq M_{r-3,r-2}\otimes S_{1}$ by $S_{1}$, and then composing with the multiplication map $S_{1}\otimes S_{1}\longrightarrow S_{2}$.
Proof
First, (30) is just the case $r=3$ of (28), composed with the inclusion $I_{3}({\mathsf{C}})\subseteq S_{3}$. Likewise to prove (31) we must show $\ker(\varphi_{r})=\ker(\psi_{r})$ for $r\geq 4$. For this we just observe that the diagram
[…]
commutes, with the two arrows on the right injective.
Thus we have two descriptions of $M_{r-1,r}$. The description $M_{r-1,r}=\ker(\varphi_{r})$ is closer to the abstract definition, and will be used later to estimate the value of the Betti numbers. The description $M_{r-1,r}=\ker(\psi_{r})$ is more amenable to effective computation.
Corollary 1
The $M_{r-1,r}({\mathsf{C}})$ only depend on $I_{2}({\mathsf{C}})$.
To make things even more explicit we introduce the following matrices. Assume ${\mathsf{C}}$ is given by a generator matrix $\mathbf{G}\in\mathbb{F}^{k\times n}$, with rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$. Choose monomial bases $\mathcal{M}_{1}=(X_{a})_{1\leq a\leq k}$, $\mathcal{M}_{2}=(X_{a}X_{b})_{1\leq a\leq b\leq k}$, and $\mathcal{M}_{3}=(X_{a}X_{b}X_{c})_{1\leq a\leq b\leq c\leq k}$ of $S_{1}$, $S_{2}$, $S_{3}$, ordered with respect to some monomial order.
Definition 1
The squared matrix of $\mathbf{G}$ is the matrix
$\mathbf{G}_{2}\in\mathbb{F}^{\,\binom{k+1}{2}\,\times\,n}$ (33)
with rows indexed by $\mathcal{M}_{2}$: the row corresponding to $X_{a}X_{b}$ is $\mathbf{c}_{a}\mathbf{c}_{b}$.
Then $I_{2}({\mathsf{C}})$ is the left kernel of $\mathbf{G}_{2}$, and we choose a basis $\mathcal{B}_{2}$ of this space. Thus $\mathcal{B}_{2}$ consists of $\beta_{1,2}$ vectors, each of which has its entries indexed by $\mathcal{M}_{2}$.
Definition 2
The degree $3$ Macaulay matrix of $\mathcal{B}_{2}$ is the matrix
$\mathbf{M}_{3}\in\mathbb{F}^{\>k\beta_{1,2}\,\times\,\binom{k+2}{3}}$ (34)
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 00 elsewhere.
Then by (30), $M_{2,3}$ is the left kernel of $\mathbf{M}_{3}$, and we choose a basis $\mathcal{B}_{3}$ of this space. Thus $\mathcal{B}_{3}$ consists of $\beta_{2,3}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$.
Now let $r\geq 4$, and assume inductively for all $13\leq i\leq r-$ 1 we have constructed a basis $\mathcal{B}_{i}$ of $M_{i-1,i}$, consisting of $\beta_{i-1,i}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{i-1}$.
Definition 3
The degree $r$ blockwise Macaulay matrix of $\mathcal{B}_{r-1},\mathcal{B}_{r-2}$ is the matrix
$\mathbf{M}_{r}\in\mathbb{F}^{\>k\beta_{r-2,r-1}\,\times\,\binom{k+1}{2}\beta_{r-3,r-2}}$ (35)
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 00 elsewhere.
Then by (31), $M_{r-1,r}$ is the left kernel of $\mathbf{M}_{r}$, and we proceed.
All this is summarized in Algorithm 1.
Algorithm 1 Compute bases $\mathcal{B}_{r}$ of $M_{r-1,r}({\mathsf{C}})$ up to some degree $D$
Input: a generator matrix $\mathbf{G}$ of the $[n,k]$-code ${\mathsf{C}}$ a target degree $D\leq k$
1:construct the matrix $\mathbf{G}_{2}$ according to Definition 1
2:compute a left kernel basis $\mathcal{B}_{2}$ of $\mathbf{G}_{2}$
3:construct the matrix $\mathbf{M}_{3}(\mathcal{B}_{2})$ according to Definition 2
4:compute a left kernel basis $\mathcal{B}_{3}$ of $\mathbf{M}_{3}$
5:for $r=4\dots D$ do
6:     construct the matrix $\mathbf{M}_{r}(\mathcal{B}_{r-1},\mathcal{B}_{r-2})$ according to Definition 3
7:     compute a left kernel basis $\mathcal{B}_{r}$ of $\mathbf{M}_{r}$
8:end for
Observe that our computation only relies on mere linear algebra, and does not make use of any Gröbner basis theory. However, the two topics are clearly related. In particular, the many linear algebra optimizations used in Gröbner basis algorithms could certainly apply in our context.
Further properties
We will have a particular interest in the length of the linear strand, or equivalently, in the following quantity:
Definition 4
For a linear code ${\mathsf{C}}$, we set
$r_{\max}({\mathsf{C}})=\max\{r:\>\beta_{r-1,r}({\mathsf{C}})>0\}.$ (36)
As the minimal resolution of a $[n,k]$-code ${\mathsf{C}}$ has length $k-1$, we always have $r_{\max}({\mathsf{C}})\leq k$.
Sometimes we only control the syzygies of certain subideals of $I({\mathsf{C}})$, and from these we want to deduce information on the syzygies of the whole ideal. Intuitively, we expect syzygies between elements of the subideals could be seen as syzygies between elements of the ideal. Of course this fails in general, but it remains true if we restrict to the linear strand:
Proposition 2
Let $M$ be a finite $S$-module generated in degrees $\geq d_{0}$, and suppose its degree $d_{0}$ part $M_{d_{0}}$ decomposes as a direct sum
$M_{d_{0}}=A_{d_{0}}\oplus B_{d_{0}}.$ (37)
Let $A=\langle A_{d_{0}}\rangle_{S}$ and $B=\langle B_{d_{0}}\rangle_{S}$ be the sub-$S$-modules of $M$ generated by $A_{d_{0}}$ and $B_{d_{0}}$. Let $\widehat{M}$ be the (first) syzygy module of $M$, and $\widehat{A}$ and $\widehat{B}$ those of $A$ and $B$. Then we have a natural inclusion
$\widehat{A}_{d_{0}+1}\oplus\widehat{B}_{d_{0}+1}\;\subseteq\;\widehat{M}_{d_{0}+1}$ (38)
the cokernel of which identifies with $A_{d_{0}+1}\cap B_{d_{0}+1}$, hence a (non-canonical) isomorphism
$\widehat{M}_{d_{0}+1}\;\simeq\;\widehat{A}_{d_{0}+1}\oplus\widehat{B}_{d_{0}+1}\oplus(A_{d_{0}+1}\cap B_{d_{0}+1}).$ (39)
Proof
We have surjective maps
$(A_{d_{0}}\oplus B_{d_{0}})\otimes S_{1}\;\twoheadrightarrow\;A_{d_{0}+1}\oplus B_{d_{0}+1}\;\twoheadrightarrow\;A_{d_{0}+1}+B_{d_{0}+1}$ (40)
with $\widehat{A}_{d_{0}+1}\oplus\widehat{B}_{d_{0}+1}$ the kernel of the leftmost map, $A_{d_{0}+1}\cap B_{d_{0}+1}$ the kernel of the rightmost map, and $\widehat{M}_{d_{0}+1}$ the kernel of the composite map. We conclude with the associated kernel-cokernel exact sequence.
Corollary 2
Suppose $I_{2}({\mathsf{C}})$ contains a certain direct sum of $l$ subspaces:
$I_{2}({\mathsf{C}})\supseteq V^{(1)}\oplus\cdots\oplus V^{(l)}.$ (41)
Then the minimal resolution of ${\mathsf{C}}^{\langle\cdot\rangle}$ canonically admits the direct sum of the linear strands of the minimal resolutions of $S/\langle V^{(1)}\rangle$, …, $S/\langle V^{(l)}\rangle$ as a subcomplex. In particular for all $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\beta_{r-1,r}(S/\langle V^{(1)}\rangle)+\cdots+\beta_{r-1,r}(S/\langle V^{(l)}\rangle).$ (42)
Proof
Apply Proposition 2 repeatedly.
Recall that $\pi{\mathsf{C}}$ is a punctured code of ${\mathsf{C}}$ if it is obtained from ${\mathsf{C}}$ by discarding some given set of coordinates (equivalently, discarding some subset of the associated set of points in projective space).
Corollary 3
Let $\pi{\mathsf{C}}$ be a punctured code of ${\mathsf{C}}$. Assume $\dim(\pi{\mathsf{C}})=\dim({\mathsf{C}})$. Then for all $r\geq 2$ we have
$\beta_{r-1,r}(\pi{\mathsf{C}})\geq\beta_{r-1,r}({\mathsf{C}})$ (43)
hence
$r_{\max}(\pi{\mathsf{C}})\geq r_{\max}({\mathsf{C}}).$ (44)
Proof
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ where $k=\dim(\pi{\mathsf{C}})=\dim({\mathsf{C}})$. Then the surjective map $S\longrightarrow(\pi{\mathsf{C}})^{\langle\cdot\rangle}$ factors through ${\mathsf{C}}^{\langle\cdot\rangle}$, so that $I_{2}(\pi{\mathsf{C}})\supseteq I_{2}({\mathsf{C}})$.
Last, the following easy result is also useful:
Lemma 4
Minimal resolutions are preserved under extension of scalars. In particular if a code ${\mathsf{C}}$ is defined over $\mathbb{F}$, and if $\mathbb{F}\subseteq\mathbb{K}$ is a field extension, then $\beta_{i,j}({\mathsf{C}})=\beta_{i,j}({\mathsf{C}}_{\mathbb{K}})$ for all $i,j$.
3 Lower bounds from the Eagon-Northcott complex
Generalities
Let $R$ be a ring, and for integers $f\geq g$, temporarily switching to column vector convention, let $\mathbf{\Phi}\in R^{g\times f}$ define a linear map
$\mathbf{\Phi}:F=R^{f}\longrightarrow G=R^{g}.$ (45)
The Eagon-Northcott complex [8] of $\mathbf{\Phi}$ is the following complex of $R$-modules, defined in terms of the exterior and (dual of) symmetric powers of $F$ and $G$:
$0\to(\operatorname{Sym}^{f-g}G)^{\vee}\otimes\bigwedge^{f}F\to\cdots\to G^{\vee}\otimes\bigwedge^{g+1}F\to\bigwedge^{g}F\overset{\bigwedge^{g}\mathbf{\Phi}}{\longrightarrow}\bigwedge^{g}G\simeq R.$ (46)
Under mild hypotheses, this complex is exact [9, Th. A.2.60 & Th. 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
In this work we will only need the case $g=2$. In this case the complex has length $f-1$, and for $2\leq r\leq f$ its $(r-1)$-th module is free of rank
$\operatorname{rk}\left((\operatorname{Sym}^{r-2}G)^{\vee}\otimes\bigwedge^{r}F\right)\;=\;(r-1)\binom{f}{r}.$ (47)
All this can be made explicit, in coordinates. Let
$\mathbf{\Phi}=\left(\begin{array}[]{cccc}x_{1}&x_{2}&\dots&x_{f}\\ x^{\prime}_{1}&x^{\prime}_{2}&\dots&x^{\prime}_{f}\end{array}\right)$ (48)
for $x_{i},x^{\prime}_{i}\in R$. Then:
  • $\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ is generated by the $\binom{f}{2}$ minors
    $q_{i,j}=x_{i}x^{\prime}_{j}-x_{j}x^{\prime}_{i}$ (49)
    for $1\leq i<j\leq f$,
  • these $q_{ij}$ are annihilated by the $2\binom{f}{3}$ relations
    $\begin{split}r_{ijk}&=x_{i}q_{jk}-x_{j}q_{ik}+x_{k}q_{ij}\\ r^{\prime}_{ijk}&=x^{\prime}_{i}q_{jk}-x^{\prime}_{j}q_{ik}+x^{\prime}_{k}q_{ij}\end{split}$ (50)
    for $1\leq i<j<k\leq f$,
  • these $r_{ijk}$ and $r^{\prime}_{ijk}$ are annihilated by the $3\binom{f}{4}$ relations
[…]
    for $1\leq i<j<k<l\leq f$,
and so on. These are just the cases $r=2,3,4$ of the more general:
Proposition 3
In the multivariate polynomial ring
$R[Z^{(j)}_{r;i_{1},\dots,i_{r}}]$ (52)
where, for each $2\leq r\leq f$, indices range over the $(r-1)\binom{f}{r}$ values $11\leq j\leq r-$ 1 and $1\leq i_{1}<\dots<i_{r}\leq f$, construct polynomials $s^{(j)}_{r;i_{1},\dots,i_{r}}$ as follows. First for $r=2$ we define the constant polynomials
$s^{(1)}_{2;i_{1},i_{2}}=x_{i_{1}}x^{\prime}_{i_{2}}-x^{\prime}_{i_{1}}x_{i_{2}}$ (53)
given by the minors of $\mathbf{\Phi}$; and for $r\geq 3$ we define linear polynomials
$s^{(j)}_{r;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z^{(j)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z^{(j-1)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
where we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
$\mathbf{s}_{r}(\mathbf{s}_{r-1})=\mathbf{0}.$ (55)
Proof
Either express (46) in the standard bases of $F=R^{f}$ and $G=R^{2}$. Or more directly, consider the $\mathbf{Z}_{r}$ and $\mathbf{s}_{r}$ as $(r-1)$-tuples of (exterior) $r$-vectors, observe that (53) means $\mathbf{s}^{(1)}_{2}=\mathbf{x}\wedge\mathbf{x}^{\prime}$, that (54) means $\mathbf{s}^{(j)}_{r}=\mathbf{x}\wedge\mathbf{Z}^{(j)}_{r-1}+\mathbf{x}^{\prime}\wedge\mathbf{Z}^{(j-1)}_{r-1}$, so the alternating property gives $\mathbf{s}^{(1)}_{3}(\mathbf{s}_{2})=\mathbf{x}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, $\;\mathbf{s}^{(2)}_{3}(\mathbf{s}_{2})=\mathbf{x}^{\prime}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, and for $r\geq 4$:
[…]
Shortening (or projection from a point)
We have interest in how this Eagon-Northcott complex interacts with reduction to a subcode, and in particular with shortening. By induction it suffices to consider the case of a codimension $1$ subcode, or a $1$-shortening respectively.
So let ${\mathsf{C}}$ be a $[n,k]$-code over $\mathbb{F}$, and ${\mathsf{C}}_{\mathcal{H}}$ a codimension $1$ subcode. We will assume ${\mathsf{C}}$ projective, but then ${\mathsf{C}}_{\mathcal{H}}$ need not be so. Choose a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k-1}$ of ${\mathsf{C}}_{\mathcal{H}}$, and complete it to a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of ${\mathsf{C}}$. Let $\mathbf{G}_{\mathcal{H}}$ and $\mathbf{G}$ be the corresponding generating matrices of ${\mathsf{C}}_{\mathcal{H}}$ and ${\mathsf{C}}$. Let also $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the polynomial ring in $k$ indeterminates, and $S_{\mathcal{H}}=\mathbb{F}[X_{1},\dots,X_{k-1}]$ its subring in the first $k-1$ indeterminates. We have a commutative diagram
[…]
where horizontal maps denote evaluation, and vertical maps inclusion. Also set
$\mathcal{H}=S_{\mathcal{H},1}=\langle X_{1},\dots,X_{k-1}\rangle_{\mathbb{F}}\subseteq S_{1}$ (58)
and let $\mathbf{p}_{\mathcal{H}}=(0:\dots:0:1)^{\top}\in\mathbf{P}^{k-1}$ be the associated point. Last, let $\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}\subseteq\mathbf{P}^{k-1}$ be the set of points defined by the columns of $\mathbf{G}$. Then we have two possibilities:
  • Either $\mathbf{c}_{1},\dots,\mathbf{c}_{k-1}$ all vanish at some common position $i$, or equivalently, $\mathbf{p}_{\mathcal{H}}=\mathbf{p}_{i}$. Puncturing this position, we can identify ${\mathsf{C}}_{\mathcal{H}}$ with the $1$-shortened subcode of ${\mathsf{C}}$ at $i$. Then the set of points in $\mathbf{P}^{k-2}$ defined by the columns of $\mathbf{G}_{\mathcal{H}}$ is the image of $\{\mathbf{p}_{1},\dots,\widehat{\mathbf{p}}_{i},\dots,\mathbf{p}_{n}\}$ under the projection from $\mathbf{p}_{i}$.
  • Otherwise, if there is no such $i$, then ${\mathsf{C}}_{\mathcal{H}}$ is a “general” codimension $1$ subcode, and the set of points in $\mathbf{P}^{k-2}$ defined by the columns of $\mathbf{G}_{\mathcal{H}}$ is the image of $\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$ under the projection from $\mathbf{p}_{\mathcal{H}}$.
(As already observed, ${\mathsf{C}}_{\mathcal{H}}$ need not be projective, so these image points need not be distinct.)
Let $f\geq g\geq d\geq 2$ be integers, and let $\mathbf{\Phi}\in S_{1}^{g\times f}$ be a matrix whose entries are homogeneous linear forms, i.e. elements of $S_{1}$. Recall from [9, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero. Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
$V_{\mathbf{\Phi},\mathcal{H}}=V_{\mathbf{\Phi}}\cap\mathcal{H}^{g}.$ (59)
Set $f_{\mathcal{H}}=\dim(V_{\mathbf{\Phi},\mathcal{H}})\geq\dim(V_{\mathbf{\Phi}})-g$, and then let $\mathbf{\Phi}_{\mathcal{H}}\in\mathcal{H}^{g\times f_{\mathcal{H}}}$ be a matrix whose columns form a basis of $V_{\mathbf{\Phi},\mathcal{H}}$.
Proposition 4
Under these hypotheses:
  1. 1.
    If $I_{d}({\mathsf{C}})$ contains the $d\times d$ minors of $\mathbf{\Phi}$, then $I_{d}({\mathsf{C}}_{\mathcal{H}})$ contains the $d\times d$ minors of $\mathbf{\Phi}_{\mathcal{H}}$.
  2. 2.
    If $\mathbf{\Phi}$ is $1$-generic, then $\mathbf{\Phi}_{\mathcal{H}}$ is $1$-generic, and $f_{\mathcal{H}}\geq f-g$.
  3. 3.
    If $I_{d}({\mathsf{C}})$ contains the $d\times d$ minors of $\mathbf{\Phi}$, if $\mathbf{\Phi}$ is $1$-generic, and if ${\mathsf{C}}_{\mathcal{H}}$ is a $1$-shortened subcode of ${\mathsf{C}}$, then $f_{\mathcal{H}}\geq f-d+1$.
Proof
By (57) we have $I_{d}({\mathsf{C}}_{\mathcal{H}})=I_{d}({\mathsf{C}})\cap S_{\mathcal{H}}$. Columns of $\mathbf{\Phi}_{\mathcal{H}}$ belong to the column span of $\mathbf{\Phi}$, so $d\times d$ minors of $\mathbf{\Phi}_{\mathcal{H}}$ are linear combinations of $d\times d$ minors of $\mathbf{\Phi}$. On the other hand, $\mathbf{\Phi}_{\mathcal{H}}$ has coefficients in $\mathcal{H}$, so its $d\times d$ minors belong to $S_{\mathcal{H}}$. This proves 1.
Assume $\mathbf{\Phi}$ is $1$-generic. As columns of $\mathbf{\Phi}_{\mathcal{H}}$ are linearly independent and belong to the column space of $\mathbf{\Phi}$, we can write $\mathbf{\Phi}_{\mathcal{H}}=\mathbf{\Phi}\mathbf{M}$ where $\mathbf{M}\in\mathbb{F}^{f\times f_{\mathcal{H}}}$ has $\operatorname{rk}(\mathbf{M})=f_{\mathcal{H}}$. Then for $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f_{\mathcal{H}}}$ nonzero we have $\mathbf{a}\mathbf{\Phi}_{\mathcal{H}}\mathbf{b}^{\top}=\mathbf{a}\mathbf{\Phi}\mathbf{M}\mathbf{b}^{\top}\neq 0$ because $\mathbf{M}\mathbf{b}^{\top}$ is nonzero. This proves that $\mathbf{\Phi}_{\mathcal{H}}$ is $1$-generic. Then it is easily seen that a $1$-generic matrix has linearly independent columns, so $\dim(V_{\mathbf{\Phi}})=f$, hence $f_{\mathcal{H}}\geq f-g$. This finishes the proof of 2.
Now we prove 3. Assume ${\mathsf{C}}_{\mathcal{H}}$ is the $1$-shortened subcode of ${\mathsf{C}}$ at position $i$, hence $\mathbf{p}_{i}=\mathbf{p}_{\mathcal{H}}=(0:\dots:0:1)^{\top}$. As $\mathbf{\Phi}$ is $1$-generic, we have $\dim(V_{\mathbf{\Phi}})=f$. Let $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}:S\longrightarrow\mathbb{F}$ denote evaluation at $\mathbf{p}_{\mathcal{H}}$. Applied coordinatewise, we also have evaluation maps $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}:S^{g}\longrightarrow\mathbb{F}^{g}$ and $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}:S^{g\times f}\longrightarrow\mathbb{F}^{g\times f}$. Also we can restrict $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}$ to subspaces, so for instance $\mathcal{H}=\ker(\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}:S_{1}\longrightarrow\mathbb{F})$. It then follows
$V_{\mathbf{\Phi},\mathcal{H}}=V_{\mathbf{\Phi}}\cap\mathcal{H}^{g}=\ker(\,\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}:V_{\mathbf{\Phi}}\longrightarrow\mathbb{F}^{g}\,),$ (60)
while the image $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}(V_{\mathbf{\Phi}})\subseteq\mathbb{F}^{g}$ is the column span of $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}(\mathbf{\Phi})\in\mathbb{F}^{g\times f}$. As the $d\times d$ minors of $\mathbf{\Phi}$ belong to $I_{d}({\mathsf{C}})$, they vanish at each column of $\mathbf{G}$, in particular they vanish at $\mathbf{p}_{\mathcal{H}}=\mathbf{p}_{i}$. But this means precisely that the $d\times d$ minors of $\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}(\mathbf{\Phi})$ all vanish, or equivalently,
$\dim(\operatorname{ev}_{\mathbf{p}_{\mathcal{H}}}(V_{\mathbf{\Phi}}))\leq d-1.$ (61)
Joint with (60) this gives $f_{\mathcal{H}}=\dim(V_{\mathbf{\Phi},\mathcal{H}})\geq\dim(V_{\mathbf{\Phi}})-(d-1)=f-d+1$.
Application to alternant and Goppa codes
Consider a code ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$, assumed to be proper. By (21), after extension of scalars, ${\mathsf{C}}_{\mathbb{F}_{q^{m}}}$ admits as a basis the $k=mt$ vectors $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$ for $10\leq a\leq t-$ 1 and $10\leq u\leq m-$ 1. Rename the $k=mt$ variables in our polynomial ring $S$ accordingly, so that our evaluation map now is
$S=\mathbb{F}_{q^{m}}[X^{(u)}_{a}]\longrightarrow{\mathsf{C}}_{\mathbb{F}_{q^{m}}}^{\langle\cdot\rangle}$ (62)
where $X^{(u)}_{a}$ evaluates as $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$.
If $\mathbf{M}$ is a matrix (or more generally an expression) in the $X^{(u)}_{a}$, we denote by $\mathbf{M}^{(v)}$ the same matrix (or expression) with each $X^{(u)}_{a}$ replaced by $X^{(u+v)}_{a}$, where $u+v$ is considered mod $m$. We also write $\mathbf{M}^{\prime}=\mathbf{M}^{(1)}$, $\mathbf{M}^{\prime\prime}=\mathbf{M}^{(2)}$, etc.
Set
$e=\lfloor\log_{q}(t-1)\rfloor,$ (63)
and for any $0\leq u\leq e$ define a $2\times(t-q^{u})$ matrix
$\mathbf{B}_{u}=\left(\begin{array}[]{cccc}X^{(0)}_{0}&X^{(0)}_{1}&\dots&X^{(0)}_{t-1-q^{u}}\\ X^{(0)}_{q^{u}}&X^{(0)}_{q^{u}+1}&\dots&X^{(0)}_{t-1}\end{array}\right).$ (64)
We then define the block matrix
$\mathbf{\Phi}=\left(\begin{array}[]{c|c|c|c}\mathbf{B}^{(e)}_{0}&\mathbf{B}^{(e-1)}_{1}&\cdots&\mathbf{B}^{(0)}_{e}\end{array}\right)$ (65)
of total size $2\times f$ where $f=(e+1)t-\frac{q^{e+1}-1}{q-1}$. Observe that this matrix only depends on $q$ and $t$, debatably on $m$, but certainly not on $n$ nor on the specific choice of ${\mathsf{C}}$.
Lemma 5
The $2\times 2$ minors of $\mathbf{\Phi}$ belong to $I_{2}({\mathsf{C}}_{\mathbb{F}_{q^{m}}})$.
Proof
The $2\times 2$ minor defined by the columns $\left[\begin{array}[]{c}X^{(e-u)}_{a}\\ X^{(e-u)}_{q^{u}+a}\end{array}\right]$ of $\mathbf{B}^{(e-u)}_{u}$ and $\left[\begin{array}[]{c}X^{(e-v)}_{b}\\ X^{(e-v)}_{q^{v}+b}\end{array}\right]$ of $\mathbf{B}^{(e-v)}_{v}$ is
$X^{(e-u)}_{a}X^{(e-v)}_{q^{v}+b}-X^{(e-u)}_{q^{u}+a}X^{(e-v)}_{b}$ (66)
which evaluates under (62) to
$(\mathbf{y}\mathbf{x}^{a})^{q^{e-u}}(\mathbf{y}\mathbf{x}^{q^{v}+b})^{q^{e-v}}-(\mathbf{y}\mathbf{x}^{q^{u}+a})^{q^{e-u}}(\mathbf{y}\mathbf{x}^{b})^{q^{e-v}}=\mathbf{0}.$ (67)
Lemma 6
The matrix $\mathbf{\Phi}$ is $1$-generic.
Proof
Let $\mathbf{a}\in\overline{\mathbb{F}}_{q}^{2}$ and $\mathbf{b}\in\overline{\mathbb{F}}_{q}^{f}$ be nonzero. Let $a_{i}$ be the rightmost nonzero entry of $\mathbf{a}$ and $b_{j}$ the rightmost nonzero entry of $\mathbf{b}$. Then the $X_{c}^{(u)}$ variable that multiplies $a_{i}b_{j}$ in $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ does not appear elsewhere in $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$, so $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}\neq 0$.
Theorem 3.1
Let ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$ be proper, $e=\lfloor\log_{q}(t-1)\rfloor$, $f=(e+1)t-\frac{q^{e+1}-1}{q-1}$. For any $s\geq 0$, let ${\mathsf{C}}_{s}$ be a $s$-shortened subcode of ${\mathsf{C}}$. Then for all $r\geq 2$ we have $\beta_{r-1,r}({\mathsf{C}}_{s})\geq(r-1)\binom{f-s}{r}$, hence
$r_{\max}({\mathsf{C}}_{s})\geq f-s.$ (68)
Proof
By Lemma 4 we can extend scalars to $\mathbb{F}_{q^{m}}$. By Lemma 5, $I_{2}({\mathsf{C}})$ then contains the $2\times 2$ minors of the $2\times f$ matrix $\mathbf{\Phi}$, and this matrix $\mathbf{\Phi}$ is $1$-generic by Lemma 6. By Proposition 4 applied $s$ times (with $g=d=2$), $I_{2}({\mathsf{C}}_{s})$ then contains the $2\times 2$ minors of a $2\times f_{s}$ matrix $\mathbf{\Phi}_{s}$, where $f_{s}\geq f-s$, and this matrix $\mathbf{\Phi}_{s}$ is $1$-generic also. Then by [9, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$. We then conclude with Corollary 2 (with $l=1$ and $V^{(1)}=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi}_{s})$).
Remark 1
It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$. See the corresponding discussion in the Supplementary material.
Examples show the bound (68) on $r_{\max}$ is tight in general for alternant codes. However one can improve it in the Goppa case. Let us focus on $q=2$ for simplicity. Let $\varphi:a\mapsto a^{2}$ be the Frobenius map, acting on any $\mathbb{F}_{2}$-algebra. For any polynomial $g\in\mathbb{F}_{2^{m}}[X]$, set $L_{g}=g(X)^{-1}\mathbb{F}_{2^{m}}[X]_{<\deg(g)}\subseteq\mathbb{F}_{2^{m}}(X)$.
Lemma 7
Let $g(X)\in\mathbb{F}_{2^{m}}[X]$ be squarefree (i.e. separable). Then
$L_{g}\,+\,\varphi(L_{g})\;=\;L_{g^{2}},$ (69)
the sum being direct
Proof
Set $t=\deg(g)$. Then $L_{g}$ and $\varphi(L_{g})$ both are $t$-dimensional subspaces of the $2t$-dimensional space $L_{g^{2}}$. To conclude we only have to prove $L_{g}\cap\varphi(L_{g})=0$. However we have $F(X)\in L_{g}\cap\varphi(L_{g})$ if and only if $F(X)=\frac{A(X)}{g(X)}=\frac{B(X)^{2}}{g(X)^{2}}$ for some $A,B$ of degree $<t$. But then this implies $g(X)|B(X)^{2}$ with $g$ squarefree of degree $t$, which is impossible unless $B=0$.
From this we readily deduce:
Lemma 8
Let $\mathbf{x}\in(\mathbb{F}_{2^{m}})^{n}$ be a support, and $g(X)\in\mathbb{F}_{2^{m}}[X]$ squarefree of degree $t\leq n/2$, not vanishing on any entry of $\mathbf{x}$. Then
$\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x},g(\mathbf{x})^{-1})+\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x}^{2},g(\mathbf{x})^{-2})=\operatorname{{\mathsf{GRS}}}_{2t}(\mathbf{x},g(\mathbf{x})^{-2}),$ (70)
the sum being direct.
Proposition 5
Let ${\mathsf{C}}=\operatorname{{\mathsf{Gop}}}(\mathbf{x},g)^{\perp}\in\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}_{2,m,n,t}$ be proper, with $g$ squarefree. Set $\mathbf{y}=g(\mathbf{x})^{-1}$. Then
${\mathsf{C}}_{\mathbb{F}_{q^{m}}}=\bigoplus_{i=0}^{m/2-1}\operatorname{{\mathsf{GRS}}}_{2t}(\mathbf{x}^{4^{i}},\mathbf{y}^{4^{i}}),\quad\text{or}$ (71)
${\mathsf{C}}_{\mathbb{F}_{q^{m}}}=\left(\bigoplus_{i=0}^{(m-1)/2-1}\operatorname{{\mathsf{GRS}}}_{2t}(\mathbf{x}^{4^{i}},\mathbf{y}^{4^{i}})\right)\oplus\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x}^{2^{m-1}},\mathbf{y}^{2^{m-1}})$ (72)
depending on whether $m$ is even or odd.
Corollary 4
Let ${\mathsf{C}}=\operatorname{{\mathsf{Gop}}}(\mathbf{x},g)^{\perp}\in\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}_{2,m,n,t}$ be proper, with $g$ squarefree. Set $\widehat{e}=\lfloor\log_{4}(2t-1)\rfloor$ and $\widehat{f}=(2\widehat{e}+2)t-\frac{4^{\widehat{e}+1}-1}{3}$. For any $s\geq 0$, let ${\mathsf{C}}_{s}$ be a $s$-shortened subcode of ${\mathsf{C}}$. Then for all $r\geq 2$ we have $\beta_{r-1,r}({\mathsf{C}}_{s})\geq(r-1)\binom{\widehat{f}-s}{r}$, hence
$r_{\max}({\mathsf{C}}_{s})\geq\widehat{f}-s.$ (73)
Proof
Same as Theorem 3.1, with $\mathbf{\Phi}$ adapted to fit Proposition 5.
4 Regularity $2$ and the small defect heuristic
Regularity $2$ and consequences
If ${\mathsf{C}}$ is a $[n,k]$-code, we let
$B_{j}=\sum_{i\geq 0}(-1)^{i}\beta_{i,j}$ (74)
be the alternating sum of its Betti numbers degree $j$, and $B(z)=\sum_{j\geq 0}B_{j}z^{j}$ their generating polynomial (it is indeed a finite sum).
Let also $H_{{\mathsf{C}}}(z)=\sum_{r\geq 0}z^{r}\dim{\mathsf{C}}^{\langle r\rangle}$ be the Hilbert series of ${\mathsf{C}}$.
Proposition 6
We have
$B(z)=(1-z)^{k}H_{{\mathsf{C}}}(z).$ (75)
Proof
Generating series reformulation of [9, Cor. 1.10].
Definition 5 ([25, Def. 1.5 & Th. 2.35])
The Castelnuovo-Mumford regularity of a projective $[n,k]$-code ${\mathsf{C}}$ is the smallest integer $r$ such that ${\mathsf{C}}^{\langle r\rangle}=\mathbb{F}^{n}$.
Definition 6 (cf. [9, §4A], after [22, Lect. 14])
The Castelnuovo-Mumford regularity of ${\mathsf{C}}^{\langle\cdot\rangle}$ is $\max\{r:\>\exists i,\,\beta_{i,i+r}({\mathsf{C}})>0\}$.
Proposition 7 ([9, Th. 4.2])
These two definitions coincide.
The square code distinguisher, and the filtration attack from [1] that extends it, work for codes ${\mathsf{C}}$ with ${\mathsf{C}}^{\langle 2\rangle}\varsubsetneq\mathbb{F}^{n}$, i.e. of regularity $>2$. This means that codes of regularity $2$ are hard to deal with under this approach. On the opposite, for us, codes of regularity $2$ are nice because Definition 6 means their Betti diagram is simple: it has only two nontrivial rows. Observe that most codes of interest have regularity $2$ (with the notable exception of self-dual codes).
Definition 7
If $f:U\longrightarrow V$ is a linear map between finite dimensional $\mathbb{F}$-vector spaces, we define its index
$\begin{split}\operatorname{ind}(f)&=\dim(U)-\dim(V)\\ &=\dim\ker(f)-\dim\operatorname{coker}(f)\end{split}$ (76)
and its defect
$\begin{split}\operatorname{def}(f)&=\min(\dim(U),\dim(V))-\operatorname{rk}(f)\\ &=\min(\dim\ker(f),\dim\operatorname{coker}(f)).\end{split}$ (77)
For any real $x$ we set $x^{+}=\max(x,0)$ and $x^{-}=(-x)^{+}$, so $x=x^{+}-x^{-}$. Then we always have
$\dim\ker(f)$ $\geq\operatorname{ind}(f)^{+}$ (78)
$\dim\operatorname{coker}(f)$ $\geq\operatorname{ind}(f)^{-}$ (79)
and then
$\operatorname{def}(f)=\dim\ker(f)-\operatorname{ind}(f)^{+}=\dim\operatorname{coker}(f)-\operatorname{ind}(f)^{-}$ (80)
measures the distance to equality in these two inequalities.
For $r\geq 3$, recall the linear map $\varphi_{r}:M_{r-2,r-1}\otimes S_{1}\longrightarrow M_{r-2,r}$ from (27). It will also be handy to let $\varphi_{2}=\operatorname{ev}_{\mathbf{G},2}:S_{2}\longrightarrow{\mathsf{C}}^{\langle 2\rangle}$ be the evaluation map.
Theorem 4.1
For all $r\geq 2$ we have
$\dim\ker(\varphi_{r})$ $=\beta_{r-1,r}$ (81)
$\dim\operatorname{coker}(\varphi_{r})$ $=\beta_{r-2,r}$ (82)
and moreover, if ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then
$\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}.$ (83)
Proof
For $r=2$ this is proved directly. Now assume $r\geq 3$. Then the first two equalities are reformulations of Lemma 2. Now assume moreover ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$. Then $H_{{\mathsf{C}}}(z)=1+kz+n\frac{z^{2}}{1-z}$, so $B(z)=(1+kz)(1-z)^{k}+nz^{2}(1-z)^{k-1}$ by Proposition 6, thus $B_{r}=(-1)^{r-1}(r-1)\binom{k+1}{r}+(-1)^{r}n\binom{k-1}{r-2}=(-1)^{r-1}\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$. On the other hand, ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$ also means ${\mathsf{C}}$ has regularity $2$ in the sense of Definition 6, so (74) reduces to $B_{r}=(-1)^{r-1}\beta_{r-1,r}+(-1)^{r-2}\beta_{r-2,r}=(-1)^{r-1}\operatorname{ind}(\varphi_{r})$ and we conclude.
Corollary 5
If ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then the bottom right entry of its Betti diagram is
$\beta_{k-1,k+1}=n-k.$ (84)
Proof
The minimal resolution of ${\mathsf{C}}$ has length $k-1$, so $\beta_{i,j}=0$ for $i>k-1$, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
Regularity $2$ helps in computing the full Betti diagram. We already gave examples in Figures 3-3. In order to provide a few more, let us first recall that a $[n,k]$ code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
Lemma 9 ([14, Th. 1], reformulated)
Let ${\mathsf{C}}$ be a $[n,k]$-code with $n\leq 2k-1$. Assume ${\mathsf{C}}$ is MDS. Then for all $r\leq 2k+1-n$ we have $\beta_{r-2,r}({\mathsf{C}})=0$.
Proposition 8
Let ${\mathsf{C}}$ be a $[k+1,k]$ MDS code, for instance a parity code or a $[k+1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are $\beta_{0,0}=\beta_{k-1,k+1}=1$, and
$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (85)
for $12\leq r\leq k-$ 1. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
Proof
The parameters imply that ${\mathsf{C}}$ has regularity $2$. By Lemma 9 we have $\beta_{r-2,r}=0$ for $r\leq k$. Then $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-(k+1)\right)\binom{k-1}{r-2}$ and we conclude with a straightforward calculation.
Proposition 9
Let ${\mathsf{C}}$ be a $[2k-1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are:
  • $\beta_{0,0}=1$
  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
In particular the ideal $I({\mathsf{C}})$ is generated by $\beta_{1,2}=\frac{(k-1)(k-2)}{2}$ quadratic forms and $\beta_{1,3}=k-1$ cubic forms.
Proof
Using the basis $\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}$ of ${\mathsf{C}}$, we see $I_{2}({\mathsf{C}})$ contains the $2\times 2$ minors of $\mathbf{\Phi}=\left(\begin{array}[]{cccc}X_{0}&X_{1}&\dots&X_{k-2}\\ X_{1}&X_{2}&\dots&X_{k-1}\end{array}\right)$, i.e. $I_{2}({\mathsf{C}})\supseteq\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ of dimension $\binom{k-1}{2}$.
Now set $n=2k-1$.
Because $n\geq 2k-1$, we have $\dim{\mathsf{C}}^{\langle 2\rangle}=2k-1$, hence $\dim I_{2}({\mathsf{C}})=\binom{k-1}{2}$, so $I_{2}({\mathsf{C}})=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$. Corollary 1 and the Eagon-Northcott complex of $\mathbf{\Phi}$ then give $\beta_{r-1,r}=(r-1)\binom{k-1}{r}$.
Because $n\leq 2k-1$, we have ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$. Then $\beta_{r-2,r}=\beta_{r-1,r}-\operatorname{ind}(\varphi_{r})=(r-1)\binom{k-1}{r}-\left(\frac{k(k+1)}{r}-(2k-1)\right)\binom{k-1}{r-2}$, and a straightforward calculation allows to conclude.
Figures 5 and 5 illustrate Propositions 8 and 9.
$\begin{array}[]{c|cccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&27&105&189&189&105&27&-\\ 2&-&-&-&-&-&-&-&1\end{array}$
Figure 4: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
$\begin{array}[]{c|ccccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&21&70&105&84&35&6&-\\ 2&-&7&42&105&140&105&42&7\end{array}$
Figure 5: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
Exercise 1
Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for $k+1<n<2k-1$.
The small defect heuristic
A direct application of Definition 7 and Theorem 4.1 gives:
Proposition 10
If ${\mathsf{C}}$ is a $[n,k]$-code with ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then for all $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r})$ (86)
$\beta_{r-2,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{-}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r}).$ (87)
It is tempting to see this result as an estimate for the Betti numbers of random codes, with a leading term given by $\operatorname{ind}(\varphi_{r})^{\pm}$ (compare with the bound in Lemma 3), and a secondary term given by $\operatorname{def}(\varphi_{r})$. Indeed, it is well known that under a reasonable probability model, a random linear map tends to have a small defect. Unfortunately, if ${\mathsf{C}}$ is a random code (say, uniform among codes of given $[n,k]$), it is not easy to control the distribution of the maps $\varphi_{r}$.
Still, in the special case $r=2$, [3] manages to give lower bounds, exponentially close to $1$, on the probability that $\operatorname{def}(\varphi_{2})=0$. One might wonder how this generalizes. Let us review some arguments that support such a generalization, but also impose some limitations on its validity.
Reformulated in our setting, the minimal resolution conjecture of [18] postulates that, over an infinite field, a Zariski-generic code has $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$, or equivalently, that $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$. However, two points require our attention:
  1. 1.
    This conjecture is known to be false in general [10].
  2. 2.
    We work over a finite field, not an infinite one.
Concerning point 1, we will argue that the conjecture is still “true enough” for us. First, a nonzero defect might not be a problem in our Betti number estimates, as long as it remains small. Moreover, as noted in the introduction of [10], the conjecture has been proved for a large range of values of $n$ and $k$. In fact, although [10] provides an infinity of counterexamples, these remain limited to specific parameters, namely of the form $n=k+O(\sqrt{k})$. And indeed, perhaps the most valuable result for us is [16], which proves that the conjecture is true when $n$ is large enough with respect to $k$.
Concerning point 2, it is true that codes behave quite differently over a finite field and over an infinite field. For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not. This makes it desirable to investigate links between Betti numbers and distance properties of a code.
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&880&1353&990&-&-&-&-&-&-\\ 2&-&-&-&-&-&330&1617&1870&1221&485&110&11\\ \end{array}$
Figure 6: an idealized $[23,12]$-code according to the minimal resolution conjecture
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&881&1371&1122&315&63&6&-&-&-\\ 2&-&-&1&18&132&645&1680&1876&1221&485&110&11\\ \end{array}$
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1392&1181&300&49&4&-&-&-\\ 2&-&-&4&39&191&630&1666&1874&1221&485&110&11\\ \end{array}$
Figure 7: some actual $[23,12]_{2}$-codes with $d=4$, $d^{\perp}=4$ (compare also with Figure 3)
For any code ${\mathsf{C}}$, let us denote by $A_{i}({\mathsf{C}})$ the number of codewords of Hamming weight $i$ in ${\mathsf{C}}$.
Experimental fact 1
Let ${\mathsf{C}}$ be a $[n,k]_{q}$-code of regularity $2$, with minimum distance $d=\operatorname{d}_{\min}({\mathsf{C}})$ and dual minimum distance $d^{\perp}=\operatorname{d}_{\min}({\mathsf{C}}^{\perp})$.
  1. 1.
    For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})=A_{d^{\perp}}({\mathsf{C}}^{\perp})$.
  2. 2.
    Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{k-d,k+1-d}({\mathsf{C}})=A_{d}({\mathsf{C}})$.
It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and for $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$, when applicable (i.e. when these intervals are nonempty).
Conversely, for a random ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$. In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
As an illustration, Figure 13 in the Supplementary material presents statistics on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes.
Consequently, we postulate:
Heuristic 1
Fix a field cardinality $q$, assume $n$ is not too close to $k$ in order to stay away from the counterexamples to the minimal resolution conjecture, and $n<\binom{k+1}{2}$ in order to ensure regularity $2$. Then for random $[n,k]_{q}$-codes, with high probability:
  1. 1.
    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$
  2. 2.
    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$.
Remark 2
Consider this Heuristic in the asymptotic regime. Setting $R=k/n$, we can take $d=\operatorname{d}_{GV}(q,n,k)\approx H_{q}^{-1}(1-R)n$ and $d^{\perp}=\operatorname{d}_{GV}(q,n,n-k)\approx H_{q}^{-1}(R)n$ the corresponding Gilbert-Varshamov distance. Then the condition in 1. translates as $H_{q}^{-1}(R)>R^{2}$, and the condition in 2. translates as $H_{q}^{-1}(1-R)>R(1-R)$, both of which are satisfied when $R$ is small enough. In particular for $q=2$, we find that 1. is satisfied for $R<0.141$ and 2. is satisfied for $R<0.277$.
5 The distinguisher, with and without shortening
General principles
Distinguishers for alternant or Goppa codes tend to work better when $t$ is small and $n$ is large. When benchmarking, it is thus common to:
  • first fix $q,m$, and then for $n=q^{m}$, find the largest distinguishable $t$
  • once such a $t$ is found, fix $q,m,t$, and find the smallest distinguishable $n$.
Also, distinguishers typically work by computing certain code invariants. We might have theoretical bounds on the values of these invariants, that are essential for an asymptotic analysis. However these bounds need not be tight. Hence for a given set of finite parameters, we can also adopt a more empirical approach: sample a certain number of codes, compute their invariants, and observe when the distinguisher “just works”.
So for given $q,m,t$ and a type of codes (alternant or Goppa), we set $n=q^{m}$ and we compute the Betti numbers of a certain number of dual codes. Most of the time, it turns out that these numbers are the same for all samples. We will denote by $\beta_{r-1,r}^{*}$ these common values. By Corollary 3, these $\beta_{r-1,r}^{*}$ still provide lower bounds on $\beta_{r-1,r}$ for smaller $n$.
On the other hand, for random codes of dimension $k=mt$, if Heuristic 1 applies, (86) gives that $\beta_{r-1,r}$ should be equal, or very close, to $\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$. Thus we expect to distinguish when this value is smaller than $\beta_{r-1,r}^{*}$, or equivalently when
$n\geq\left\lceil\frac{k(k+1)}{r}-\frac{\beta_{r-1,r}^{*}-1}{\binom{k-1}{r-2}}\right\rceil$ (88)
(provided $\beta_{r-1,r}^{*}>0$).
When $r=2$, this gives the distinguishability threshold of the square code distinguisher of [11]. Using syzygies of higher degree $r$ allows to make the term $\frac{k(k+1)}{r}$ smaller and reach a broader range of parameters.
As for shortening, we do not expect it to improve the distinguishability threshold. This technique is used only, but crucially, to reduce the overall complexity when parameters grow larger.
Now let us see how our syzygy distinguisher fares, first on the examples proposed in [6], and then on the Classic McEliece system.
Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{97}$, while the distinguisher of [6] works down to $n_{\text{CMT}}=\mathbf{80}$.
We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, which is better than the lower bound in Theorem 3.1 ($e=0$, $r_{\max}\geq t-1=3$). Moreover, their Betti numbers are:
$\beta_{1,2}^{*}=40,\qquad\beta_{2,3}^{*}=80,\qquad\beta_{3,4}^{*}=12.$ (89)
We then observe that the $\beta_{2,3}$-distinguisher works down to $n_{\beta_{2,3}}=\mathbf{86}$, and the $\beta_{3,4}$-distinguisher works down to $n_{\beta_{3,4}}=\mathbf{68}$, both of which coincide with (88). More precisely, computing the Betti numbers of dual Goppa and random codes around these values of $n$ consistently yields:
$\begin{array}[]{|c|ccccccccccccc|}\hline\cr n&\cdots&88&87&\mathbf{86}&85&84&\cdots&70&69&\mathbf{68}&67&66&\cdots\\ \hline\cr\beta_{2,3}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&80&80&\mathbf{80}&85&100&&310&325&340&355&370&\\ \beta_{2,3}^{\text{random}}&&40&55&\mathbf{70}&85&100&&310&325&340&355&370&\\ \hline\cr\beta_{3,4}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&12&12&12&12&12&&12&12&\mathbf{12}&105&210&\\ \beta_{3,4}^{\text{random}}&&0&0&0&0&0&&0&0&\mathbf{0}&105&210&\\ \hline\cr\end{array}$
We see that they stick to their “predicted” values: $\beta_{r-1,r}=\max(\beta_{r-1,r}^{*},\operatorname{ind}(\varphi_{r})^{+})$ for dual Goppa, and $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})^{+}$ for random codes.
Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{62}$, while the distinguisher of [6] works down to $n_{\text{CMT}}=\mathbf{59}$.
We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, which is better than the lower bound in Corollary 4 ($\widehat{e}=1$, $r_{\max}\geq 7$). Moreover, their top Betti numbers are:
$\beta_{5,6}^{*}=1020,\qquad\beta_{6,7}^{*}=288,\qquad\beta_{7,8}^{*}=42.$ (90)
From (88) we expect to distinguish at $\beta_{5,6}$ for $n\geq 57$, at $\beta_{6,7}$ for $n\geq 49$, and at $\beta_{7,8}$ for $n\geq 43$.
And indeed at $n_{\beta_{5,6}}=\mathbf{57}$ we consistently find $\beta_{5,6}\geq 1020$ for dual Goppa codes, while $\beta_{5,6}<500$ for random codes with quite high probability.
For smaller $n$ we have to pass to $\beta_{6,7}$. The $\beta_{6,7}$-distinguisher works well for $n=56$, but the quality gradually falls (distinguishing errors occur more frequently) as $n$ becomes smaller. It is difficult to point a precise threshold where the distinguisher ceases to work. Arguably we still have a positive advantage at $n_{\beta_{6,7}}=\mathbf{50}$, but not anymore at $n=49$. We could then try with $\beta_{7,8}$, but this fails too.
What happens? It turns out the conditions in Heuristic 1 are not satisfied anymore to ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$. From Experimental fact 1, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$. For $r=7$ and $k=mt=18$ this gives $d\geq 13$. For smaller $d$ we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$. The distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ w.h.p., and experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably. Now, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
The Classic McEliece 348864 system
We consider a dual code ${\mathsf{C}}\in\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$ proper of dimension $k=mt=768$. Theorem 3.1 gives $e=5$, $r_{\max}({\mathsf{C}})\geq 321$, which Corollary 4 improves to $\widehat{e}=3$, $r_{\max}({\mathsf{C}})\geq 427$.
For $n=3488$ and $k=768$ we have $\left\lfloor\frac{k(k+1)}{n}\right\rfloor=169$. Moreover, $d=\operatorname{d}_{GV}(2,n,k)=810$, so $k+1-d<0$. By Heuristic 1 we expect to distinguish at $\beta_{r-1,r}$ with $r=170$, with a large margin of safety. For this, Algorithm 1 computes the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $k\beta_{i-2,i-1}\times\binom{k+1}{2}\beta_{i-3,i-2}$. Using $\operatorname{ind}(\varphi_{i})=\left(\frac{k(k+1)}{i}-n\right)\binom{k-1}{i-2}$ as a lower bound for these $\beta_{i-1,i}$, we find that the maximal dimension of these matrices is at least $2^{596}$. Using $\omega\approx 2.372$ for the exponent of linear algebra gives a complexity at least $2^{1414}$.
However we have plenty of space in the interval $170\leq r\leq 427$ to apply shortening. Set $s=377$, $n_{s}=n-s=3111$, $k_{s}=k-s=391$. If ${\mathsf{C}}_{s}$ is a $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$, Corollary 4 ensures $r_{\max}({\mathsf{C}}_{s})\geq 427-s=50$. On the other hand, we have $\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor=49$, and $d_{s}=\operatorname{d}_{GV}(2,n_{s},k_{s})=921$, so $k_{s}+1-d_{s}\ll 0$. Thus we expect to distinguish at $\beta_{r_{s}-1,r_{s}}$ with $r_{s}=50$. Now, for $i\leq r_{s}$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}$. As $d_{s}^{\perp}=\operatorname{d}_{GV}(2,n_{s},n_{s}-k_{s})=55>r_{s}$, these $\beta_{i-1,i}$ can be estimated as $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$. We then find that the maximal dimension of these matrices is around $2^{223}$, and the complexity around $2^{529}$.
Asymptotics
Fix a base field cardinality $q$, for instance $q=2$, and a (dual) rate $R$. In [4] it is suggested to work with a primal code of rate between $0.7$ and $0.8$, so passing to the dual gives $0.2\leq R\leq 0.3$. For $n\to\infty$ we set:
  • $m=\lceil\log_{q}(n)\rceil=\log_{q}(n)+O(1)$
  • $k\approx Rn\;$ such that:
  • $t=\frac{k}{m}\;$ is an integer.
Theorem 5.1
Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}$, or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes. Assume we have a lower bound of the form
$r_{\max}({\mathsf{C}}_{s})\geq k-s^{*}-s$ (91)
with $s^{*}=s^{*}(\mathcal{T}_{q,m,n,t})=o(n)$, valid for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of ${\mathsf{C}}$. Let then $r=o(n)$ be an integer such that $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$. Now let ${\mathsf{C}}$ be a $[n,k]_{q}$-code. Then computing
$\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r}),$ (92)
where ${\mathsf{C}}_{k-s^{*}-r}$ is a $(k-s^{*}-r)$-shortening of ${\mathsf{C}}$, allows to distinguish with high probability whether ${\mathsf{C}}$ is a random code ($\beta_{r-1,r}=0$) or an element of $\mathcal{T}_{q,m,n,t}$ ($\beta_{r-1,r}>0$), by doing at most $r$ kernel computations of linear maps between vector spaces whose maximal dimension can be estimated as $(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$.
In fact it is possible to take $s^{*}\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ and $r\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n$, which gives for the total complexity of the process an estimated upper bound of
$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}.$ (93)
Observe [5, §3.4] that the security parameter in the Classic McEliece system, based on the complexity of generic decoding algorithms, is linear in $t\propto n/\log(n)$. Then in (93) we have $\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$, which means that our complexity is subexponential in the security parameter.
Proof
The shortened code ${\mathsf{C}}_{k-s^{*}-r}$ has dimension $s^{*}+r$ and length $n-k+s^{*}+r$. As $s^{*}$ and $r$ are $o(n)$ we have $s^{*}+r\ll n-k+s^{*}+r$, and $\operatorname{d}_{GV}$ is close to the length. Moreover $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$ by hypothesis. Thus Heuristic 1 applies, and we expect $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})=0$ with high probabilty when ${\mathsf{C}}$ is random. On the other hand, (91) gives $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})>0$ for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$. Using Algorithm 1 we obtain $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $(s^{*}+r)\beta_{i-2,i-1}({\mathsf{C}}_{k-s^{*}-r})\times\binom{s^{*}+r+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{k-s^{*}-r})$. As the rate $\frac{s^{*}+r}{n-k+s^{*}+r}$ goes to zero, by Heuristic 1 again we can estimate the largest dimension of this matrix as $\binom{s^{*}+r+1}{2}\operatorname{ind}(\varphi_{i-2,{\mathsf{C}}_{k-s^{*}-r}})\leq(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$, for random codes — and we can assume this holds also for codes in $\mathcal{T}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
Now for a concrete choice of $s^{*}$ and $r$, set
  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$
so
$f=\left(1-\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}+O\left(\frac{1}{\log_{q}(n)}\right)\right)k.$ (94)
By Theorem 3.1, for all $s$ and all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ we have $r_{\max}({\mathsf{C}}_{s})\geq f-s$, so we can take $s^{*}=k-f\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ as claimed. (In the binary Goppa case we could use Corollary 4 instead of Theorem 3.1, but this leads to the same estimate on $f$, hence on $s^{*}$.) Then for $\varepsilon>0$, setting $r=\left\lceil(1+\varepsilon)\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\right\rceil\ll s^{*}$, we have $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}\approx\frac{(s^{*})^{2}}{n-k}\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\ll r$, and then we can let $\varepsilon\to 0$.
Last, for the complexity estimate (93), using $r\ll s^{*}$ again and Stirling’s formula, we get $\log_{q}\left((s^{*}+r)^{3}\binom{s^{*}+r}{r-3}\right)\approx r\log_{q}\left(\frac{s^{*}}{r}\right)$, and we conclude.
We observe it is possible to slightly relax the condition $m=\lceil\log_{q}(n)\rceil$: complexity remains subexponential as long as $\frac{m}{\log_{q}(n)}\to 1$ fast enough.
6 Conclusion and open problems
We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms, more precisely, subexponential in the error-correcting capability of the code. However this is only an asymptotic result. For concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the security parameter by a non-negligible factor.
Problem 1
To what extent can the implementation of the distinguisher be improved? For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be:
  • somewhat sparse
  • already partially reduced.
This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster. More generally, optimizations used in related domains, for instance in Gröbner basis algorithms, could probably be imported.
(Interestingly, the author was led to introduce the shortening technique in this distinguisher by serendipity, after witnessing a strange phenomenon while partially implementing the optimization alluded in Problem 1.)
Problem 2
Can our distinguisher, and in particular the Betti number computation, benefit from a quantum speedup?
Problem 3
Our lower bounds on the Betti numbers, and in particular the $r_{\max}$, of dual alternant or Goppa codes, are not tight in general. Can they be improved? Better, can one give a complete, explicit description of the minimal resolution of these codes? (See the Supplementary material for elements in this direction.) How would this impact the complexity of the distinguisher?
We initiated the study of syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
Problem 4
Pursue this study in a systematic way. This should cover basic operations on codes that have a standard geometric interpretation, such as puncturing (discarding one subset from a set of points), shortening (projection from one or several of the points), duality of codes (“Gale duality”), tensor product of codes (Segre embedding), and powers of codes (Veronese embedding), but also other operations such as the $(u|u+v)$, subfield subcode, trace code, or concatenated code constructions; and crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
Problem 5
Can syzygies serve in decoding algorithms?
Problem 6
Can our syzygy distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack? For instance, one would like to combine our results with:
  • filtration arguments
  • extraction of short relations, e.g. via MinRank techniques.
(Observe that some of the explicit syzygies described in Proposition 3 are short, while Proposition 4 shows that short relations pass to shortened subcodes.)
It is very common that new mathematical tools (Euclidean lattices, elliptic curves, pairings, isogenies…) are introduced in the cryptographic realm first for cryptanalytic purposes, i.e. to break systems. Once digested by the community, they are then used in a more constructive way, to build new cryptosystems.
Problem 7
Can one devise syzygy-based cryptography?
Supplementary material
More on syzygies of shortened dual alternant and Goppa codes
It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$ in Theorem 3.1, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$.
Let us first consider the simpler case $s=0$.
The Eagon-Northcott complex of each such conjugate gives a subcomplex of the minimal resolution of ${\mathsf{C}}$, however one should take care of the fact that in general these subcomplexes are not in direct sum. Proposition 3 gives a basis $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ for each syzygy space $M_{r-1,r}(\mathbf{\Phi}^{(u)})$. The indices $i_{1},\dots,i_{r}$ correspond to columns of $\mathbf{\Phi}^{(u)}=\left(\begin{array}[]{c|c|c|c}\mathbf{B}^{(u+e)}_{0}&\mathbf{B}^{(u+e-1)}_{1}&\cdots&\mathbf{B}^{(u)}_{e}\end{array}\right)$. It is then easily seen:
Proposition 11
The $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ with column $i_{1}$ not in $\mathbf{B}^{(e+u)}_{0}$ also belong to $M_{r-1,r}(\mathbf{\Phi}^{(u-1)})$. Conversely, for $e<m/2$, those with $i_{1}$ in $\mathbf{B}^{(e+u)}_{0}$ are linearly independent, hence form a basis of the subspace $V_{r}=M_{r-1,r}(\mathbf{\Phi}^{(0)})+\cdots+M_{r-1,r}(\mathbf{\Phi}^{(r)})$ of $M_{r-1,r}({\mathsf{C}})$.
This gives the improved lower bound:
Corollary 6
Set $e=\lfloor\log_{q}(t-1)\rfloor$ and $f=(e+1)t-\frac{q^{e+1}-1}{q-1}$. Then for any $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{r})=m(r-1)\left(\binom{f}{r}-\binom{f-(t-1)}{r}\right).$ (95)
For $r=2$ this specializes to
$\beta_{1,2}({\mathsf{C}})\;\geq\;\dim(V_{2})=\frac{m(t-1)}{2}\left((2e+1)t-2\frac{q^{e+1}-1}{q-1}\right),$ (96)
a result already proved in [11] by (essentially) the same method. It is observed moreover than in many cases the lower bound (96) is an equality, so we actually have
$I_{2}({\mathsf{C}})=V_{2}.$ (97)
Likewise for $r>f-(t-1)$ this gives
$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{f})=m(r-1)\binom{f}{r}$ (98)
and we will see examples where this is an equality.
However for arbitrary $r$, the improved lower bound (95) still isn’t tight, because the inclusion $V_{r}\subseteq M_{r-1,r}({\mathsf{C}})$ is strict in general. For instance, one can show that $M_{2,3}({\mathsf{C}})$ contains syzygies of the form
$\begin{split}X_{a}(&X_{b+1}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}+1}^{(u)}X_{c}^{(u+v)})-X_{a+q^{u}}(X_{b}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}}^{(u)}X_{c}^{(u+v)})+\\ &+X_{c}^{(u+v)}(X_{a}X_{b+q^{v}+1}^{(u)}-X_{a+q^{u}}X_{b+q^{v}}^{(u)})-X_{c+1}^{(u+v)}(X_{a}X_{b+1}^{(u)}-X_{a+q^{u}}X_{b}^{(u)})=0\end{split}$ (99)
that do not belong to $V_{3}$ in general.
In principle, under (97) it should be possible to use the explicit description of $V_{2}$ together with Proposition 2 to compute (the linear strand of) the minimal resolution of ${\mathsf{C}}$. However the combinatorics appears quite complicated.
Now for general $s$, (68) and (73) can be extended to:
Experimental fact 2
Let ${\mathsf{C}}$ be a dual alternant code. Then for all $s$ we have
$r_{\max}({\mathsf{C}}_{s})\geq r_{\max}({\mathsf{C}})-s.$ (100)
It would be tempting to conjecture that (100) holds for all codes, but it turns out that one can find counterexamples. However, these counterexamples are quite rare. So maybe an interesting problem instead should be to give criteria for (100) to hold.
Observe that shortening is dual to puncturing. As the behaviour of minimal resolutions of codes under puncturing is quite well understood (Corollary 3), maybe the study of shortening should go hand-in-hand with that of code duality. Experiments suggest a loose link between $\beta_{r-2,r}({\mathsf{C}}^{\perp})$ and $\beta_{k-r,k+1-r}(C)$, at least for some regimes of parameters.
Back to dual alternant or Goppa codes, we actually observe a strong regular pattern (for Goppa codes the author only tested the irreducible case, but the result is likely to generalize):
Experimental fact 3
Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$ or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes, namely, either dual alternant codes, or dual Goppa codes with irreducible Goppa polynomial. Let $q$ be a field cardinality, and $t\geq 3$ an integer.
  1. 1.
    For all $m$ large enough,
    $r_{\max}({\mathsf{C}})=r^{*}_{\mathcal{T},q,t}$ (101)
    is the same for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,q^{m},t}$, i.e. it generically does not depend on $m$ nor on the choice of ${\mathsf{C}}$, but only on $\mathcal{T},q,t$.
Now $\mathcal{T},q,t$ being fixed, we set $r^{*}=r^{*}_{\mathcal{T},q,t}$. Also, given $m,t$, we set $k=mt$.
  1. 2.
    For all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, and for all $s\leq r^{*}-2$, if $r^{*}-s>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
    $r_{\max}({\mathsf{C}}_{s})=r^{*}-s.$ (102)
  2. 3.
    For $20\leq i\leq r^{*}-$ 2 there are functions $b_{i}(r)$ (actually depending on $\mathcal{T},q,t$, so $b_{i}(r)=b_{i,\mathcal{T},q,t}(r)$) such that, for all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, for all $s\leq r^{*}-2$, and for all $r$ in the interval $2\leq r\leq r^{*}-s$, if $r>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,b_{r^{*}-s-r}(r).$ (103)
  3. 4.
    We have $b_{0}(r)=r-1$ independently of $\mathcal{T},q,t$, so (103) reduces to
    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=m\,(r^{*}-s-1).$ (104)
    More generally for $i\geq 1$ we have $b_{i}(r)=(r-1)\binom{r+i}{i}$ for a certain number of (but now, not for all) values of $\mathcal{T},q,t$, so then (103) reduces to
    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,(r-1)\binom{r^{*}-s}{r}.$ (105)
We observe that (68) gives a lower bound on $r^{*}$. For alternant codes, experimentally, this lower bound is an equality. However for Goppa codes this lower bound seems to be always a strict inequality. And even for binary Goppa codes, the improved lower bound (73) still is a strict inequality.
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&251&1400&3230&2480&\mathbf{1400}&\mathbf{480}&\mathbf{70}\\ 1&202&880&1170&\mathbf{840}&\mathbf{350}&\mathbf{60}&-\\ 2&154&440&\mathbf{450}&\mathbf{240}&\mathbf{50}&-&-\\ 3&107&\mathbf{200}&\mathbf{150}&\mathbf{40}&-&-&-\\ 4&66&\mathbf{80}&\mathbf{30}&-&-&-&-\\ 5&31&\mathbf{20}&-&-&-&-&-\\ 6&\mathbf{10}&-&-&-&-&-&-\\ 7&-&-&-&-&-&-&-\end{array}$
Figure 8: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{2,10,5}\;$ ($r^{*}=8$)
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ 8&18&-&-&-&-&-&-\end{array}$
Figure 9: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{3,5,6}\;$ ($r^{*}=8$)
$\begin{array}[]{c|ccccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}&\beta_{8,9}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 8&280&3224&7464&4272&\mathbf{3360}&\mathbf{1728}&\mathbf{504}&\mathbf{64}\\ 9&249&2510&1800&\mathbf{1792}&\mathbf{1120}&\mathbf{384}&\mathbf{56}&-\\ 10&219&1856&\mathbf{840}&\mathbf{672}&\mathbf{280}&\mathbf{48}&-&-\\ 11&190&1260&\mathbf{360}&\mathbf{192}&\mathbf{40}&-&-&-\\ 12&162&720&\mathbf{120}&\mathbf{32}&-&-&-&-\\ 13&135&234&\mathbf{24}&-&-&-&-&-\\ 14&109&\mathbf{16}&-&-&-&-&-&-\end{array}$
Figure 10: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,8,5}\;$ ($r^{*}=17$)
$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr 0&180&1293&3090&3144&960&\mathbf{36}\\ 1&151&810&1350&732&\mathbf{30}&-\\ 2&123&450&450&\mathbf{24}&-&-\\ 3&96&210&24&-&-&-\\ 4&70&30&-&-&-&-\\ 5&45&-&-&-&-&-\\ 6&21&-&-&-&-&-\\ 7&-&-&-&-&-&-\end{array}$
Figure 11: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{3,6,5}\;$ ($r^{*}=7$)
$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 32&719&8474&\mathbf{450}&\mathbf{240}&\mathbf{50}&-\\ 33&662&6216&\mathbf{150}&\mathbf{40}&-&-\\ 34&606&4070&\mathbf{30}&-&-&-\\ 35&551&2034&-&-&-&-\\ \end{array}$
Figure 12: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}\;$ ($r^{*}=38$)
These Figures illustrate the last Experimental fact. Boldface values match (105). This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
Experimental data on defects for random codes
Figure 13 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes. For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling). The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
Here we have $\frac{k(k+1)}{n}\approx 4.86$. Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable), while $\operatorname{def}(\varphi_{r})=0$ with high probability as we move away from these intervals.
$d=11$
$d=12$
$d=13$
$d=14$
$d^{\perp}\!=\!3$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.269&[1,3]\\ 4&23.821&[15,55]\\ 5&6.927&[5,21]\\ 6&1.341&[1,7]\\ 7&0.042&[0,1]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.245&[1,3]\\ 4&23.171&[15,52]\\ 5&1.948&[1,8]\\ 6&0.086&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.201&[1,3]\\ 4&21.975&[15,48]\\ 5&0.345&[0,5]\\ 6&0.006&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.164&[1,3]\\ 4&20.902&[14,47]\\ 5&0.067&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
$d^{\perp}\!=\!4$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&6.178&[1,14]\\ 5&6.514&[5,20]\\ 6&1.263&[1,7]\\ 7&0.035&[0,1]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.963&[1,14]\\ 5&1.882&[1,8]\\ 6&0.090&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.525&[1,12]\\ 5&0.357&[0,5]\\ 6&0.010&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&4.885&[1,11]\\ 5&0.053&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
$d^{\perp}\!=\!5$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&5.847&[5,15]\\ 6&1.153&[1,6]\\ 7&0.020&[0,1]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&1.485&[1,6]\\ 6&0.055&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.197&[0,2]\\ 6&0.002&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.033&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
Figure 13: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes

### v5

@@
-The syzygy distinguisher
+& Télécom Paris, C2
 Abstract
@@
 The invariants that allow us to distinguish are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code.
-Since its introduction in 1978, this is the first time an analysis of the McEliece cryptosystem breaks the exponential barrier.
+Since its introduction in 1978, this is the first time an analysis111 we exclude results such as [28] that use an attack model for which direct countermeasures exist of the McEliece cryptosystem breaks the exponential barrier.
 1 Introduction
@@
 However, to an attacker, the public code looks like a random code, and removing the noise is untractable.
-Slightly more formally, a security proof for the system relies on two assumptions:
+Slightly more formally, a traditional security proof for the system relies on two assumptions:
   1. 1.
@@
 Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order!
-This gives a quite convincing feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
+This might give a feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
 Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e.
 a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms.
-Moreover, for given finite parameters, it is more efficient than previous distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
+Moreover, for given finite parameters, it is more efficient than state-of-the-art distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
 Principles and organization
@@
 If this quantity is smaller than the said lower bound, we can distinguish.
-Likewise we claim that Theorem 2.8 of [9] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes, where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher.
+Likewise we claim that Theorem 2.8 of [10] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes,222erratum: among MDS $[7,4]$-codes (which then implies of full square) where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher.
 For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$.
@@
   • On the other hand, we estimate the Betti numbers of random codes in terms of raw code parameters (length, dimension, distance).
-This is done in section 4, partially relying on heuristics for which we do not have full proofs, but still providing experimental evidence, and sometimes also partial theoretical arguments.
+This is done in section 4, partially relying on a natural heuristic: random codes are not expected to admit more syzygies than those forced by these parameters.
+We do not have full proofs for this fact, but we provide experimental evidence and partial theoretical arguments that support it.
 Prior to that, section 2 explains how these invariants can be effectively computed.
@@
 This leads to different theories, although seeking links between the two could be an interesting project.
-Acknowledgements
-The author learned a lot from the code-based cryptography seminar organized by Jean-Pierre Tillich at Inria Paris.
-More important perhaps, without his friendly pressure, probably this work would have stayed eternally in limbo.
-He should be thanked for all that.
-The very favourable environment towards research at ANSSI in general, and at LCR in particular, was also highly helpful.
-Henri Gilbert’s advices led to improvements in several aspects of the presentation.
-Anonymous members of LCR unsuccessfully suggested the title “The Dyztynguysher” for this work.
-The author is part of the ANR-21-CE39-0009 BARRACUDA and ANR-21-CE39-0008 KLEPTOMANIAC research projects. Some of the computations were done on the lame servers in Télécom Paris.
 Notation and conventions
@@
 Restricting to homogeneous polynomials of degree $r$, we have a surjective map $S_{r}\longrightarrow{\mathsf{C}}^{\langle r\rangle}$, whose kernel we denote $I_{r}({\mathsf{C}})$.
-We then define the homogeneous coordinate ring of ${\mathsf{C}}$:
+We then define the homogeneous coordinate ring of ${\mathsf{C}}$ as the formal direct sum
 ${\mathsf{C}}^{\langle\cdot\rangle}=\bigoplus_{r\geq 0}{\mathsf{C}}^{\langle r\rangle}$ (12)
-and its homogeneous ideal:
+and likewise its homogeneous ideal
 $I({\mathsf{C}})=\bigoplus_{r\geq 0}I_{r}({\mathsf{C}}).$ (13)
@@
 If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$.
+Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $[g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
 Iterating this construction, we obtain a minimal resolution
@@
   • $M_{i+1}=\ker(\;F_{i}\longrightarrow M_{i}\;)$.
-Thus the $F_{i}$ are of the form
+The number $\beta_{i,j}$ of degree $j$ elements in a minimal system of generators of $M_{i}$ does not depend on the choices made, and is called the $(i,j)$-th graded Betti number of $M_{0}$.
+Keeping track of the grading, we have
 $F_{i}=\bigoplus_{j\geq 0}S(-j)^{\beta_{i,j}}$ (24)
-where $S(-j)$ is the free $S$-module of rank $1$ generated in degree $j$, so $S(-j)_{d}=S_{d-j}$; and $\beta_{i,j}$, the $(i,j)$-th graded Betti number of $M_{0}$, is the number of elements of degree $j$ in a minimal system of generators of $M_{i}$.
-From now on let ${\mathsf{C}}$ be a $[n,k]$-code, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring.
+where $S(-j)$ is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
+It is customary to display the graded Betti numbers in the form of a Betti diagram, as follows:
+$\begin{array}[]{c|ccccc}&0&1&\dots&i&\dots\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots\\ \vdots&\vdots&\vdots&&\vdots&\\ \end{array}$
+with null entries marked as “--” for readability.
+Lemma 1
+Assume $M_{i}$ is generated in degrees $\geq D$, i.e.
+$\beta_{i,j}=0$ for all $j\leq D-1$.
+Then $M_{i+1}$ is generated in degrees $\geq D+1$, and by induction all the upper-right quadrant of the Betti diagram defined by $\beta_{i,D-1}$ vanishes.
+Proof
+Let $g_{1},\dots,g_{N}$ form a minimal system of generators of $M_{i}$.
+By hypothesis all $g_{u}$ have degree $\geq D$.
+Then, in a homogeneous relation $\sum_{u}f_{u}g_{u}=0$, no $f_{u}$ can be a nonzero constant: otherwise the corresponding $g_{u}$ could be expressed as a linear combination of the others, hence could be removed from the system of generators, contradicting minimality.
+Thus all nonzero $f_{u}$ have degree $\geq 1$, and the element of $M_{i+1}$ defined by this homogeneous relation has degree $\geq D+1$.
+As $M_{i}$ and $F_{i}$ are graded $S$-modules, we will also write $M_{i,j}$ and $F_{i,j}$ for their degree $j$ homogeneous components.
+From now on let ${\mathsf{C}}$ be a $[n,k]$-code, with generator matrix $\mathbf{G}$, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring.
 Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$.
@@
 $0\longrightarrow F_{k-1}\longrightarrow\cdots\longrightarrow F_{2}\longrightarrow F_{1}\longrightarrow F_{0}=S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}\longrightarrow 0$ (25)
-with $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$, $F_{0}=S$, $M_{1}=I({\mathsf{C}})$, hence $\beta_{0,0}({\mathsf{C}})=1$, $\beta_{0,j}({\mathsf{C}})=0$ for $j\neq 0$, and $\beta_{1,j}({\mathsf{C}})$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$.
-It is customary to display the graded Betti numbers in the form of a Betti diagram, as follows:
-$\begin{array}[]{c|cccccc}&0&1&\dots&i&\dots&k-1\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&&\vdots\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots&\beta_{k-1,r+k-1}\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots&\beta_{k-1,r+k}\\ \vdots&\vdots&\vdots&&\vdots&&\vdots\\ \end{array}$
-with null entries marked as “--” for readability. In our situation, as $F_{0}=S$, the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
-By the minimality property of the resolution, if for some $i$ and $D$ we have $\beta_{i,j}=0$ for all $j\leq D$, then also $\beta_{i+1,j}=0$ for all $j\leq D+1$.
-Applied recursively, this means all the upper-right quadrant of the Betti diagram defined by $\beta_{i,D}$ vanishes.
-In particular, as the evaluation map $S_{1}\longrightarrow C$ is an isomorphism, we see that $I({\mathsf{C}})$ is generated in degrees $\geq 2$, which implies:
+and the Betti diagram has $k$ columns, indexed from 00 to $k-1$.
+As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
+Likewise $M_{1}=I({\mathsf{C}})$, and $\beta_{1,j}$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$.
+As the evaluation map $S_{1}\longrightarrow C$ is an isomorphism, we see that $I({\mathsf{C}})$ is generated in degrees $\geq 2$.
+Lemma 1 then implies:
 Lemma 2
@@
 $F_{i,i+1}=M_{i,i+1}=\mathbb{F}^{\beta_{i,i+1}}.$ (26)
-Hence the 00-th row of the Betti diagram is $(1,-,-,\dots)$.
+Hence the 00-th row of the Betti diagram is always $(1,-,-,\dots)$.
 Figures 3-3 offer a few such diagrams for contemplation.
@@
 $\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
-Figure 3: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
+Figure 3: the $[23,12]_{2}$ Golay code
+Observe that replacing $\mathbf{G}$ with $\mathbf{S}\mathbf{G}$, for $\mathbf{S}\in\mathbb{F}^{k\times k}$ an invertible matrix, corresponds to a linear change of coordinates for the variables $X_{1},\dots,X_{k}$.
+This is an automorphism of $S$, hence will not affect the Betti numbers.
+Thus these Betti numbers really are invariants of ${\mathsf{C}}$, independently of the choice of $\mathbf{G}$.
+More generally recall that two $[n,k]$ codes ${\mathsf{C}}_{1},{\mathsf{C}}_{2}$, defined by generator matrices $\mathbf{G}_{1},\mathbf{G}_{2}$, are linearly isometric, or monomially equivalent, if $\mathbf{G}_{1}=\mathbf{S}\mathbf{G}_{2}\mathbf{P}\mathbf{D}$ for $\mathbf{S}\in\mathbb{F}^{k\times k}$ an invertible matrix, and $\mathbf{P},\mathbf{D}\in\mathbb{F}^{n\times n}$ a permutation matrix and an invertible diagonal matrix.
+Lemma 3
+Monomially equivalent codes have the same Betti numbers.
+Proof
+We already discussed the action of $\mathbf{S}$. As for the action of $\mathbf{P}\mathbf{D}$, it preserves the kernel of the graded evaluation map $S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$.
 Effective computation
@@
 Let $\mathcal{G}$ be a minimal system of homogeneous generators of $M_{r-2}$, and for each $j$ let $B_{r-2,j}\subseteq M_{r-2,j}$ be the linear subspace generated by the degree $j$ elements of $\mathcal{G}$.
-Then, as $M_{r-2}$ is generated in degrees $\geq r-1$, we have $B_{r-2,r-1}=M_{r-2,r-1}$, and $B_{r-2,r}$ is a supplementary subspace to $S_{1}\cdot B_{r-2,r-1}=\operatorname{im}(\varphi_{r})$ in $M_{r-2,r}$.
+Then, as $M_{r-2}$ is generated in degrees $\geq r-1$, we have $B_{r-2,r-1}=M_{r-2,r-1}$, and $B_{r-2,r}$ is a complementary subspace to $S_{1}\cdot B_{r-2,r-1}=\operatorname{im}(\varphi_{r})$ in $M_{r-2,r}$.
 This proves (29).
@@
 The $M_{r-1,r}({\mathsf{C}})$ only depend on $I_{2}({\mathsf{C}})$.
-To make things even more explicit we introduce the following matrices.
+To make things entirely explicit we now introduce the following matrices.
 Assume ${\mathsf{C}}$ is given by a generator matrix $\mathbf{G}\in\mathbb{F}^{k\times n}$, with rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$.
@@
 The squared matrix of $\mathbf{G}$ is the matrix
-$\mathbf{G}_{2}\in\mathbb{F}^{\,\binom{k+1}{2}\,\times\,n}$ (33)
+$\mathbf{M}_{2}\in\mathbb{F}^{\,\binom{k+1}{2}\,\times\,n}$ (33)
 with rows indexed by $\mathcal{M}_{2}$: the row corresponding to $X_{a}X_{b}$ is $\mathbf{c}_{a}\mathbf{c}_{b}$.
-Then $I_{2}({\mathsf{C}})$ is the left kernel of $\mathbf{G}_{2}$, and we choose a basis $\mathcal{B}_{2}$ of this space.
+Then $I_{2}({\mathsf{C}})$ is the left kernel of $\mathbf{M}_{2}$, and we choose a basis $\mathcal{B}_{2}$ of this space.
 Thus $\mathcal{B}_{2}$ consists of $\beta_{1,2}$ vectors, each of which has its entries indexed by $\mathcal{M}_{2}$.
@@
 with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 00 elsewhere.
+(Said otherwise, the rows of $\mathbf{M}_{3}$ are formed by multiplying all $\mathbf{q}\in\mathcal{B}_{2}$, seen as quadratic forms $\mathbf{q}=\sum_{1\leq b\leq c\leq k}q_{X_{b}X_{c}}X_{b}X_{c}$, by all variables $X_{a}$.)
 Then by (30), $M_{2,3}$ is the left kernel of $\mathbf{M}_{3}$, and we choose a basis $\mathcal{B}_{3}$ of this space.
@@
 with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 00 elsewhere.
+(Said otherwise, the rows of $\mathbf{M}_{r}$ are formed by multiplying all $\mathbf{s}\in\mathcal{B}_{r-1}$ by all variables $X_{a}$, where each $\mathbf{s}$ is seen as a formal sum $\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}L_{\mathbf{t}}[\mathbf{t}]$ whose coefficients are linear forms $L_{\mathbf{t}}=\sum_{1\leq b\leq k}s_{X_{b},\mathbf{t}}X_{b}$, resulting in formal sums $X_{a}\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}(X_{a}L_{\mathbf{t}})[\mathbf{t}]$ whose coefficients are quadratic forms.)
 Then by (31), $M_{r-1,r}$ is the left kernel of $\mathbf{M}_{r}$, and we proceed.
@@
 Input: a generator matrix $\mathbf{G}$ of the $[n,k]$-code ${\mathsf{C}}$ a target degree $D\leq k$
-1:construct the matrix $\mathbf{G}_{2}$ according to Definition 1
-2:compute a left kernel basis $\mathcal{B}_{2}$ of $\mathbf{G}_{2}$
+1:construct the matrix $\mathbf{M}_{2}$ according to Definition 1
+2:compute a left kernel basis $\mathcal{B}_{2}$ of $\mathbf{M}_{2}$
 3:construct the matrix $\mathbf{M}_{3}(\mathcal{B}_{2})$ according to Definition 2
@@
 Sometimes we only control the syzygies of certain subideals of $I({\mathsf{C}})$, and from these we want to deduce information on the syzygies of the whole ideal.
-Intuitively, we expect syzygies between elements of the subideals could be seen as syzygies between elements of the ideal.
+Intuitively, we expect that syzygies between elements of the subideals could be seen as syzygies between elements of the ideal.
 Of course this fails in general, but it remains true if we restrict to the linear strand:
@@
 $0\to(\operatorname{Sym}^{f-g}G)^{\vee}\otimes\bigwedge^{f}F\to\cdots\to G^{\vee}\otimes\bigwedge^{g+1}F\to\bigwedge^{g}F\overset{\bigwedge^{g}\mathbf{\Phi}}{\longrightarrow}\bigwedge^{g}G\simeq R.$ (46)
-Under mild hypotheses, this complex is exact [9, Th.
+Under mild hypotheses, this complex is exact [10, Th.
 A.2.60 & Th.
@@
 $s^{(j)}_{r;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z^{(j)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z^{(j-1)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
-where we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
+where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
 $\mathbf{s}_{r}(\mathbf{s}_{r-1})=\mathbf{0}.$ (55)
@@
 elements of $S_{1}$.
-Recall from [9, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero.
+Recall from [10, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero.
 Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
 By Proposition 4 applied $s$ times (with $g=d=2$), $I_{2}({\mathsf{C}}_{s})$ then contains the $2\times 2$ minors of a $2\times f_{s}$ matrix $\mathbf{\Phi}_{s}$, where $f_{s}\geq f-s$, and this matrix $\mathbf{\Phi}_{s}$ is $1$-generic also.
-Then by [9, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$.
+Then by [10, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$.
 We then conclude with Corollary 2 (with $l=1$ and $V^{(1)}=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi}_{s})$).
@@
 It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$.
-See the corresponding discussion in the Supplementary material.
+See the corresponding discussion at the beginning of the Supplementary material.
 Examples show the bound (68) on $r_{\max}$ is tight in general for alternant codes.
@@
 Proof
-Generating series reformulation of [9, Cor. 1.10].
-Definition 5 ([25, Def. 1.5 & Th. 2.35])
+Generating series reformulation of [10, Cor. 1.10].
+Definition 5 ([26, Def. 1.5 & Th. 2.35])
 The Castelnuovo-Mumford regularity of a projective $[n,k]$-code ${\mathsf{C}}$ is the smallest integer $r$ such that ${\mathsf{C}}^{\langle r\rangle}=\mathbb{F}^{n}$.
-Definition 6 (cf. [9, §4A], after [22, Lect. 14])
+Definition 6 (cf. [10, §4A], after [23, Lect. 14])
 The Castelnuovo-Mumford regularity of ${\mathsf{C}}^{\langle\cdot\rangle}$ is $\max\{r:\>\exists i,\,\beta_{i,i+r}({\mathsf{C}})>0\}$.
-Proposition 7 ([9, Th. 4.2])
+Proposition 7 ([10, Th. 4.2])
 These two definitions coincide.
+From now on we will just say “regularity” for short.
 The square code distinguisher, and the filtration attack from [1] that extends it, work for codes ${\mathsf{C}}$ with ${\mathsf{C}}^{\langle 2\rangle}\varsubsetneq\mathbb{F}^{n}$, i.e.
@@
 $\begin{split}\operatorname{ind}(f)&=\dim(U)-\dim(V)\\ &=\dim\ker(f)-\dim\operatorname{coker}(f)\end{split}$ (76)
-and its defect
+and its rank defect
 $\begin{split}\operatorname{def}(f)&=\min(\dim(U),\dim(V))-\operatorname{rk}(f)\\ &=\min(\dim\ker(f),\dim\operatorname{coker}(f)).\end{split}$ (77)
@@
 and moreover, if ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then
-$\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}.$ (83)
+$\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ (83)
 Proof
@@
 The minimal resolution of ${\mathsf{C}}$ has length $k-1$, so $\beta_{i,j}=0$ for $i>k-1$, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
-Regularity $2$ helps in computing the full Betti diagram.
-We already gave examples in Figures 3-3.
-In order to provide a few more, let us first recall that a $[n,k]$ code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
-Lemma 9 ([14, Th. 1], reformulated)
-Let ${\mathsf{C}}$ be a $[n,k]$-code with $n\leq 2k-1$. Assume ${\mathsf{C}}$ is MDS. Then for all $r\leq 2k+1-n$ we have $\beta_{r-2,r}({\mathsf{C}})=0$.
-Proposition 8
-Let ${\mathsf{C}}$ be a $[k+1,k]$ MDS code, for instance a parity code or a $[k+1,k]$ $\operatorname{{\mathsf{GRS}}}$ code.
-Then the nonzero Betti numbers of ${\mathsf{C}}$ are $\beta_{0,0}=\beta_{k-1,k+1}=1$, and
-$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (85)
-for $12\leq r\leq k-$ 1. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
-Proof
-The parameters imply that ${\mathsf{C}}$ has regularity $2$.
-By Lemma 9 we have $\beta_{r-2,r}=0$ for $r\leq k$.
-Then $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-(k+1)\right)\binom{k-1}{r-2}$ and we conclude with a straightforward calculation.
-Proposition 9
-Let ${\mathsf{C}}$ be a $[2k-1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are:
-  • $\beta_{0,0}=1$
-  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
-  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
-In particular the ideal $I({\mathsf{C}})$ is generated by $\beta_{1,2}=\frac{(k-1)(k-2)}{2}$ quadratic forms and $\beta_{1,3}=k-1$ cubic forms.
-Proof
-Using the basis $\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}$ of ${\mathsf{C}}$, we see $I_{2}({\mathsf{C}})$ contains the $2\times 2$ minors of $\mathbf{\Phi}=\left(\begin{array}[]{cccc}X_{0}&X_{1}&\dots&X_{k-2}\\ X_{1}&X_{2}&\dots&X_{k-1}\end{array}\right)$, i.e.
-$I_{2}({\mathsf{C}})\supseteq\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ of dimension $\binom{k-1}{2}$.
-Now set $n=2k-1$.
-Because $n\geq 2k-1$, we have $\dim{\mathsf{C}}^{\langle 2\rangle}=2k-1$, hence $\dim I_{2}({\mathsf{C}})=\binom{k-1}{2}$, so $I_{2}({\mathsf{C}})=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$.
-Corollary 1 and the Eagon-Northcott complex of $\mathbf{\Phi}$ then give $\beta_{r-1,r}=(r-1)\binom{k-1}{r}$.
-Because $n\leq 2k-1$, we have ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$.
-Then $\beta_{r-2,r}=\beta_{r-1,r}-\operatorname{ind}(\varphi_{r})=(r-1)\binom{k-1}{r}-\left(\frac{k(k+1)}{r}-(2k-1)\right)\binom{k-1}{r-2}$, and a straightforward calculation allows to conclude.
-Figures 5 and 5 illustrate Propositions 8 and 9.
-$\begin{array}[]{c|cccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&27&105&189&189&105&27&-\\ 2&-&-&-&-&-&-&-&1\end{array}$
-Figure 4: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
-$\begin{array}[]{c|ccccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&21&70&105&84&35&6&-\\ 2&-&7&42&105&140&105&42&7\end{array}$
-Figure 5: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
-Exercise 1
-Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for $k+1<n<2k-1$.
+Regularity $2$ helps in computing the full Betti diagram of codes. See Propositions 10 and 11 in the Supplementary material for more.
 The small defect heuristic
-A direct application of Definition 7 and Theorem 4.1 gives:
-Proposition 10
-If ${\mathsf{C}}$ is a $[n,k]$-code with ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then for all $r\geq 2$ we have
-$\beta_{r-1,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r})$ (86)
-$\beta_{r-2,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{-}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r}).$ (87)
-It is tempting to see this result as an estimate for the Betti numbers of random codes, with a leading term given by $\operatorname{ind}(\varphi_{r})^{\pm}$ (compare with the bound in Lemma 3), and a secondary term given by $\operatorname{def}(\varphi_{r})$.
-Indeed, it is well known that under a reasonable probability model, a random linear map tends to have a small defect.
-Unfortunately, if ${\mathsf{C}}$ is a random code (say, uniform among codes of given $[n,k]$), it is not easy to control the distribution of the maps $\varphi_{r}$.
-Still, in the special case $r=2$, [3] manages to give lower bounds, exponentially close to $1$, on the probability that $\operatorname{def}(\varphi_{2})=0$.
-One might wonder how this generalizes.
-Let us review some arguments that support such a generalization, but also impose some limitations on its validity.
-Reformulated in our setting, the minimal resolution conjecture of [18] postulates that, over an infinite field, a Zariski-generic code has $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$, or equivalently, that $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$.
+By (78) and Theorem 4.1, for any $[n,k]$-code ${\mathsf{C}}$ of regularity $2$ and for any $r\geq 2$ we have
+$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}.$ (85)
+We might interpret this lower bound as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$.
+Then, when ${\mathsf{C}}$ is random, it is quite natural to expect that no other syzygy will appear, or equivalently that (85) will be an equality, i.e.
+$\operatorname{def}(\varphi_{r})=0$.
+This is compatible with results such as those in [6] and the references within, that give estimates on the distribution of the rank, hence the defect, of random linear maps, under various probability laws.
+Unfortunately, even if the probability distribution of ${\mathsf{C}}$ is nice (say, uniform among codes of given $[n,k]$), it is not easy to control the distribution of $\varphi_{r}$, so we will not be able to give proofs.
+Moreover it could be that the iterative algebraic process in the construction of $\varphi_{r}$ would lead to some unexpected constraints on its rank.
+And indeed, in this section we will identify some parameter ranges for which the defect cannot be small; but conversely, we will also give arguments that support the validity of this small defect heuristic at least under some proper conditions.
+A first argument in support of the heuristic is that it is unconditionally true when $r=2$.
+Indeed, in this special case, [3] manages to give lower bounds, exponentially close to $1$, on the probability that $\operatorname{def}(\varphi_{2})=0$.
+Another argument is the minimal resolution conjecture of [19].
+It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$ or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$.
 However, two points require our attention:
   1. 1.
-    This conjecture is known to be false in general [10].
+    This conjecture is now known to be false in general [11].
   2. 2.
     We work over a finite field, not an infinite one.
-Concerning point 1, we will argue that the conjecture is still “true enough” for us.
+Concerning point 1, we will argue that the conjecture is still “true enough” for our use.
 First, a nonzero defect might not be a problem in our Betti number estimates, as long as it remains small.
 Moreover, as noted in the introduction of [11], the conjecture has been proved for a large range of values of $n$ and $k$.
-In fact, although [10] provides an infinity of counterexamples, these remain limited to specific parameters, namely of the form $n=k+O(\sqrt{k})$.
+In fact, although [11] provides an infinity of counterexamples, these remain limited to very specific parameters, namely of the form $n=k+O(\sqrt{k})$.
 And indeed, perhaps the most valuable result for us is [17], which proves that the conjecture is true when $n$ is large enough with respect to $k$.
-Concerning point 2, it is true that codes behave quite differently over a finite field and over an infinite field.
+Concerning point 2, clearly, codes behave quite differently over a finite field and over an infinite field.
 For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not.
 This makes it desirable to investigate links between Betti numbers and distance properties of a code.
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&880&1353&990&-&-&-&-&-&-\\ 2&-&-&-&-&-&330&1617&1870&1221&485&110&11\\ \end{array}$
-Figure 6: an idealized $[23,12]$-code according to the minimal resolution conjecture
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&881&1371&1122&315&63&6&-&-&-\\ 2&-&-&1&18&132&645&1680&1876&1221&485&110&11\\ \end{array}$
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1392&1181&300&49&4&-&-&-\\ 2&-&-&4&39&191&630&1666&1874&1221&485&110&11\\ \end{array}$
-Figure 7: some actual $[23,12]_{2}$-codes with $d=4$, $d^{\perp}=4$ (compare also with Figure 3)
 For any code ${\mathsf{C}}$, let us denote by $A_{i}({\mathsf{C}})$ the number of codewords of Hamming weight $i$ in ${\mathsf{C}}$.
@@
   1. 1.
-    For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})=A_{d^{\perp}}({\mathsf{C}}^{\perp})$.
+    For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$.
+    Moreover if $d^{\perp}\leq\frac{k(k+1)}{n}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
   2. 2.
-    Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{k-d,k+1-d}({\mathsf{C}})=A_{d}({\mathsf{C}})$.
+    Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$.
+    Moreover if $k+1-d\geq\frac{k(k+1)}{n}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
 It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and for $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$, when applicable (i.e. when these intervals are nonempty).
-Conversely, for a random ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$.
+Experimental fact 2
+Conversely, for a random code ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$.
 In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
-As an illustration, Figure 13 in the Supplementary material presents statistics on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes.
+Figures 13-14 in the Supplementary material illustrate various aspects of these Experimental facts.
 Consequently, we postulate:
@@
   1. 1.
-    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$
+    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$
   2. 2.
-    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$.
+    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$.
 Remark 2
 Consider this Heuristic in the asymptotic regime.
-Setting $R=k/n$, we can take $d=\operatorname{d}_{GV}(q,n,k)\approx H_{q}^{-1}(1-R)n$ and $d^{\perp}=\operatorname{d}_{GV}(q,n,n-k)\approx H_{q}^{-1}(R)n$ the corresponding Gilbert-Varshamov distance.
+Setting $R=k/n$, we can take $d=\operatorname{d}_{GV}(q,n,k)\approx H_{q}^{-1}(1-R)n$ and $d^{\perp}=\operatorname{d}_{GV}(q,n,n-k)\approx H_{q}^{-1}(R)n$ the corresponding Gilbert-Varshamov distances, where $H_{q}$ is the $q$-ary entropy function.
 Then the condition in 1.
-translates as $H_{q}^{-1}(R)>R^{2}$, and the condition in 2.
-translates as $H_{q}^{-1}(1-R)>R(1-R)$, both of which are satisfied when $R$ is small enough.
+translates as $H_{q}^{-1}(1-R)>R(1-R)$, and the condition in 2.
+translates as $H_{q}^{-1}(R)>R^{2}$, both of which are satisfied when $R$ is small enough.
 In particular for $q=2$, we find that 1.
-is satisfied for $R<0.141$ and 2.
-is satisfied for $R<0.277$.
-5 The distinguisher, with and without shortening
-General principles
-Distinguishers for alternant or Goppa codes tend to work better when $t$ is small and $n$ is large. When benchmarking, it is thus common to:
-  • first fix $q,m$, and then for $n=q^{m}$, find the largest distinguishable $t$
-  • once such a $t$ is found, fix $q,m,t$, and find the smallest distinguishable $n$.
-Also, distinguishers typically work by computing certain code invariants.
+is satisfied for $R<0.277$ and 2.
+is satisfied for $R<0.141$.
+5 The distinguisher
+Basic version (without shortening)
+Distinguishers typically work by computing certain code invariants.
 We might have theoretical bounds on the values of these invariants, that are essential for an asymptotic analysis.
@@
 By Corollary 3, these $\beta_{r-1,r}^{*}$ still provide lower bounds on $\beta_{r-1,r}$ for smaller $n$.
-On the other hand, for random codes of dimension $k=mt$, if Heuristic 1 applies, (86) gives that $\beta_{r-1,r}$ should be equal, or very close, to $\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$.
-Thus we expect to distinguish when this value is smaller than $\beta_{r-1,r}^{*}$, or equivalently when
-$n\geq\left\lceil\frac{k(k+1)}{r}-\frac{\beta_{r-1,r}^{*}-1}{\binom{k-1}{r-2}}\right\rceil$ (88)
-(provided $\beta_{r-1,r}^{*}>0$).
+For a given $r$ with $\beta_{r-1,r}^{*}>0$, our distinguisher then works as follows.
+Taking as input a (dual) code ${\mathsf{C}}$ of dimension $k=mt$, compute $\beta_{r-1,r}({\mathsf{C}})$ with Algorithm 1.
+Then if $\beta_{r-1,r}({\mathsf{C}})\geq\beta_{r-1,r}^{*}$ declare ${\mathsf{C}}$ is of the special type in question.
+Otherwise declare ${\mathsf{C}}$ is random.
+Observe that for random codes of dimension $k$, if the conditions in Heuristic 1 are satisfied, we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$.
+Thus the distinguisher succeeds when this value is smaller than $\beta_{r-1,r}^{*}$, or equivalently when
+$n\geq\left\lceil\frac{k(k+1)}{r}-\frac{\beta_{r-1,r}^{*}-1}{\binom{k-1}{r-2}}\right\rceil.$ (86)
 When $r=2$, this gives the distinguishability threshold of the square code distinguisher of [12].
 Using syzygies of higher degree $r$ allows to make the term $\frac{k(k+1)}{r}$ smaller and reach a broader range of parameters.
-As for shortening, we do not expect it to improve the distinguishability threshold. This technique is used only, but crucially, to reduce the overall complexity when parameters grow larger.
-Now let us see how our syzygy distinguisher fares, first on the examples proposed in [6], and then on the Classic McEliece system.
-Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
-For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{97}$, while the distinguisher of [6] works down to $n_{\text{CMT}}=\mathbf{80}$.
-We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, which is better than the lower bound in Theorem 3.1 ($e=0$, $r_{\max}\geq t-1=3$).
-Moreover, their Betti numbers are:
-$\beta_{1,2}^{*}=40,\qquad\beta_{2,3}^{*}=80,\qquad\beta_{3,4}^{*}=12.$ (89)
-We then observe that the $\beta_{2,3}$-distinguisher works down to $n_{\beta_{2,3}}=\mathbf{86}$, and the $\beta_{3,4}$-distinguisher works down to $n_{\beta_{3,4}}=\mathbf{68}$, both of which coincide with (88).
-More precisely, computing the Betti numbers of dual Goppa and random codes around these values of $n$ consistently yields:
-$\begin{array}[]{|c|ccccccccccccc|}\hline\cr n&\cdots&88&87&\mathbf{86}&85&84&\cdots&70&69&\mathbf{68}&67&66&\cdots\\ \hline\cr\beta_{2,3}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&80&80&\mathbf{80}&85&100&&310&325&340&355&370&\\ \beta_{2,3}^{\text{random}}&&40&55&\mathbf{70}&85&100&&310&325&340&355&370&\\ \hline\cr\beta_{3,4}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&12&12&12&12&12&&12&12&\mathbf{12}&105&210&\\ \beta_{3,4}^{\text{random}}&&0&0&0&0&0&&0&0&\mathbf{0}&105&210&\\ \hline\cr\end{array}$
-We see that they stick to their “predicted” values: $\beta_{r-1,r}=\max(\beta_{r-1,r}^{*},\operatorname{ind}(\varphi_{r})^{+})$ for dual Goppa, and $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})^{+}$ for random codes.
-Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
-For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{62}$, while the distinguisher of [6] works down to $n_{\text{CMT}}=\mathbf{59}$.
-We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, which is better than the lower bound in Corollary 4 ($\widehat{e}=1$, $r_{\max}\geq 7$).
-Moreover, their top Betti numbers are:
-$\beta_{5,6}^{*}=1020,\qquad\beta_{6,7}^{*}=288,\qquad\beta_{7,8}^{*}=42.$ (90)
-From (88) we expect to distinguish at $\beta_{5,6}$ for $n\geq 57$, at $\beta_{6,7}$ for $n\geq 49$, and at $\beta_{7,8}$ for $n\geq 43$.
-And indeed at $n_{\beta_{5,6}}=\mathbf{57}$ we consistently find $\beta_{5,6}\geq 1020$ for dual Goppa codes, while $\beta_{5,6}<500$ for random codes with quite high probability.
-For smaller $n$ we have to pass to $\beta_{6,7}$.
-The $\beta_{6,7}$-distinguisher works well for $n=56$, but the quality gradually falls (distinguishing errors occur more frequently) as $n$ becomes smaller.
-It is difficult to point a precise threshold where the distinguisher ceases to work.
-Arguably we still have a positive advantage at $n_{\beta_{6,7}}=\mathbf{50}$, but not anymore at $n=49$.
-We could then try with $\beta_{7,8}$, but this fails too.
-What happens?
-It turns out the conditions in Heuristic 1 are not satisfied anymore to ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$.
-From Experimental fact 1, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$.
-For $r=7$ and $k=mt=18$ this gives $d\geq 13$.
-For smaller $d$ we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$.
-The distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ w.h.p., and experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably.
-Now, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
-The Classic McEliece 348864 system
-We consider a dual code ${\mathsf{C}}\in\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$ proper of dimension $k=mt=768$.
-Theorem 3.1 gives $e=5$, $r_{\max}({\mathsf{C}})\geq 321$, which Corollary 4 improves to $\widehat{e}=3$, $r_{\max}({\mathsf{C}})\geq 427$.
-For $n=3488$ and $k=768$ we have $\left\lfloor\frac{k(k+1)}{n}\right\rfloor=169$.
-Moreover, $d=\operatorname{d}_{GV}(2,n,k)=810$, so $k+1-d<0$.
-By Heuristic 1 we expect to distinguish at $\beta_{r-1,r}$ with $r=170$, with a large margin of safety.
-For this, Algorithm 1 computes the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $k\beta_{i-2,i-1}\times\binom{k+1}{2}\beta_{i-3,i-2}$.
-Using $\operatorname{ind}(\varphi_{i})=\left(\frac{k(k+1)}{i}-n\right)\binom{k-1}{i-2}$ as a lower bound for these $\beta_{i-1,i}$, we find that the maximal dimension of these matrices is at least $2^{596}$.
-Using $\omega\approx 2.372$ for the exponent of linear algebra gives a complexity at least $2^{1414}$.
-However we have plenty of space in the interval $170\leq r\leq 427$ to apply shortening.
-Set $s=377$, $n_{s}=n-s=3111$, $k_{s}=k-s=391$.
-If ${\mathsf{C}}_{s}$ is a $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$, Corollary 4 ensures $r_{\max}({\mathsf{C}}_{s})\geq 427-s=50$.
-On the other hand, we have $\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor=49$, and $d_{s}=\operatorname{d}_{GV}(2,n_{s},k_{s})=921$, so $k_{s}+1-d_{s}\ll 0$.
-Thus we expect to distinguish at $\beta_{r_{s}-1,r_{s}}$ with $r_{s}=50$.
-Now, for $i\leq r_{s}$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}$.
-As $d_{s}^{\perp}=\operatorname{d}_{GV}(2,n_{s},n_{s}-k_{s})=55>r_{s}$, these $\beta_{i-1,i}$ can be estimated as $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$.
-We then find that the maximal dimension of these matrices is around $2^{223}$, and the complexity around $2^{529}$.
+We compared our new distinguisher (in this basic version) with the one from [7], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$.
+Results can be found at the end of the Supplementary material.
+In both cases, we improve the distinguishability threshold.
+Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (86) (with $r=r_{\max}^{*}=4$), i.e.
+the conditions in Heuristic 1 are always met.
+On the other hand, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$, the conditions in Heuristic 1 are the limiting factor (but still improving over [7]).
+A general description of the set of distinguishable parameters is complicated, as it has to encompass both (86) and the conditions in the heuristic.
+However, a very interesting phenomenon will appear in the asymptotic analysis: all these constraints will be automatically satisfied, i.e.
+asymptotically all parameters will become distinguishable.
+Full version (with shortening)
+We can also use shortened dual codes as input in our distinguisher.
+In general, shortening $s$ times a $[n,k]$-code ${\mathsf{C}}$ produces a $[n_{s},k_{s}]$-code ${\mathsf{C}}_{s}$ with $n_{s}=n-s$, $k_{s}=k-s$.
+Having smaller parameters will decrease the complexity.
+Moreover the (dual) rate $R_{s}=k_{s}/n_{s}$ also decreases when we shorten; as observed in Remark 2, this will help satisfying the conditions in Heuristic 1.
+Example 1
+Consider the class $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ of dual binary Goppa codes with $m=10$, $n=1024$, $t=10$, irreducible Goppa polynomial.
+Without shortening, these codes are out of reach of our distinguisher: we would have to run Algorithm 1 up to $r=10$ and do linear algebra in dimension roughly $10^{20}$, which is infeasible.
+On the other hand, we can effectively distinguish after shortening $40$ times: we experimentally find that $40$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ have $\beta_{3,4}=30$, while $\beta_{3,4}=0$ for random codes with the same parameters.
+Computational cost (several days each, on a workstation with several hundreds of GB of RAM) prevented the author to run more than a few tests, but no deviation was observed.
+In particular $\beta_{3,4}=0$ consistently in the random case confirms Heuristic 1 for these parameters.
+Proposition 8
+Fix a class $\mathcal{T}_{q,m,n,t}$ of dual alternant or Goppa codes with some given parameters $q,m,n,t$, and assume that for some integer $r^{*}$ we have a lower bound of the form
+$r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ (87)
+for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of ${\mathsf{C}}$.
+Set $k=mt$, and let ${\mathsf{C}}$ be either a random $[n,k]_{q}$-code, or an element of $\mathcal{T}_{q,m,n,t}$.
+Choose an integer $s$ such that
+$\frac{k_{s}(k_{s}+1)}{n_{s}}<r^{*}-s$ (88)
+where $n_{s}=n-s$ and $k_{s}=k-s$.
+  1. 1.
+    Let $d_{s}\approx\operatorname{d}_{GV}(q,n_{s},k_{s})$ be the typical minimum distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$.
+Then, under part 1.
+of Heuristic 1, computing
+    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s}),$ (89)
+    where ${\mathsf{C}}_{s}$ is any $s$-shortened subcode of ${\mathsf{C}}$, allows to distinguish whether ${\mathsf{C}}$ is a random code (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ w.h.p.) or an element of $\mathcal{T}_{q,m,n,t}$ (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$).
+  2. 2.
+    Moreover, let $d_{s}^{\perp}\approx\operatorname{d}_{GV}(q,n_{s},n_{s}-k_{s})$ be the typical dual distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
+Then, under Heuristic 1, for large enough parameters the complexity of this distinguisher is dominated by the sum
+    $\kappa=\sum_{i\leq\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1}\max\left(k_{s}\operatorname{ind}(\varphi_{i-1}),\binom{k_{s}+1}{2}\operatorname{ind}(\varphi_{i-2})\right)^{\omega}$ (90)
+    where $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$ and $\omega\approx 2.372$ is the exponent of linear algebra.
+Proof
+If ${\mathsf{C}}$ is random, so is ${\mathsf{C}}_{s}$.
+Then, as $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 1.
+of Heuristic 1 we indeed expect $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ for $r^{*}-s>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
+On the other hand, for ${\mathsf{C}}$ in $\mathcal{T}_{q,m,n,t}$, the lower bound $r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ precisely means $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$.
+Now using Algorithm 1 we obtain $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $2\leq i\leq r^{*}-s$.
+The contributions for $i<4$ are negligible.
+For $i\geq 4$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}({\mathsf{C}}_{s})\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{s})$ and the contributions for $i>\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1$ are negligible because of part 1.
+Then, as $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 2.
+of Heuristic 1 we expect $\beta_{i-1,i}({\mathsf{C}}_{s})=\operatorname{ind}(\varphi_{i})$ for $2\leq i\leq\frac{k_{s}(k_{s}+1)}{n_{s}}$ for random ${\mathsf{C}}$; and we can assume this holds also for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
+From this we conclude.
+Example 2
+The following table shows how the distinguisher from Proposition 8 applies to Classic McEliece parameters.
+We take $r^{*}=\widehat{f}$ as given by Corollary 4, and then choose the maximum possible shortening order $s$ satisfying (88).
+$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr(n,m,t)&(3488,12,64)&(4608,13,96)&(6688,13,128)&(6960,13,119)&(8192,13,128)\\ \hline\cr r^{*}&427&683&939&867&939\\ \hline\cr s&377&568&816&769&848\\ \hline\cr r^{*}-s&50&115&123&98&91\\ \hline\cr[n_{s},k_{s}]&[3111,391]&[4040,680]&[5872,848]&[6191,778]&[7344,816]\\ \hline\cr\frac{k_{s}(k_{s}+1)}{{n_{s}}_{\phantom{0\!\!}}}&49.27&114.62&122.61&97.89&90.78\\ \hline\cr\operatorname{d}_{GV}&921&1069&1650&1828&2256\\ \hline\cr\operatorname{d}_{GV}^{\perp}&55&62&122&108&110\\ \hline\cr\kappa&{2^{528}}^{\phantom{i\!}}&(2^{1080})&(2^{1224})&2^{1030}&2^{997}\\ \hline\cr\end{array}$
+We observe that the condition $\operatorname{d}_{GV}(n_{s},k_{s})>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$ is always satisfied with a large margin, so we could choose a smaller $s$ and still be able to distinguish.
+Concerning the complexity estimate $\kappa$, it relies on the condition $\operatorname{d}_{GV}(n_{s},n_{s}-k_{s})>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
+We see this condition is satisfied for the parameter sets $(3488,12,64)$, $(6960,13,119)$, $(8192,13,128)$, but for $(4608,13,96)$ and $(6688,13,128)$ it is not (then, it would still be possible to give a complexity estimate, using Lemma 5 instead of Heuristic 1).
+These complexities improve those from [7], although they remain practically unreacheable and well beyond security levels.
 Asymptotics
 Fix a base field cardinality $q$, for instance $q=2$, and a (dual) rate $R$.
-In [4] it is suggested to work with a primal code of rate between $0.7$ and $0.8$, so passing to the dual gives $0.2\leq R\leq 0.3$.
-For $n\to\infty$ we set:
+In [4] it is suggested to take a primal code of rate between $0.7$ and $0.8$, so passing to the dual gives $0.2\leq R\leq 0.3$.
+However here we allow any $R$.
+Then for $n\to\infty$ set:
   • $m=\lceil\log_{q}(n)\rceil=\log_{q}(n)+O(1)$
@@
   • $t=\frac{k}{m}\;$ is an integer.
+A key observation then is that the lower bound $f$ on $r_{\max}$ of dual alternant codes in Theorem 3.1, or equivalently, the number $f$ of columns of the matrix $\mathbf{\Phi}$ in (65), is very close to $k$:
+Lemma 11
+We have
+$f=\left(1-\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}+O\left(\frac{1}{\log_{q}(n)}\right)\right)k.$ (91)
+Proof
+Direct consequence of:
+  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
+  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$.
+In the case of binary Goppa codes, the $\widehat{f}$ from Corollary 4 improves the $f$ from Theorem 3.1.
+However one could show that both have the same asymptotics.
+So, in order to distinguish these codes from random codes, we will not care about the extra structure: treating Goppa codes as alternant codes will suffice.
+(Still, this could be used to distinguish Goppa from “general” alternant codes.)
+Now, under Heuristic 1, we have:
 Theorem 5.1
-Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}$, or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes.
-Assume we have a lower bound of the form
-$r_{\max}({\mathsf{C}}_{s})\geq k-s^{*}-s$ (91)
-with $s^{*}=s^{*}(\mathcal{T}_{q,m,n,t})=o(n)$, valid for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of ${\mathsf{C}}$.
-Let then $r=o(n)$ be an integer such that $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$.
-Now let ${\mathsf{C}}$ be a $[n,k]_{q}$-code.
-Then computing
-$\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r}),$ (92)
-where ${\mathsf{C}}_{k-s^{*}-r}$ is a $(k-s^{*}-r)$-shortening of ${\mathsf{C}}$, allows to distinguish with high probability whether ${\mathsf{C}}$ is a random code ($\beta_{r-1,r}=0$) or an element of $\mathcal{T}_{q,m,n,t}$ ($\beta_{r-1,r}>0$), by doing at most $r$ kernel computations of linear maps between vector spaces whose maximal dimension can be estimated as $(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$.
-In fact it is possible to take $s^{*}\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ and $r\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n$, which gives for the total complexity of the process an estimated upper bound of
-$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}.$ (93)
-Observe [5, §3.4] that the security parameter in the Classic McEliece system, based on the complexity of generic decoding algorithms, is linear in $t\propto n/\log(n)$.
-Then in (93) we have $\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$, which means that our complexity is subexponential in the security parameter.
+Asymptotically, $q$-ary alternant (including Goppa) codes of dual rate $R$ can be distinguished from random codes, with complexity at most
+$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}$ (92)
+where $\omega\approx 2.372$ is the exponent of linear algebra.
 Proof
-The shortened code ${\mathsf{C}}_{k-s^{*}-r}$ has dimension $s^{*}+r$ and length $n-k+s^{*}+r$.
-As $s^{*}$ and $r$ are $o(n)$ we have $s^{*}+r\ll n-k+s^{*}+r$, and $\operatorname{d}_{GV}$ is close to the length.
-Moreover $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$ by hypothesis.
-Thus Heuristic 1 applies, and we expect $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})=0$ with high probabilty when ${\mathsf{C}}$ is random.
-On the other hand, (91) gives $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})>0$ for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$.
-Using Algorithm 1 we obtain $\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $(s^{*}+r)\beta_{i-2,i-1}({\mathsf{C}}_{k-s^{*}-r})\times\binom{s^{*}+r+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{k-s^{*}-r})$.
-As the rate $\frac{s^{*}+r}{n-k+s^{*}+r}$ goes to zero, by Heuristic 1 again we can estimate the largest dimension of this matrix as $\binom{s^{*}+r+1}{2}\operatorname{ind}(\varphi_{i-2,{\mathsf{C}}_{k-s^{*}-r}})\leq(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$, for random codes — and we can assume this holds also for codes in $\mathcal{T}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
-Now for a concrete choice of $s^{*}$ and $r$, set
-  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
-  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$
-so
-$f=\left(1-\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}+O\left(\frac{1}{\log_{q}(n)}\right)\right)k.$ (94)
-By Theorem 3.1, for all $s$ and all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ we have $r_{\max}({\mathsf{C}}_{s})\geq f-s$, so we can take $s^{*}=k-f\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ as claimed.
-(In the binary Goppa case we could use Corollary 4 instead of Theorem 3.1, but this leads to the same estimate on $f$, hence on $s^{*}$.) Then for $\varepsilon>0$, setting $r=\left\lceil(1+\varepsilon)\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\right\rceil\ll s^{*}$, we have $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}\approx\frac{(s^{*})^{2}}{n-k}\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\ll r$, and then we can let $\varepsilon\to 0$.
-Last, for the complexity estimate (93), using $r\ll s^{*}$ again and Stirling’s formula, we get $\log_{q}\left((s^{*}+r)^{3}\binom{s^{*}+r}{r-3}\right)\approx r\log_{q}\left(\frac{s^{*}}{r}\right)$, and we conclude.
-We observe it is possible to slightly relax the condition $m=\lceil\log_{q}(n)\rceil$: complexity remains subexponential as long as $\frac{m}{\log_{q}(n)}\to 1$ fast enough.
+We use Proposition 8 with $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, and $r^{*}=f$ as provided by Theorem 3.1.
+The analysis will be easier if rephrased in terms of $k_{r^{*}}=k-r^{*}$ and $r=r^{*}-s$.
+First, Lemma 11 gives
+$k_{r^{*}}\;=\;k-f\;\sim\;R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n.$ (93)
+Let us then turn to condition (88). Fix an $\varepsilon>0$, and set
+$r=\left\lceil(1+\varepsilon)\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n\right\rceil\ll k_{r^{*}}$ (94)
+and $s=r^{*}-r$. Then $k_{s}=k_{r^{*}}+r$ and $n_{s}=n-k+k_{r^{*}}+r$, so for $n$ large enough we get
+$\frac{k_{s}(k_{s}+1)}{n_{s}}=\frac{(k_{r^{*}}+r)(k_{r^{*}}+r+1)}{n-k+k_{r^{*}}+r}\approx\frac{k_{r^{*}}^{2}}{n-k}\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n<r=r^{*}-s.$ (95)
+Thus, for any $\varepsilon>0$, condition (88) is satisfied with this value of $r$ as soon as $n$ is large enough, say larger than some $n(\varepsilon)$.
+But this means precisely that as $n\to\infty$, we can let $\varepsilon=\varepsilon(n)\to 0$ while still satisfying the condition.
+We then get
+$r\;\sim\;\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n.$ (96)
+Moreover the shortened code ${\mathsf{C}}_{s}$ has rate $\frac{k_{s}}{n_{s}}=\frac{k_{r^{*}}+r}{n-k+k_{r^{*}}+r}=o(1)$, so by Remark 2 both conditions in Heuristic 1 are satisfied.
+Thus Proposition 8 applies.
+[…]
+In [5, §3.4], the security level $2^{b}$ — or less formally, the number $b$ of targeted “bits of security” — of the Classic McEliece system, essentially defined as the complexity of the currently best attack (namely, by information set decoding), is shown asymptotically to be linear in $t\propto n/\log(n)$.
+Then in (92) we have
+$\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$
+which means that the complexity of our distinguisher is subexponential in $b$ (although, admittedly, only very slightly so).
 6 Conclusion and open problems
-We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms, more precisely, subexponential in the error-correcting capability of the code.
-However this is only an asymptotic result.
-For concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the security parameter by a non-negligible factor.
+We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms; more precisely, subexponential in the error-correcting capability of the code.
+However this is only an asymptotic result: for concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the best attacks by a non-negligible factor.
 Problem 1
-To what extent can the implementation of the distinguisher be improved?
-For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be:
-  • somewhat sparse
-  • already partially reduced.
+Improve the implementation of this distinguisher.333addendum: actually, the improvements discussed below can already be found in works of Kreuzer-Lusteck and Albano-La Scala
+For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be (i) somewhat sparse and (ii) already partially reduced.
 This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster.
-More generally, optimizations used in related domains, for instance in Gröbner basis algorithms, could probably be imported.
-(Interestingly, the author was led to introduce the shortening technique in this distinguisher by serendipity, after witnessing a strange phenomenon while partially implementing the optimization alluded in Problem 1.)
+Alternatively, instead of iteratively computing a whole range of the linear strand, one could directly compute a targeted Betti number, e.g.
+using Koszul cohomology (the author would like to thank W.
+Castryck for this suggestion).
+It is then likely that this Koszul cohomology computation could also benefit from similar optimizations (i) and (ii) exploiting sparsity and structure.
 Problem 2
@@
 Better, can one give a complete, explicit description of the minimal resolution of these codes?
-(See the Supplementary material for elements in this direction.) How would this impact the complexity of the distinguisher?
-We initiated the study of syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
+(See at the beginning of the Supplementary material for elements in this direction.)
+We initiated the study of higher syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
 Problem 4
 Pursue this study in a systematic way.
-This should cover basic operations on codes that have a standard geometric interpretation, such as puncturing (discarding one subset from a set of points), shortening (projection from one or several of the points), duality of codes (“Gale duality”), tensor product of codes (Segre embedding), and powers of codes (Veronese embedding), but also other operations such as the $(u|u+v)$, subfield subcode, trace code, or concatenated code constructions; and crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
+This should cover basic operations on codes (some of which have a nice geometric interpretation, some of which less so), and also crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
 Problem 5
-Can syzygies serve in decoding algorithms?
+Can higher syzygies serve in decoding algorithms?
 Problem 6
-Can our syzygy distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack?
-For instance, one would like to combine our results with:
+Betti numbers actually are invariants of the monomial equivalence class of a code (Lemma 3). Can they find other applications, for instance, to problems of code equivalence?
+Problem 7
+Can our distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack?
+For instance, one would like to combine our approach with:
   • filtration arguments
@@
 Once digested by the community, they are then used in a more constructive way, to build new cryptosystems.
-Problem 7
+Problem 8
 Can one devise syzygy-based cryptography?
+{credits}
+6.0.1 Acknowledgements
+The author learned a lot from the code-based cryptography seminar organized by Jean-Pierre Tillich at Inria Paris.
+More important perhaps, without his friendly pressure, probably this work would have stayed eternally in limbo.
+He should be thanked for all that.
+The very favourable environment towards research at ANSSI in general, and at LCR in particular, was also highly helpful.
+Henri Gilbert’s advices led to improvements in several aspects of the presentation.
+Anonymous members of LCR unsuccessfully suggested the title “The Dyztynguysher” for this work.
+The author is part of the ANR-21-CE39-0009 BARRACUDA and ANR-21-CE39-0008 KLEPTOMANIAC research projects. Some of the computations were done on the lame servers in Télécom Paris.
 Supplementary material
@@
 Now for general $s$, (68) and (73) can be extended to:
-Experimental fact 2
+Experimental fact 3
 Let ${\mathsf{C}}$ be a dual alternant code. Then for all $s$ we have
@@
 Back to dual alternant or Goppa codes, we actually observe a strong regular pattern (for Goppa codes the author only tested the irreducible case, but the result is likely to generalize):
-Experimental fact 3
+Experimental fact 4
 Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$ or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes, namely, either dual alternant codes, or dual Goppa codes with irreducible Goppa polynomial.
@@
 Figure 4: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{2,10,5}\;$ ($r^{*}=8$)
-$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ 8&18&-&-&-&-&-&-\end{array}$
+$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ \end{array}$
 Figure 5: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{3,5,6}\;$ ($r^{*}=8$)
@@
 Figure 8: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}\;$ ($r^{*}=38$)
-These Figures illustrate the last Experimental fact.
+Figures 8-8 illustrate the last Experimental fact.
 Boldface values match (107).
 This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
+Computing the Betti diagram of some families of codes
+Regularity $2$ helps in computing the full Betti diagram of codes, as one row can be deduced from the other.
+In order to illustrate this, let us first recall that a $[n,k]$-code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
+Lemma 12 ([15, Th. 1], reformulated)
+Let ${\mathsf{C}}$ be a $[n,k]$-code with $n\leq 2k-1$. Assume ${\mathsf{C}}$ is MDS. Then for all $r\leq 2k+1-n$ we have $\beta_{r-2,r}({\mathsf{C}})=0$.
+Proposition 10
+Let ${\mathsf{C}}$ be a $[k+1,k]$ MDS code, for instance a parity code or a $[k+1,k]$ $\operatorname{{\mathsf{GRS}}}$ code.
+Then the nonzero Betti numbers of ${\mathsf{C}}$ are $\beta_{0,0}=\beta_{k-1,k+1}=1$, and
+$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (108)
+for $12\leq r\leq k-$ 1. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
+Proof
+The parameters imply that ${\mathsf{C}}$ has regularity $2$.
+By Lemma 12 we have $\beta_{r-2,r}=0$ for $r\leq k$.
+Then $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-(k+1)\right)\binom{k-1}{r-2}$ and we conclude with a straightforward calculation.
+Proposition 11
+Let ${\mathsf{C}}$ be a $[2k-1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are:
+  • $\beta_{0,0}=1$
+  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
+  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
+In particular the ideal $I({\mathsf{C}})$ is generated by $\beta_{1,2}=\frac{(k-1)(k-2)}{2}$ quadratic forms and $\beta_{1,3}=k-1$ cubic forms.
+Proof
+Using the basis $\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}$ of ${\mathsf{C}}$, we see $I_{2}({\mathsf{C}})$ contains the $2\times 2$ minors of $\mathbf{\Phi}=\left(\begin{array}[]{cccc}X_{0}&X_{1}&\dots&X_{k-2}\\ X_{1}&X_{2}&\dots&X_{k-1}\end{array}\right)$, i.e.
+$I_{2}({\mathsf{C}})\supseteq\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ of dimension $\binom{k-1}{2}$.
+Now set $n=2k-1$.
+Because $n\geq 2k-1$, we have $\dim{\mathsf{C}}^{\langle 2\rangle}=2k-1$, hence $\dim I_{2}({\mathsf{C}})=\binom{k-1}{2}$, so $I_{2}({\mathsf{C}})=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$.
+Corollary 1 and the Eagon-Northcott complex of $\mathbf{\Phi}$ then give $\beta_{r-1,r}=(r-1)\binom{k-1}{r}$.
+Because $n\leq 2k-1$, we have ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$.
+Then $\beta_{r-2,r}=\beta_{r-1,r}-\operatorname{ind}(\varphi_{r})=(r-1)\binom{k-1}{r}-\left(\frac{k(k+1)}{r}-(2k-1)\right)\binom{k-1}{r-2}$, and a straightforward calculation allows to conclude.
+Figures 10 and 10 illustrate Propositions 10 and 11.
+$\begin{array}[]{c|cccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&27&105&189&189&105&27&-\\ 2&-&-&-&-&-&-&-&1\end{array}$
+Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
+$\begin{array}[]{c|ccccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&21&70&105&84&35&6&-\\ 2&-&7&42&105&140&105&42&7\end{array}$
+Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
+Exercise 1
+Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for $k+1<n<2k-1$.
 Experimental data on defects for random codes
-Figure 13 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes.
+Figures 13-13 compare Betti diagrams for various types of codes of length $n=23$ and dimension $k=12$. Observe that these parameters give $\frac{k(k+1)}{n}\approx 6.78$.
+First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e.
+with all defects equal to zero.
+It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram.
+Such codes are MDS, i.e.
+they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified.
+Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g.
+$q=5$ suffices (the author did not search for optimality).
+These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified.
+Moreover, the converse Experimental fact 2 is also verified as all defects are zero.
+$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&880&1353&990&-&-&-&-&-&-\\ 2&-&-&-&-&-&330&1617&1870&1221&485&110&11\\ \end{array}$
+Figure 11: an idealized $[23,12]$-code according to the minimal resolution conjecture
+$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1397&1224&490&121&18&1&-&-\\ 2&-&-&4&44&234&820&1738&1888&1222&485&110&11\\ \end{array}$
+Figure 12: a (pseudo)random $[23,12]_{2}$-code ($d=3$, $d^{\perp}=4$)
+$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
+Figure 13: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
+Second, Figure 13 presents the Betti diagram of the binary code whose generator matrix
+$G=\left(\begin{array}[]{ccccccccccccccccccccccc}1&1&0&0&1&0&0&1&0&0&0&0&1&1&1&1&1&1&0&1&1&0&1\\ 0&1&0&1&0&0&0&1&0&0&0&1&0&0&0&0&1&0&1&1&0&1&0\\ &&&&&\vdots&&&&&&&&&&&&\vdots&&&&&\\ 0&1&0&0&1&0&1&0&0&0&1&0&1&0&0&1&0&1&0&0&0&0&0\end{array}\right)$
+of size $12\times 23$, is constructed from the binary expansion of $\pi$.
+We consider this code as representative of random codes.
+One can check that it has $d=3$, $d^{\perp}=4$.
+Then Experimental fact 1 is verified, as the defect is positive for $4\leq r\leq 10$.
+Moreover at the extremities of this interval we have equality $\beta_{2,4}=A_{4}(C^{\perp})=4$ and $\beta_{9,10}=A_{3}(C)=1$, as predicted.
+And the defect is zero for $r<4$ and $r>10$ so Experimental fact 2 is verified also.
+Last, Figure 13 reproduces Figure 3, the binary Golay code.
+As $d,d^{\perp}\geq 7$, Experimental fact 1 is trivially verified.
+On the other hand, Experimental fact 2 is not verified, but this only illustrates the fact that the Golay code is certainly not representative of random codes.
+Indeed it has a lot of algebraic structure, which explains it admitting special syzygies.
+More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes.
 For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling).
@@
 Here we have $\frac{k(k+1)}{n}\approx 4.86$.
-Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable), while $\operatorname{def}(\varphi_{r})=0$ with high probability as we move away from these intervals.
+Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
 $d=11$
@@
 Figure 14: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes
+Distinguishing Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
+For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{97}$, while the distinguisher of [7] works down to $n_{\text{CMT}}=\mathbf{80}$.
+We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, which is better than the lower bound in Theorem 3.1 ($e=0$, $r_{\max}\geq t-1=3$).
+Moreover, their Betti numbers are:
+$\beta_{1,2}^{*}=40,\qquad\beta_{2,3}^{*}=80,\qquad\beta_{3,4}^{*}=12.$ (109)
+We then observe that the $\beta_{2,3}$-distinguisher works down to $n_{\beta_{2,3}}=\mathbf{86}$, and the $\beta_{3,4}$-distinguisher works down to $n_{\beta_{3,4}}=\mathbf{68}$, both of which coincide with (86).
+More precisely, computing the Betti numbers of dual Goppa and random codes around these values of $n$ consistently yields:
+$\begin{array}[]{|c|ccccccccccccc|}\hline\cr n&\cdots&88&87&\mathbf{86}&85&84&\cdots&70&69&\mathbf{68}&67&66&\cdots\\ \hline\cr\beta_{2,3}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&80&80&\mathbf{80}&85&100&&310&325&340&355&370&\\ \beta_{2,3}^{\text{random}}&&40&55&\mathbf{70}&85&100&&310&325&340&355&370&\\ \hline\cr\beta_{3,4}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&12&12&12&12&12&&12&12&\mathbf{12}&105&210&\\ \beta_{3,4}^{\text{random}}&&0&0&0&0&0&&0&0&\mathbf{0}&105&210&\\ \hline\cr\end{array}$
+We see that they stick to their “predicted” values: $\beta_{r-1,r}=\max(\beta_{r-1,r}^{*},\operatorname{ind}(\varphi_{r})^{+})$ for dual Goppa, and $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})^{+}$ for random codes.
+Distinguishing Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
+For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{62}$, while the distinguisher of [7] works down to $n_{\text{CMT}}=\mathbf{59}$.
+We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, which is better than the lower bound in Corollary 4 ($\widehat{e}=1$, $r_{\max}\geq 7$).
+Moreover, their top Betti numbers are:
+$\beta_{5,6}^{*}=1020,\qquad\beta_{6,7}^{*}=288,\qquad\beta_{7,8}^{*}=42.$ (110)
+From (86) we expect to distinguish at $\beta_{5,6}$ for $n\geq 57$, at $\beta_{6,7}$ for $n\geq 49$, and at $\beta_{7,8}$ for $n\geq 43$.
+And indeed at $n_{\beta_{5,6}}=\mathbf{57}$ we consistently find $\beta_{5,6}\geq 1020$ for dual Goppa codes, while $\beta_{5,6}<500$ for random codes with quite high probability.
+For smaller $n$ we have to pass to $\beta_{6,7}$.
+The $\beta_{6,7}$-distinguisher works well for $n=56$, but the quality gradually falls (distinguishing errors occur more frequently) as $n$ becomes smaller.
+It is difficult to point a precise threshold where the distinguisher ceases to work.
+Arguably we still have a positive advantage at $n_{\beta_{6,7}}=\mathbf{50}$, but not anymore at $n=49$.
+We could then try with $\beta_{7,8}$, but this fails too.
+What happens?
+It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$.
+Indeed, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$.
+For $r=7$ and $k=mt=18$ this gives $d\geq 13$.
+Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$.
+Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity.
+Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably.
+Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
+Remark 3
+The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher.
+Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$.
+Let us thus consider a distinguisher that decides that a given $[n,n-18]$-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise.
+Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$.
+This means that this distinguisher has more than $99\%$ success rate!
+And it remains remarkably effective even for smaller $n$.
+For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.
+In these experiments, random codes were defined by taking a uniformly random $(n-18)\times n$ generator matrix, which might quite often produce degenerate codes as $n$ approaches $18$.
+Other probability distributions preventing this problem could be used, which would slightly alter the success rate, but not the qualitative behaviour of this distance-based distinguisher.

### v6

@@
 The invariants that allow us to distinguish are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code.
-Since its introduction in 1978, this is the first time an analysis111 we exclude results such as [28] that use an attack model for which direct countermeasures exist of the McEliece cryptosystem breaks the exponential barrier.
+Since its introduction in 1978, this is the first time an analysis (in the CPA model) of the McEliece cryptosystem breaks the exponential barrier.
 1 Introduction
@@
 However, to an attacker, the public code looks like a random code, and removing the noise is untractable.
-Slightly more formally, a traditional security proof for the system relies on two assumptions:
+At the time of publication of [29] it was not customary to fully formalize the expected security properties of the scheme.
+In modern language we would say that it aims at one-wayness against chosen plaintext attacks (OW-CPA).
+There are now variants of the McEliece scheme that claim stronger security properties (e.g.
+IND-CCA2), but for this they rely on the hypothesis that the original McEliece scheme is OW-CPA.
+This makes it even more relevant today to scrutinize whether this hypothesis can be reduced to more fundamental theoretical assumptions.
+And indeed, as the informal description of the system naturally suggests, a traditional security proof reduces this hypothesis to the following two assumptions:
   1. 1.
@@
   • Key recovery attacks address the computational version: recover the Goppa structure of the code, or at least an equivalent one [16], if it exists.
-Although certain arguments, such as the fact that the class of Goppa codes is very large, make assumption 1 plausible, it remains quite ad hoc from a theoretical point of view.
+Assumption 1 was first introduced formally in [9].
+Although certain arguments, such as the fact that the class of Goppa codes is very large, make this assumption plausible, it remains quite ad hoc from a theoretical point of view.
 Its sole virtue is that it passed the test of time.
@@
 This suggests that if a weakness in assumption 1 were to be uncovered one day, then algebraic methods should be a tool of choice.
-However, up to now, the best distinguishers and key recovery algorithms such as those in [12][8][7][1] only apply to alternant or Goppa codes with very degraded parameters.
+However, up to now, the best distinguishers and key recovery algorithms such as those in [15][11][10][2][24][23] only apply to alternant or Goppa codes with very degraded parameters.
 Against McEliece with cryptographically relevant parameters, either they have exponential complexity with large constants, which makes them useless, or worse: they simply cease to work.
@@
 This might give a feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
-Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e.
+Preventing the spread of this belief, we present a new distinguisher for alternant and Goppa codes, i.e.
 a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms.
@@
 If this quantity is smaller than the said lower bound, we can distinguish.
-Likewise we claim that Theorem 2.8 of [10] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes,222erratum: among MDS $[7,4]$-codes (which then implies of full square) where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher.
-For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$.
+Likewise we claim that Theorem 2.8 of [13] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among MDS $[7,4]$-codes.
+Indeed, for such codes, the square always fills the whole space — so the square distinguisher does not apply — and we have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$.
 Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$.
@@
 and these satisfy the syzygies
-[…]
-Thus we can distinguish by computing $\beta_{2,3}({\mathsf{C}})=\dim\ker(\mathbf{M}_{3})$, which will yield $2$ for $\operatorname{{\mathsf{GRS}}}$ and 00 for non-$\operatorname{{\mathsf{GRS}}}$ $[7,4]$-codes.
+$X_{1}Q_{3}-X_{2}Q_{2}+X_{3}Q_{1}\;=\;X_{2}Q_{3}-X_{3}Q_{2}+X_{4}Q_{1}\;=\;0.$ (5)
+Thus we can distinguish by computing $\beta_{2,3}({\mathsf{C}})=\dim\ker(\mathbf{M}_{3})$, which will yield $2$ for $\operatorname{{\mathsf{GRS}}}$ and 0 for non-$\operatorname{{\mathsf{GRS}}}$ MDS $[7,4]$-codes.
 We generalize these examples to higher Betti numbers, following the exact same pattern:
@@
 Last, section 5 combines everything and chooses parameters in order to optimize asymptotic complexity.
-Of special importance in this regard is Proposition 4, which shows that the property that the ideal of a code contains minors of a matrix of linear forms passes to its shortened subcodes.
+Of special importance in this regard is Proposition 4 from section 3, which shows that the property that the ideal of a code contains minors of a matrix of linear forms passes to its shortened subcodes.
 Related (and unrelated) works
 As already noted, our distinguisher can be seen as a generalization of the square distinguisher of [15].
-Using an approach similar to ours, the work [7] also extends this square distinguisher by exploiting special properties of the space of quadratic relations, but in a different direction.
-Last, the key recovery attack in [1] combines shortening of the dual code with ideas from the square distinguisher, and a careful algebraic modeling in order to apply tools from Gröbner basis theory.
+Using an approach similar to ours, the work [10] also extends this square distinguisher by exploiting special properties of the space of quadratic relations, but in a different direction: the authors detect the existence of short relations by computing the Hilbert series of a convenient Pfaffian modeling (and in some cases they even manage to recover the key).
+Last, the key recovery attack in [2] combines shortening of the dual code with ideas from the square distinguisher to reduce parameters, and then performs a careful algebraic modeling exploiting the symmetries of a system, that can finally be solved with Gröbner basis algorithms.
 All these results have limited range of applicability, but they introduced numerous techniques that influenced the present work.
+After the first version of this text appeared, other works continued this series of results.
+A refined analysis of the Pfaffian distinguisher can now be found in [24].
+Using purely geometric methods, [23] presents a new key recovery attack; but quite surprisingly, this attack works precisely in the regimes where those of [10] and [2] apply!
+Last, the recent work [41] proposes a distinguisher based on higher degree relations vanishing with higher multiplicities.
 From a geometric point of view, the Betti numbers and the syzygies we consider are those of a set of points in projective space (namely, defined by the columns of a parity check matrix of the code).
@@
 This leads to different theories, although seeking links between the two could be an interesting project.
+Additional material and new results since the Eurocypt 2025 version
+There are significative differences between this text and its prior version [37].
+First, the supplementary material to [37] is now included in the main text, sometimes after some reworking.
+Moreover, several new results were introduced.
+  1. (a)
+    The discussion covering possible further improvements in the description of the minimal resolution of algebraic codes has been incorporated in section 3.
+After some reworking, it now includes Corollary 4 that provides an explicit family of linearly independent quadratic relations for dual alternant codes that matches the lower bound of [30, Th.
+20] (thus giving an alternative proof for this result while sticking to the approach of [15]).
+  2. (b)
+    A tight lower bound on the number of columns of a matrix whose $2\times 2$ minors give quadratic relations (and determine the $r_{\max}$) of Goppa codes is now stated in Remark 1 in section 3 (although its proof is deferred to the forthcoming work [35]).
+  3. (c)
+    The computation of the Betti diagram of $[k,k+1]$ MDS codes and of [k,2k1][k,2k-1] $\operatorname{{\mathsf{GRS}}}$ codes has been incorporated in section 4.
+  4. (d)
+    The presentation of experimental data on defects of random codes also has been incorporated in section 4.
+  5. (e)
+    The new Proposition 12 in section 4 proves a (very small) part of Experimental fact 1, linking the $r_{\max}$ of a code to its minimum distance.
+  6. (f)
+    The two examples of application of the distinguisher to some Goppa codes with small parameters have been incorporated in section 5.
+  7. (g)
+    A new variant of the distinguisher, based on Koszul cohomology and the (block) Wiedemann algorithm, has been introduced also in section 5.
+Some of these new results allow a noticeable improvement on the complexity of the distinguisher.
+Point (b) allows to predict the optimal number $s$ of coordinates to shorten.
+Point (g) allows to use linear algebra with exponent $2$ instead of $\omega$ (and also to relax the dependency of the complexity estimate from some of the heuristics).
+When applied e.g.
+to the Classic McEliece 348864 parameters, each allows to gain roughly 60 bits of complexity, and combined, they reduce the complexity estimate from $2^{528}$ to $2^{401}$.
 Notation and conventions
@@
 $\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}+\cdots+x_{n}y_{n}.$ (7)
-A $k$-dimensional subspace ${\mathsf{C}}\subseteq\mathbb{F}^{n}$ is called a $[n,k]$-code (and a $[n,k]_{q}$-code in case $\mathbb{F}=\mathbb{F}_{q}$).
+A $k$-dimensional subspace ${\mathsf{C}}\subseteq\mathbb{F}^{n}$ is called a $[n,k]$-code (and a $n,k]_{q}-$code in case $\mathbb{F}=\mathbb{F}_{q}$).
 The orthogonal space $\mathsf{C}$$^{\perp}$ is called the dual code of $\mathsf{C}$}.
@@
 $\begin{split}\operatorname{{\mathsf{Alt}}}_{t}(\mathbf{x},\mathbf{y})&=\operatorname{{\mathsf{GRS}}}_{t}(\mathbf{x},\mathbf{y})^{\perp}\cap(\mathbb{F}_{q})^{n}\\ &=\{\mathbf{c}\in(\mathbb{F}_{q})^{n}:\>c_{1}y_{1}x_{1}^{j}+\cdots+c_{n}y_{n}x_{n}^{j}=0\quad(0\leq j<t)\},\end{split}$ (17)
-with parameters $[n,(\geq)n-mt]_{q}$.
+with parameters $n,(\geq)n-mt]_{q}$.
 Last, given a polynomial $g(X)\in\mathbb{F}_{q^{m}}[X]$ that does not vanish on any entry of $\mathbf{x}$, the $q$-ary Goppa code with support $\mathbf{x}$ and Goppa polynomial $g$ is
@@
 If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$.
-Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $[g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
+Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
 Iterating this construction, we obtain a minimal resolution
@@
 $F_{i}=\bigoplus_{j\geq 0}S(-j)^{\beta_{i,j}}$ (24)
-where $S(-j)$ is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
+where S(j)S(-j) is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
 It is customary to display the graded Betti numbers in the form of a Betti diagram, as follows:
-$\begin{array}[]{c|ccccc}&0&1&\dots&i&\dots\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots\\ \vdots&\vdots&\vdots&&\vdots&\\ \end{array}$
-with null entries marked as “--” for readability.
+[…]
+with null entries marked as “-” for readability.
 Lemma 1
@@
 From now on let $\mathsf{C}$} be a $[n,k]$-code, with generator matrix $\mathbf{G}$, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring.
-Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$.
+Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length k1k-1.
 So we have an exact sequence
 $0\longrightarrow F_{k-1}\longrightarrow\cdots\longrightarrow F_{2}\longrightarrow F_{1}\longrightarrow F_{0}=S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}\longrightarrow 0$ (25)
-and the Betti diagram has $k$ columns, indexed from 00 to $k-1$.
-As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
+and the Betti diagram has $k$ columns, indexed from 0 to k1k-1.
+As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 0-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
 Likewise $M_{1}=I({\mathsf{C}})$, and $\beta_{1,j}$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$.
@@
 $F_{i,i+1}=M_{i,i+1}=\mathbb{F}^{\beta_{i,i+1}}.$ (26)
-Hence the 00-th row of the Betti diagram is always $(1,-,-,\dots)$.
+Hence the 0-th row of the Betti diagram is always $(1,-,-,\dots)$.
 Figures 3-3 offer a few such diagrams for contemplation.
-$\begin{array}[]{c|cccc}&0&1&2&3\\ \hline\cr 0&1&-&-&-\\ 1&-&3&-&-\\ 2&-&1&6&3\\ \end{array}$
-Figure 1: the $[7,4]_{2}$ Hamming code
-$\begin{array}[]{c|cccccc}&0&1&2&3&4&5\\ \hline\cr 0&1&-&-&-&-&-\\ 1&-&10&16&-&-&-\\ 2&-&1&5&26&20&5\\ \end{array}$
-Figure 2: the $[11,6]_{3}$ Golay code
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
-Figure 3: the $[23,12]_{2}$ Golay code
+[…]
+Figure 1: the $7,4]_{2}$ Hamming code
+[…]
+Figure 2: the $11,6]_{3}$ Golay code
+[…]
+Figure 3: the $23,12]_{2}$ Golay code
 Observe that replacing $\mathbf{G}$ with $\mathbf{S}\mathbf{G}$, for $\mathbf{S}\in\mathbb{F}^{k\times k}$ an invertible matrix, corresponds to a linear change of coordinates for the variables $X_{1},\dots,X_{k}$.
@@
 We already discussed the action of $\mathbf{S}$. As for the action of $\mathbf{P}\mathbf{D}$, it preserves the kernel of the graded evaluation map $S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$.
-Effective computation
+Effective computation of linear syzygies
 There are several algorithms to compute minimal resolutions and graded Betti numbers in general.
@@
 For this we just observe that the diagram
-[…]
+$M_{r-2,r}$$M_{r-2,r-1}\otimes S_{1}$$F_{r-3,r}$$M_{r-3,r-2}\otimes S_{2}=F_{r-3,r-2}\otimes S_{2}$$\scriptstyle{\varphi_{r}}$$\scriptstyle{\psi_{r}}$ (32)
 commutes, with the two arrows on the right injective.
@@
 $\mathbf{M}_{3}\in\mathbb{F}^{\>k\beta_{1,2}\,\times\,\binom{k+2}{3}}$ (34)
-with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 00 elsewhere.
+with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 0 elsewhere.
 (Said otherwise, the rows of $\mathbf{M}_{3}$ are formed by multiplying all $\mathbf{q}\in\mathcal{B}_{2}$, seen as quadratic forms $\mathbf{q}=\sum_{1\leq b\leq c\leq k}q_{X_{b}X_{c}}X_{b}X_{c}$, by all variables $X_{a}$.)
@@
 Thus $\mathcal{B}_{3}$ consists of $\beta_{2,3}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$.
-Now let $r\geq 4$, and assume inductively for all $13\leq i\leq r-$ 1 we have constructed a basis $\mathcal{B}_{i}$ of $M_{i-1,i}$, consisting of $\beta_{i-1,i}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{i-1}$.
+Now let $r\geq 4$, and assume inductively for all $3\leq i\leq r-1$ we have constructed a basis $\mathcal{B}_{i}$ of $M_{i-1,i}$, consisting of $\beta_{i-1,i}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{i-1}$.
 Definition 3
@@
 $\mathbf{M}_{r}\in\mathbb{F}^{\>k\beta_{r-2,r-1}\,\times\,\binom{k+1}{2}\beta_{r-3,r-2}}$ (35)
-with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 00 elsewhere.
+with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 0 elsewhere.
 (Said otherwise, the rows of $\mathbf{M}_{r}$ are formed by multiplying all $\mathbf{s}\in\mathcal{B}_{r-1}$ by all variables $X_{a}$, where each $\mathbf{s}$ is seen as a formal sum $\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}L_{\mathbf{t}}[\mathbf{t}]$ whose coefficients are linear forms $L_{\mathbf{t}}=\sum_{1\leq b\leq k}s_{X_{b},\mathbf{t}}X_{b}$, resulting in formal sums $X_{a}\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}(X_{a}L_{\mathbf{t}})[\mathbf{t}]$ whose coefficients are quadratic forms.)
@@
 8:end for
-Observe that our computation only relies on mere linear algebra, and does not make use of any Gröbner basis theory.
+An implementation can be found in [34], which was used to compute all the examples in this text.
+Observe that this algorithm only relies on mere linear algebra, and does not make use of any Gröbner basis theory.
 However, the two topics are clearly related.
-In particular, the many linear algebra optimizations used in Gröbner basis algorithms could certainly apply in our context.
+In particular, the many linear algebra optimizations used in Gröbner basis algorithms certainly apply here also.
+This will be discussed later in the text, when we’ll deal with complexity estimates.
 Further properties
@@
 $r_{\max}({\mathsf{C}})=\max\{r:\>\beta_{r-1,r}({\mathsf{C}})>0\}.$ (36)
-As the minimal resolution of a $[n,k]$-code ${\mathsf{C}}$ has length $k-1$, we always have $r_{\max}({\mathsf{C}})\leq k$.
+As the minimal resolution of a $[n,k]$-code $\mathsf{C}$} has length k1k-1, we always have $r_{\max}({\mathsf{C}})\leq k$.
 Sometimes we only control the syzygies of certain subideals of $I({\mathsf{C}})$, and from these we want to deduce information on the syzygies of the whole ideal.
@@
 $0\to(\operatorname{Sym}^{f-g}G)^{\vee}\otimes\bigwedge^{f}F\to\cdots\to G^{\vee}\otimes\bigwedge^{g+1}F\to\bigwedge^{g}F\overset{\bigwedge^{g}\mathbf{\Phi}}{\longrightarrow}\bigwedge^{g}G\simeq R.$ (46)
-Under mild hypotheses, this complex is exact [10, Th.
+Under mild hypotheses, this complex is exact [13, Th.
 A.2.60 & Th.
 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
-In this work we will only need the case $g=2$. In this case the complex has length $f-1$, and for $2\leq r\leq f$ its $(r-1)$-th module is free of rank
+In this work we will only need the case $g=2$. In this case the complex has length f1f-1, and for $2\leq r\leq f$ its (r1)(r-1)-th module is free of rank
 $\operatorname{rk}\left((\operatorname{Sym}^{r-2}G)^{\vee}\otimes\bigwedge^{r}F\right)\;=\;(r-1)\binom{f}{r}$. (47)
@@
 In the multivariate polynomial ring
-$R[Z^{(j)}_{r;i_{1},\dots,i_{r}}]$ (52)
-where, for each $2\leq r\leq f$, indices range over the $(r-1)\binom{f}{r}$ values $11\leq j\leq r-$ 1 and $1\leq i_{1}<\dots<i_{r}\leq f$, construct polynomials $s^{(j)}_{r;i_{1},\dots,i_{r}}$ as follows.
+$R[Z_{r,j;i_{1},\dots,i_{r}}]$ (52)
+where, for each $2\leq r\leq f$, indices range over the $(r-1)\binom{f}{r}$ values $1\leq j\leq r-1$ and $1\leq i_{1}<\dots<i_{r}\leq f$, construct polynomials $s_{r,j;i_{1},\dots,i_{r}}$ as follows.
 First for $r=2$ we define the constant polynomials
-$s^{(1)}_{2;i_{1},i_{2}}=x_{i_{1}}x^{\prime}_{i_{2}}-x^{\prime}_{i_{1}}x_{i_{2}}$ (53)
+$s_{2,1;i_{1},i_{2}}=x_{i_{1}}x^{\prime}_{i_{2}}-x^{\prime}_{i_{1}}x_{i_{2}}$ (53)
 given by the minors of $\mathbf{\Phi}$; and for $r\geq 3$ we define linear polynomials
-$s^{(j)}_{r;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z^{(j)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z^{(j-1)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
-where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
+$s_{r,j;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z_{r-1,j;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z_{r-1,j-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
+where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z_{r,j;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
 $\mathbf{s}_{r}(\mathbf{s}_{r-1})=\mathbf{0}.$ (55)
@@
 Either express (46) in the standard bases of $F=R^{f}$ and $G=R^{2}$.
-Or more directly, consider the $\mathbf{Z}_{r}$ and $\mathbf{s}_{r}$ as $(r-1)$-tuples of (exterior) $r$-vectors, observe that (53) means $\mathbf{s}^{(1)}_{2}=\mathbf{x}\wedge\mathbf{x}^{\prime}$, that (54) means $\mathbf{s}^{(j)}_{r}=\mathbf{x}\wedge\mathbf{Z}^{(j)}_{r-1}+\mathbf{x}^{\prime}\wedge\mathbf{Z}^{(j-1)}_{r-1}$, so the alternating property gives $\mathbf{s}^{(1)}_{3}(\mathbf{s}_{2})=\mathbf{x}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, $\;\mathbf{s}^{(2)}_{3}(\mathbf{s}_{2})=\mathbf{x}^{\prime}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, and for $r\geq 4$:
+Or more directly, consider the $\mathbf{Z}_{r}$ and $\mathbf{s}_{r}$ as (r1)(r-1)-tuples of (exterior) $r$-vectors, observe that (53) means $\mathbf{s}_{2,1}=\mathbf{x}\wedge\mathbf{x}^{\prime}$, that (54) means $\mathbf{s}_{r,j}=\mathbf{x}\wedge\mathbf{Z}_{r-1,j}+\mathbf{x}^{\prime}\wedge\mathbf{Z}_{r-1,j-1}$, so the alternating property gives $\mathbf{s}_{3,1}(\mathbf{s}_{2})=\mathbf{x}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, $\mathbf{s}_{3,2}(\mathbf{s}_{2})=\mathbf{x}^{\prime}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, and for $r\geq 4$:
 […]
@@
 Let $\mathbf{G}_{\mathcal{H}}$ and $\mathbf{G}$ be the corresponding generating matrices of $\mathsf{C}$$_{\mathcal{H}}$ and $\mathsf{C}$}.
-Let also $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the polynomial ring in $k$ indeterminates, and $S_{\mathcal{H}}=\mathbb{F}[X_{1},\dots,X_{k-1}]$ its subring in the first $k-1$ indeterminates.
+Let also $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the polynomial ring in $k$ indeterminates, and $S_{\mathcal{H}}=\mathbb{F}[X_{1},\dots,X_{k-1}]$ its subring in the first k1k-1 indeterminates.
 We have a commutative diagram
@@
 elements of $S_{1}$.
-Recall from [10, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero.
+Recall from [13, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero.
 Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
 Joint with (60) this gives $f_{\mathcal{H}}=\dim(V_{\mathbf{\Phi},\mathcal{H}})\geq\dim(V_{\mathbf{\Phi}})-(d-1)=f-d+1$.
-Application to alternant and Goppa codes
+Application to alternant codes
 Consider a code ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$, assumed to be proper.
-By (21), after extension of scalars, ${\mathsf{C}}_{\mathbb{F}_{q^{m}}}$ admits as a basis the $k=mt$ vectors $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$ for $10\leq a\leq t-$ 1 and $10\leq u\leq m-$ 1.
+By (21), after extension of scalars, ${\mathsf{C}}_{\mathbb{F}_{q^{m}}}$ admits as a basis the $k=mt$ vectors $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$ for $0\leq a\leq t-1$ and $0\leq u\leq m-1$.
 Rename the $k=mt$ variables in our polynomial ring $S$ accordingly, so that our evaluation map now is
@@
 Proof
-The $2\times 2$ minor defined by the columns $\left[\begin{array}[]{c}X^{(e-u)}_{a}\\ X^{(e-u)}_{q^{u}+a}\end{array}\right]$ of $\mathbf{B}^{(e-u)}_{u}$ and $\left[\begin{array}[]{c}X^{(e-v)}_{b}\\ X^{(e-v)}_{q^{v}+b}\end{array}\right]$ of $\mathbf{B}^{(e-v)}_{v}$ is
-$X^{(e-u)}_{a}X^{(e-v)}_{q^{v}+b}-X^{(e-u)}_{q^{u}+a}X^{(e-v)}_{b}$ (66)
+The $2\times 2$ minor defined by the columns $\left[\begin{array}[]{c}X^{(e-u)}_{a}\\ X^{(e-u)}_{a+q^{u}}\end{array}\right$] of $\mathbf{B}^{(e-u)}_{u}$ and $\left[\begin{array}[]{c}X^{(e-v)}_{b}\\ X^{(e-v)}_{b+q^{v}}\end{array}\right$] of $\mathbf{B}^{(e-v)}_{v}$ is
+$X^{(e-u)}_{a}X^{(e-v)}_{b+q^{v}}-X^{(e-v)}_{b}X^{(e-u)}_{a+q^{u}}$ (66)
 which evaluates under (62) to
-$(\mathbf{y}\mathbf{x}^{a})^{q^{e-u}}(\mathbf{y}\mathbf{x}^{q^{v}+b})^{q^{e-v}}-(\mathbf{y}\mathbf{x}^{q^{u}+a})^{q^{e-u}}(\mathbf{y}\mathbf{x}^{b})^{q^{e-v}}=\mathbf{0}.$ (67)
+$(\mathbf{y}\mathbf{x}^{a})^{q^{e-u}}(\mathbf{y}\mathbf{x}^{b+q^{v}})^{q^{e-v}}-(\mathbf{y}\mathbf{x}^{b})^{q^{e-v}}(\mathbf{y}\mathbf{x}^{a+q^{u}})^{q^{e-u}}=\mathbf{0}.$ (67)
 Lemma 8
@@
 By Proposition 4 applied $s$ times (with $g=d=2$), $I_{2}({\mathsf{C}}_{s})$ then contains the $2\times 2$ minors of a $2\times f_{s}$ matrix $\mathbf{\Phi}_{s}$, where $f_{s}\geq f-s$, and this matrix $\mathbf{\Phi}_{s}$ is $1$-generic also.
-Then by [10, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$.
+Then by [13, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$.
 We then conclude with Corollary 2 (with $l=1$ and $V^{(1)}=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi}_{s})$).
-Remark 1
+Further results on syzygies of dual alternant codes
 It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$.
-See the corresponding discussion at the beginning of the Supplementary material.
-Examples show the bound (68) on $r_{\max}$ is tight in general for alternant codes.
-However one can improve it in the Goppa case.
+Unfortunately we only have very partial results in this direction.
+These partial results will not be used in the remainder of the text, but we present them nevertheless in the hope that they could motivate further study.
+Let us first consider the simpler case $s=0$, i.e. without shortening. Let then $\mathsf{C}$} be a proper dual alternant code as in Theorem 3.1.
+The Eagon-Northcott complex of each conjugate $\mathbf{\Phi}^{(u)}$ is a subcomplex of the minimal resolution of $\mathsf{C}$}, so that in each degree $r$ we get a subspace
+$V_{r}=M_{r-1,r}(\mathbf{\Phi}^{(0)})+\cdots+M_{r-1,r}(\mathbf{\Phi}^{(m-1)})\subseteq M_{r-1,r}({\mathsf{C}}),$ (69)
+hence $\beta_{r-1,r}({\mathsf{C}})\geq\dim V_{r}$. Two questions then arise naturally:
+  • (Q1)
+    Is the sum in (69) direct? If not, what are the linear dependencies between the $M_{r-1,r}(\mathbf{\Phi}^{(u)})$ in $M_{r-1,r}({\mathsf{C}})$?
+  • (Q2)
+    Is the inclusion $V_{r}\subseteq M_{r-1,r}({\mathsf{C}})$ an equality, i.e. do the $M_{r-1,r}(\mathbf{\Phi}^{(u)})$ together generate $M_{r-1,r}({\mathsf{C}})$?
+We can answer these questions in degree $r=2$.
+Definition 5
+For any $u,v$ we let
+$I_{2}^{(u,v)}\subseteq I_{2}({\mathsf{C}})$ (70)
+be the subspace generated by the quadratic forms
+$q^{(u,v)}_{a,b,c,d}=X_{a}^{(u)}X_{d}^{(v)}-X_{b}^{(v)}X_{c}^{(u)}$ (71)
+for all $a,b,c,d$ such that
+$aq^{u}+dq^{v}=bq^{v}+cq^{u}\mod q^{m}-1$ (72)
+(so that these quadratic forms indeed evaluate to $\mathbf{0}$).
+Observe that $I_{2}^{(u,v)}=I_{2}^{(v,u)}$.
+Thus, in order to avoid these $I_{2}^{(u,v)}$ appearing twice, we will impose some conditions on the pairs $u,v$ we consider.
+Say, we consider only those where $v=u-w\mod m$ with $0\leq w<m/2$, and also in case $m$ even those where $v=u-m/2\mod m$ with $0\leq u<m/2$.
+Lemma 9
+With such a restriction, the $I_{2}^{(u,v)}$ are in direct sum in $I_{2}({\mathsf{C}})$.
+Proof
+Consider a linear combination $\sum_{u,v}q^{(u,v)}$ for general $q^{(u,v)}\in I_{2}^{(u,v)}$.
+If some $q^{(u,v)}$ is nonzero, then it contains a monomial $X_{i}^{(u)}X_{j}^{(v)}$ for some $i,j$.
+However this monomial does not appear in the $q^{(u^{\prime},v^{\prime})}$ with $(u^{\prime},v^{\prime})\neq(u,v)$, so the linear combination is nonzero.
+Lemma 10
+If $v=u-w\mod m$ with $1\leq w\leq m/2$, then a basis of $I_{2}^{(u,v)}$ is given by those $2\times 2$ minors of the matrix
+$\left(\begin{array}[]{c|c}\mathbf{B}^{(u)}_{0}&\mathbf{B}^{(v)}_{w}\end{array}\right)$
+obtained by picking one column in the first block $\mathbf{B}^{(u)}_{0}$ and the other column in the second block $\mathbf{B}^{(v)}_{w}$.
+As a consequence, $\dim I_{2}^{(u,v)}=(t-1)(t-q^{w})$ if $w\leq e$, and 0 otherwise.
+Likewise if $u=v$, a basis of $I_{2}^{(u,u)}$ is given by the $2\times 2$ minors of $\mathbf{B}^{(u)}_{0}$, and $\dim I_{2}^{(u,v)}=\frac{(t-1)(t-2)}{2}$.
+Proof
+First consider the case $v=u-w\mod m$ with $1\leq w\leq m/2$. Let
+$q^{(u,v)}_{a,b}=X^{(u)}_{a}X^{(v)}_{b+q^{w}}-X^{(v)}_{b}X^{(u)}_{a+1}$ (73)
+be the $2\times 2$ minor constructed from the column $\left[\begin{array}[]{c}X^{(u)}_{a}\\ X^{(u)}_{a+1}\end{array}\right$] of $\mathbf{B}^{(u)}_{0}$ and the column $\left[\begin{array}[]{c}X^{(v)}_{b}\\ X^{(v)}_{b+q^{w}}\end{array}\right$] of $\mathbf{B}^{(v)}_{w}$.
+Then keeping the notation of (71) we have $q^{(u,v)}_{a,b}=q^{(u,v)}_{a,b,c,d}$ with $c=a+1$ and $d=b+q^{w}$, so condition (72) is satisfied and $q^{(u,v)}_{a,b}\in I_{2}^{(u,v)}$.
+Now we show that these $q^{(u,v)}_{a,b}$ generate $I_{2}^{(u,v)}$.
+For this it suffices to show that all the $q^{(u,v)}_{a,b,c,d}$ (for $a,b,c,d$ satisfying (72)) are linear combinations of the $q^{(u,v)}_{a,b}$.
+After a possible change of sign we can assume $c>a$.
+Then we proceed by induction on the difference cac-a.
+First, in the case $c=a+1$, we necessarily have $d=b+q^{w}$, and $q^{(u,v)}_{a,b,c,d}=q^{(u,v)}_{a,b}$.
+Now for general $c>a$ we write
+$q^{(u,v)}_{a,b,c,d}=q^{(u,v)}_{a,b+q^{w},c-1,d}+q^{(u,v)}_{c-1,b}$ (74)
+and apply the induction hypothesis to $q^{(u,v)}_{a,b+q^{w},c-1,d}$.
+Last, linear independence of the $q^{(u,v)}_{a,b}$ follows from the theory of $1$-genericity, but it can also be proved directly as follows.
+Consider a linear combination $\sum_{a,b}\lambda_{a,b}q^{(u,v)}_{a,b}$, and choose a minimal $a$ such that $\lambda_{a,b}\neq 0$.
+Then $q^{(u,v)}_{a,b}$ contains the monomial $X^{(u)}_{a}X^{(v)}_{b+q^{w}}$, but this monomial does not appear in the other $q^{(u,v)}_{a^{\prime},b^{\prime}}$ with $\lambda_{a^{\prime},b^{\prime}}\neq 0$ (indeed, either $a^{\prime}=a$ but then $b^{\prime}\neq b$, or $a^{\prime}\geq a+1$ by minimality).
+Thus the linear combination is nonzero.
+The proof of the case $u=v$ is entirely similar, details are left to the reader.
+Now we translate these results in terms of the $2\times 2$ minors of the conjugate matrices $\mathbf{\Phi}^{(u)}$.
+For $1\leq i_{1}<i_{2}\leq f$ we denote by $q^{(u)}_{i_{1},i_{2}}$ the minor defined by the $i_{1}$-th and $i_{2}$-th columns of $\mathbf{\Phi}^{(u)}=\left(\begin{array}[]{c|c|c|c}\mathbf{B}^{(u+e)}_{0}&\mathbf{B}^{(u+e-1)}_{1}&\cdots&\mathbf{B}^{(u)}_{e}\end{array}\right)$.
+Observe that if $i_{1}$ falls in the block $\mathbf{B}^{(u+e-j_{1})}_{j_{1}}$ of $\mathbf{\Phi}^{(u)}$, and $i_{2}$ in the block $\mathbf{B}^{(u+e-j_{2})}_{j_{2}}$, then
+$q^{(u)}_{i_{1},i_{2}}\in I_{2}^{(u+e-j_{1},u+e-j_{2})}.$ (75)
+It follows:
+Proposition 5
+For a given $u$, the minors $q^{(u)}_{i_{1},i_{2}}$ with first column in $\mathbf{B}^{(u+e)}_{0}$ (i.e.
+with $j_{1}=0$, or equivalently $i_{1}\leq t-1$) are linearly independent.
+Actually, they form a basis of the “primitive” subspace
+$I_{2}(\mathbf{\Phi}^{(u)})_{\operatorname{prim}}=I_{2}^{(u+e,u+e)}\oplus I_{2}^{(u+e,u+e-1)}\oplus\cdots\oplus I_{2}^{(u+e,u)}\subseteq I_{2}(\mathbf{\Phi}^{(u)}),$ (76)
+of dimension $\dim I_{2}(\mathbf{\Phi}^{(u)})_{\operatorname{prim}}=\binom{f}{2}-\binom{f-(t-1)}{2}$.
+Proof
+Direct consequence of Lemmata 9 and 10. For the dimension, we count $\binom{f}{2}$ pairs $i_{1},i_{2}$ with $1\leq i_{1}<i_{2}\leq f$, and we exclude the $\binom{f-(t-1)}{2}$ of them with $i_{1}>t-1$.
+Proposition 6
+With notation as above, the minors $q^{(u)}_{i_{1},i_{2}}$ with first column not in $\mathbf{B}^{(u+e)}_{0}$ (i.e.
+with $j_{1}>0$, or equivalently, $i_{1}>t-1$) belong to the intersection $I_{2}(\mathbf{\Phi}^{(u)})\cap I_{2}(\mathbf{\Phi}^{(u-j_{1})})_{\operatorname{prim}}$.
+Proof
+Actually, $q^{(u)}_{i_{1},i_{2}}\in I_{2}(\mathbf{\Phi}^{(u)})\cap I_{2}^{(u+e-j_{1},u+e-j_{2})}$.
+We can now answer Question (Q1) for $r=2$:
+Corollary 4
+If $e<m/2$, the sum $V_{2}=I_{2}(\mathbf{\Phi}^{(0)})+\cdots+I_{2}(\mathbf{\Phi}^{(m-1)})$ can be written as a direct sum
+$V_{2}=I_{2}(\mathbf{\Phi}^{(0)})_{\operatorname{prim}}\oplus\cdots\oplus I_{2}(\mathbf{\Phi}^{(m-1)})_{\operatorname{prim}}$ (77)
+so
+$\begin{split}\beta_{1,2}({\mathsf{C}})\geq\dim V_{2}&=m\left(\binom{f}{2}-\binom{f-(t-1)}{2}\right)\\ &=\frac{m(t-1)}{2}\left((2e+1)t-2\frac{q^{e+1}-1}{q-1}\right).\end{split}$ (78)
+A basis of $V_{2}$ is given by the quadratic forms $q^{(u)}_{i_{1},i_{2}}$ with $u,i_{1},i_{2}$ ranging among those with $0\leq u\leq m-1$, $1\leq i_{1}<i_{2}\leq f$, and $i_{1}\leq t-1$.
+Proof
+Follows by the two Propositions above, and Lemma 9 again. The condition $e<m/2$ is there to ensure that no $I_{2}^{(u,v)}$ appears twice when we expand the sum.
+The lower bound (78) was first observed without proof as [15, Exp.
+fact 1], and was eventually proved as [30, Th.
+20].
+Actually, [15] makes computations with quadratic forms very similar to ours, but what lacks there is a proof of linear independence (or said otherwise, in order to prove linear independence, the authors have to assume that (78) holds and is an equality).
+On the other hand, the proof in [30] uses a different approach: the authors do not work with the space of quadratic relations (the kernel of the degree $2$ evaluation map), but with the square code (the image of this evaluation map).
+So what we just did is provide an alternative proof for [30, Th.
+20], while reverting to the original approach of [15].
+As for Question (Q2), it is experimentally observed in [15] that (78) is an equality with overwhelming probability, so for $r=2$ the answer to (Q2) is positive in general.
+Now we turn to the case $r\geq 3$, where we can only propose conjectures.
+Proposition 3 gives a basis $s_{r,j;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ for each syzygy space $M_{r-1,r}(\mathbf{\Phi}^{(u)})$, where the indices $i_{1},\dots,i_{r}$ correspond to columns of $\mathbf{\Phi}^{(u)}$.
+A natural generalization of the answer to (Q1) would be as follows:
+  • For a given $u$, the $s_{r,j;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ with column $i_{1}$ in $\mathbf{B}^{(e+u)}_{0}$ are linearly independent, hence form a basis of a subspace $M_{r-1,r}(\mathbf{\Phi}^{(u)})_{\operatorname{prim}}\subseteq M_{r-1,r}(\mathbf{\Phi}^{(u)})$, of dimension $\binom{f}{r}-\binom{f-(t-1)}{r}$.
+  • On the other hand, those with $i_{1}$ not in $\mathbf{B}^{(e+u)}_{0}$ belong to $M_{r-1,r}(\mathbf{\Phi}^{(u)})\cap\sum_{u^{\prime}\neq u}M_{r-1,r}(\mathbf{\Phi}^{(u^{\prime})})_{\operatorname{prim}}$.
+  • If $e<m/2$, the sum $V_{r}=M_{r-1,r}(\mathbf{\Phi}^{(0)})+\cdots+M_{r-1,r}(\mathbf{\Phi}^{(m-1)})$ can be written as the direct sum of the $M_{r-1,r}(\mathbf{\Phi}^{(u)})_{\operatorname{prim}}$.
+Hence
+    $\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{r})=m(r-1)\left(\binom{f}{r}-\binom{f-(t-1)}{r}\right)$ (79)
+    and in particular for r>f(t1)r>f-(t-1):
+    $\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{r})=m(r-1)\binom{f}{r}$. (80)
+Although most of the arguments used to prove the case $r=2$ seem to generalize to $r\geq 3$, the linear algebra manipulations become more and more intricate, and the author was unfortunately not able to derive a satisfactory proof.
+On the other hand, we will see examples where (80) not only holds but is an equality, which suggests that the answer to (Q2) can be positive in some cases.
+However the following example shows that for $r=3$ the answer to (Q2) is negative in general, and the lower bound (79) still isn’t tight:
+Example 1
+With $\mathsf{C}$} as above, $I_{2}({\mathsf{C}})$ contains the four quadratic forms
+$X_{b+1}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}+1}^{(u)}X_{c}^{(u+v)}$ (81)
+$X_{b}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}}^{(u)}X_{c}^{(u+v)}$ (82)
+$X_{a}X_{b+q^{v}+1}^{(u)}-X_{a+q^{u}}X_{b+q^{v}}^{(u)}$ (83)
+$X_{a}X_{b+1}^{(u)}-X_{a+q^{u}}X_{b}^{(u)}$ (84)
+which admit the syzygy
+$\begin{split}X_{a}(&X_{b+1}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}+1}^{(u)}X_{c}^{(u+v)})-X_{a+q^{u}}(X_{b}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}}^{(u)}X_{c}^{(u+v)})+\\ &+X_{c}^{(u+v)}(X_{a}X_{b+q^{v}+1}^{(u)}-X_{a+q^{u}}X_{b+q^{v}}^{(u)})-X_{c+1}^{(u+v)}(X_{a}X_{b+1}^{(u)}-X_{a+q^{u}}X_{b}^{(u)})=0.\end{split}$ (85)
+However this syzygy defines an element of $M_{2,3}({\mathsf{C}})$ that does not belong to $V_{3}$ in general.
+For general $s$, the author only has guesses motivated by experimental evidence. More precisely, in many examples, one finds that the lower bound (68) is an equality, i.e.
+$r_{\max}({\mathsf{C}}_{s})=f-s$ (86)
+and then also that the natural generalization of (80) holds and is an equality, i.e.
+$\beta_{r-1,r}({\mathsf{C}}_{s})=m(r-1)\binom{f-s}{r}$ (87)
+for $r$ close enough to fsf-s.
+Figures 5 and 5 illustrate these facts. Boldface values match (87) (and (86)).
+[…]
+Figure 4: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{2,10,5}\;$ ($e\!=\!2,r^{*}\!=\!8$)
+[…]
+This strongly suggests that the minimal resolution of $\mathsf{C}$} (resp.
+its shortened subcodes ${\mathsf{C}}_{s}$) contains $m$ conjugate Eagon-Northcott subcomplexes of length $f$ (resp.
+fsf-s), and that in high enough degree this minimal resolution actually coincides with the direct sum of these conjugate subcomplexes.
+Improved lower bounds in the Goppa case
+We just saw that the bound (68) on $r_{\max}$ is tight in general for alternant codes.
+However one can improve it in the Goppa case, still using the same techniques.
 Let us focus on $q=2$ for simplicity.
@@
 $L_{g}\,+\,\varphi(L_{g})\;=\;L_{g^{2}},$ (88)
-the sum being direct
+the sum being direct.
 Proof
@@
 To conclude we only have to prove $L_{g}\cap\varphi(L_{g})=0$.
-However we have $F(X)\in L_{g}\cap\varphi(L_{g})$ if and only if $F(X)=\frac{A(X)}{g(X)}=\frac{B(X)^{2}}{g(X)^{2}}$ for some $A,B$ of degree $<t$.
+However we have $F(X)\in L_{g}\cap\varphi(L_{g})$ if and only if $F(X)=\frac{A(X)}{g(X)}=\frac{B(X)^{2}}{g(X)^{2}}$ for some $A,B$ of degree <t<t.
 But then this implies $g(X)|B(X)^{2}$ with $g$ squarefree of degree $t$, which is impossible unless $B=0$.
@@
 Same as Theorem 3.1, with $\mathbf{\Phi}$ adapted to fit Proposition 7.
+Remark 1
+It turns out that this last lower bound can be improved still further, replacing the value of $\widehat{f}$ with
+$r^{*}=(a+2)t-2^{a}$ (93)
+where $a=\lceil\log_{2}(t)\rceil$.
+The proof, which will be found in the forthcoming paper [35], uses new ingredients, the main of which being a reinterpretation of the number of columns of a convenient matrix $\mathbf{\Phi}$ as $r^{*}=\dim({\mathsf{C}}_{\mathbb{F}_{2^{m}}}\cap\mathbf{x}{\mathsf{C}}_{\mathbb{F}_{2^{m}}})$.
+(A similar result also holds in the nonbinary case, but then we have to set $r^{*}=(a+1)t-q^{a}$ where $a=\lceil\log_{q}(t/(q-1))\rceil$.
+Observe that the binary case does specialize from the general case with $q=2$, as the $a+1$ becomes a $a+2$.)
+This new bound now appears to be tight in most examples, i.e.
+$r_{\max}({\mathsf{C}}_{s})=r^{*}-s$ (94)
+with $r^{*}$ given by Remark 1. Moreover, the analogue of (87) then also appears to hold experimentally, i.e.
+$\beta_{r-1,r}({\mathsf{C}}_{s})=m(r-1)\binom{r^{*}-s}{r}$ (95)
+for $r$ close enough to $r^{*}-s$.
+Figures 8-8 illustrate these facts. Boldface values match (95) (and (94)), leading to the same presumption that the one following Figures 5 and 5.
+[…]
+Figure 6: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,8,5}$ ($a\!=\!3,r^{*}\!=\!17$)
+[…]
+Figure 7: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}$ ($a\!=\!4,r^{*}\!=\!38$)
+[…]
+Figure 8: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,10}$ ($a\!=\!4,r^{*}\!=\!44$)
+Observe that Theorem 3.1 and Corollary 5 are not tight on these examples:
+  • for $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,8,5}$ they give $f=8$ and $\widetilde{f}=15$, compared to $r^{*}=17$
+  • for $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}$ they give $f=21$ and $\widetilde{f}=33$, compared to $r^{*}=38$
+  • for $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,10}$ they give $f=25$ and $\widetilde{f}=39$, compared to $r^{*}=44$
+so Remark 1 really is an improvement.
 4 Regularity $2$ and the small defect heuristic
@@
 Proof
-Generating series reformulation of [10, Cor. 1.10].
-Definition 5 ([26, Def. 1.5 & Th. 2.35])
+Generating series reformulation of [13, Cor. 1.10].
+Definition 6([36, Def. 1.5 & Th. 2.35])
 The Castelnuovo-Mumford regularity of a projective $[n,k]$-code $\mathsf{C}$} is the smallest integer $r$ such that ${\mathsf{C}}^{\langle r\rangle}=\mathbb{F}^{n}$.
-Definition 6 (cf. [10, §4A], after [23, Lect. 14])
+Definition 7(cf. [13, §4A], after [31, Lect. 14])
 The Castelnuovo-Mumford regularity of $\mathsf{C}$$^{\langle\cdot\rangle}$ is $\max\{r:\>\exists i,\,\beta_{i,i+r}({\mathsf{C}})>0\}$.
-Proposition 7 ([10, Th. 4.2])
+Proposition 9([13, Th. 4.2])
 These two definitions coincide.
@@
 The square code distinguisher, and the filtration attack from [2] that extends it, work for codes $\mathsf{C}$} with ${\mathsf{C}}^{\langle 2\rangle}\varsubsetneq\mathbb{F}^{n}$, i.e.
-of regularity $>2$.
+of regularity >2>2.
 This means that codes of regularity $2$ are hard to deal with under this approach.
@@
 Then $H_{{\mathsf{C}}}(z)=1+kz+n\frac{z^{2}}{1-z}$, so $B(z)=(1+kz)(1-z)^{k}+nz^{2}(1-z)^{k-1}$ by Proposition 8, thus $B_{r}=(-1)^{r-1}(r-1)\binom{k+1}{r}+(-1)^{r}n\binom{k-1}{r-2}=(-1)^{r-1}\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$.
-On the other hand, ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$ also means ${\mathsf{C}}$ has regularity $2$ in the sense of Definition 6, so (74) reduces to $B_{r}=(-1)^{r-1}\beta_{r-1,r}+(-1)^{r-2}\beta_{r-2,r}=(-1)^{r-1}\operatorname{ind}(\varphi_{r})$ and we conclude.
+On the other hand, ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$ also means $\mathsf{C}$} has regularity 22 in the sense of Definition 7, so (96) reduces to $B_{r}=(-1)^{r-1}\beta_{r-1,r}+(-1)^{r-2}\beta_{r-2,r}=(-1)^{r-1}\operatorname{ind}(\varphi_{r})$ and we conclude.
 Corollary 6
@@
 Proof
-The minimal resolution of ${\mathsf{C}}$ has length $k-1$, so $\beta_{i,j}=0$ for $i>k-1$, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
-Regularity $2$ helps in computing the full Betti diagram of codes. See Propositions 10 and 11 in the Supplementary material for more.
+The minimal resolution of $\mathsf{C}$} has length k1k-1, so $\beta_{i,j}=0$ for i>k1i>k-1, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
+Regularity $2$ helps in computing the full Betti diagram of codes, as one row can be deduced from the other.
+In order to illustrate this, let us first recall that a $[n,k]$-code $\mathsf{C}$} is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
+Lemma 13([19, Th. 1], reformulated)
+Let $\mathsf{C}$} be a $[n,k]$-code with $n\leq 2k-1$. Assume $\mathsf{C}$} is MDS. Then for all $r\leq 2k+1-n$ we have $\beta_{r-2,r}({\mathsf{C}})=0$.
+Proposition 10
+Let $\mathsf{C}$} be a $[k+1,k]$ MDS code, for instance a parity code or a $[k+1,k]$ $\operatorname{{\mathsf{GRS}}}$ code.
+Then the nonzero Betti numbers of $\mathsf{C}$} are $\beta_{0,0}=\beta_{k-1,k+1}=1$, and
+$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (107)
+for $2\leq r\leq k-1$. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
+Proof
+The parameters imply that $\mathsf{C}$} has regularity $2$.
+By Lemma 13 we have $\beta_{r-2,r}=0$ for $r\leq k$.
+Then $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-(k+1)\right)\binom{k-1}{r-2}$ and we conclude with a straightforward calculation.
+Proposition 11
+Let $\mathsf{C}$} be a [2k1,k][2k-1,k] $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of $\mathsf{C}$} are:
+  • $\beta_{0,0}=1$
+  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $2\leq r\leq k-1$
+  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
+In particular the ideal $I({\mathsf{C}})$ is generated by $\beta_{1,2}=\frac{(k-1)(k-2)}{2}$ quadratic forms and $\beta_{1,3}=k-1$ cubic forms.
+Proof
+Using the basis $\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}$ of $\mathsf{C}$}, we see $I_{2}({\mathsf{C}})$ contains the $2\times 2$ minors of $\mathbf{\Phi}=\left(\begin{array}[]{cccc}X_{0}&X_{1}&\dots&X_{k-2}\\ X_{1}&X_{2}&\dots&X_{k-1}\end{array}\right)$, i.e.
+$I_{2}({\mathsf{C}})\supseteq\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ of dimension $\binom{k-1}{2}$.
+Now set n=2k1n=2k-1.
+Because $n\geq 2k-1$, we have $\dim{\mathsf{C}}^{\langle 2\rangle}=2k-1$, hence $\dim I_{2}({\mathsf{C}})=\binom{k-1}{2}$, so $I_{2}({\mathsf{C}})=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$.
+Corollary 1 and the Eagon-Northcott complex of $\mathbf{\Phi}$ then give $\beta_{r-1,r}=(r-1)\binom{k-1}{r}$.
+Because $n\leq 2k-1$, we have ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$.
+Then $\beta_{r-2,r}=\beta_{r-1,r}-\operatorname{ind}(\varphi_{r})=(r-1)\binom{k-1}{r}-\left(\frac{k(k+1)}{r}-(2k-1)\right)\binom{k-1}{r-2}$, and a straightforward calculation allows to conclude.
+Figures 10 and 10 illustrate Propositions 10 and $1$.
+$\begin{array}[]{c|cccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&27&105&189&189&105&27&-\\ 2&-&-&-&-&-&-&-&1\end{array}$
+Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
+$\begin{array}[]{c|ccccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&21&70&105&84&35&6&-\\ 2&-&7&42&105&140&105&42&7\end{array}$
+Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
+Exercise 1
+Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for k+1<n<2k1k+1<n<2k-1.
 The small defect heuristic
-By (78) and Theorem 4.1, for any $[n,k]$-code ${\mathsf{C}}$ of regularity $2$ and for any $r\geq 2$ we have
-$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}.$ (85)
-We might interpret this lower bound as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$.
+Because of (103), a necessary condition for a $[n,k]$-code $\mathsf{C}$} of regularity $2$ to have $\beta_{r-1,r}=0$ is that $\operatorname{ind}(\varphi_{r})\leq 0$. This immediately yields
+$r_{\max}({\mathsf{C}})\geq\lceil r_{\operatorname{crit}}({\mathsf{C}})\rceil-1$ (108)
+where we define the critical degree of $\mathsf{C}$}
+$r_{\operatorname{crit}}({\mathsf{C}})=\frac{k(k+1)}{n}$ (109)
+as the value of $r$ at which the formula (105) for $\operatorname{ind}(\varphi_{r})$ changes sign.
+More precisely, (100) and Theorem 4.1 give the lower bound
+$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$ (110)
+that can be interpreted as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$.
 Then, when $\mathsf{C}$} is random, it is quite natural to expect that no other syzygy will appear, or equivalently that (110) will be an equality, i.e.
@@
 Another argument is the minimal resolution conjecture of [25].
-It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$ or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$.
+It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq r_{\operatorname{crit}}({\mathsf{C}})$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq r_{\operatorname{crit}}({\mathsf{C}})$, or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$.
 However, two points require our attention:
@@
 This makes it desirable to investigate links between Betti numbers and distance properties of a code.
+At this moment the author mostly has only experimental results in this direction: Actually, experiments suggest much stronger results hold.
 For any code $\mathsf{C}$}, let us denote by $A_{i}({\mathsf{C}})$ the number of codewords of Hamming weight $i$ in $\mathsf{C}$}.
+Also recall $r_{\operatorname{crit}}=\frac{k(k+1)}{n}$.
 Experimental fact 1
-Let ${\mathsf{C}}$ be a $[n,k]_{q}$-code of regularity $2$, with minimum distance $d=\operatorname{d}_{\min}({\mathsf{C}})$ and dual minimum distance $d^{\perp}=\operatorname{d}_{\min}({\mathsf{C}}^{\perp})$.
+Let $\mathsf{C}$} be a $n,k]_{q}-$code of regularity $2$, with minimum distance $d=\operatorname{d}_{\min}({\mathsf{C}})$ and dual minimum distance $d^{\perp}=\operatorname{d}_{\min}({\mathsf{C}}^{\perp})$.
   1. 1.
     For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$.
-    Moreover if $d^{\perp}\leq\frac{k(k+1)}{n}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
+    Moreover if $d^{\perp}\leq r_{\operatorname{crit}}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of $\mathsf{C}$} where this inequality is an equality).
   2. 2.
     Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$.
-    Moreover if $k+1-d\geq\frac{k(k+1)}{n}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
-It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and for $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$, when applicable (i.e. when these intervals are nonempty).
+    Moreover if $k+1-d\geq r_{\operatorname{crit}}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of $\mathsf{C}$} where this inequality is an equality).
+It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq r_{\operatorname{crit}}$ and for $r_{\operatorname{crit}}\leq\ r\leq k+1-d$, when applicable (i.e.
+when these intervals are nonempty).
 Experimental fact 2
-Conversely, for a random code ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$.
-In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
-Figures 13-14 in the Supplementary material illustrate various aspects of these Experimental facts.
-Consequently, we postulate:
+Conversely, for a random code $\mathsf{C}$} among codes of given parameters $n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min(d^{\perp},r_{\operatorname{crit}})$ and as $r\gg\max(r_{\operatorname{crit}},k+1-d)$.
+In particular $r_{\max}({\mathsf{C}})=\max(\lfloor r_{\operatorname{crit}}\rfloor,k+1-d)$, or is very close to this value, with high probability.
+Actually, as a modest first step, the author is able to prove the first assertion in point 2. of Experimental fact 1:
+Proposition 12
+Any $[n,k,d]$ code $\mathsf{C}$} has $r_{\max}({\mathsf{C}})\geq k+1-d$, i.e. $\beta_{r-1,r}({\mathsf{C}})>0$ for all $r\leq k+1-d$.
+Proof
+Let $\mathbf{c}_{1}$ be a codeword of minimum weight $\operatorname{w}(\mathbf{c}_{1})=d$.
+Let $S$ be the support of $\mathbf{c}_{1}$, and let ${\mathsf{C}}_{S}\subseteq{\mathsf{C}}$ be the subcode made of all codewords that vanish over $S$.
+Then $\dim({\mathsf{C}}_{S})\geq k-\#{S}=k-d$, so we can find kdk-d linearly independent codewords $\mathbf{c}_{2},\dots,\mathbf{c}_{k+1-d}$ in ${\mathsf{C}}_{S}$.
+Let then complete $\mathbf{c}_{1},\dots,\mathbf{c}_{k-d+1}$ into a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of $\mathsf{C}$}, and use this basis to define the graded evaluation map $S=\mathbb{F}[X_{1},\dots,X_{k}]\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$.
+By construction we have $\mathbf{c}_{1}\mathbf{c}_{2}=\cdots=\mathbf{c}_{1}\mathbf{c}_{k-d+1}=\mathbf{0}$, which means $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle\subseteq I_{2}({\mathsf{C}})$.
+By Corollary 2 we deduce $\beta_{r-1,r}({\mathsf{C}})\geq\beta_{r-1,r}(S/\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle)$ for all $r$, hence in particular $r_{\max}({\mathsf{C}})\geq r_{\max}(S/\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle)$.
+However we have $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle=X_{1}\langle X_{2},\dots,X_{k+1-d}\rangle$, so the minimal resolution of $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle$ can be identified with the Koszul complex on $X_{2},\dots,X_{k+1-d}$ (with degree shifted by $1$), so its $r_{\max}$ is k+1dk+1-d.
+Remark 2
+The quantity $r_{\max}=\max\{r:\;\beta_{r-1,r}>0\}$, and the dual quantity $r_{\min}=\min\{r:\;\beta_{r-2,r}>0\}$ that implicitely appears together with it in these Experimental facts, are at the core of deep conjectures of Green and Green-Lazarsfeld, discussed e.g.
+in the last two chapters of [13] (with the notation therein, $r_{\max}=b(X)$ and $r_{\min}=a(X)+3$).
+Let us now illustrate various aspects of these Experimental facts.
+Figures 13-13 compare Betti diagrams for three codes of length $n=23$ and dimension $k=12$ (so $r_{\operatorname{crit}}\approx 6.78$) of quite different nature.
+First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e.
+with all defects equal to zero.
+It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram.
+Such codes are MDS, i.e.
+they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq r_{\operatorname{crit}}$ and $r_{\operatorname{crit}}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified.
+Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g.
+$q=5$ suffices (the author did not search for optimality).
+These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified.
+Moreover, the converse Experimental fact 2 is also verified as all defects are zero.
+[…]
+Figure 11: an idealized $[23,12]$-code according to the minimal resolution conjecture
+[…]
+Figure 12: a (pseudo)random $23,12]_{2}-$code ($d=3$, $d^{\perp}=4$)
+[…]
+Figure 13: the $23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
+Second, Figure 13 presents the Betti diagram of the binary code whose generator matrix
+$G=\left(\begin{array}[]{ccccccccccccccccccccccc}1&1&0&0&1&0&0&1&0&0&0&0&1&1&1&1&1&1&0&1&1&0&1\\ 0&1&0&1&0&0&0&1&0&0&0&1&0&0&0&0&1&0&1&1&0&1&0\\ &&&&&\vdots&&&&&&&&&&&&\vdots&&&&&\\ 0&1&0&0&1&0&1&0&0&0&1&0&1&0&0&1&0&1&0&0&0&0&0\end{array}\right)$
+of size $12\times 23$, is constructed from the binary expansion of $\pi$.
+We consider this code as representative of random codes.
+One can check that it has $d=3$, $d^{\perp}=4$.
+Then Experimental fact 1 is verified, as the defect is positive for $4\leq r\leq 10$.
+Moreover at the extremities of this interval we have equality $\beta_{2,4}=A_{4}(C^{\perp})=4$ and $\beta_{9,10}=A_{3}(C)=1$, as predicted.
+And the defect is zero for $r<4$ and $r>10$ so Experimental fact 2 is verified also.
+Last, Figure 13 reproduces Figure 3, the binary Golay code.
+As $d,d^{\perp}\geq 7$, Experimental fact 1 is trivially verified.
+On the other hand, Experimental fact 2 is not verified, but this only illustrates the fact that the Golay code is certainly not representative of random codes.
+Indeed it has a lot of algebraic structure, which explains it admitting special syzygies.
+More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $56,16]_{2}-$codes.
+For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling).
+The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
+Here we have $r_{\operatorname{crit}}\approx 4.86$.
+Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
+Experiments with other parameters display similar results.
+$d=11$
+$d=12$
+$d=13$
+$d=14$
+$d^{\perp}\!=\!3$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.269&[1,3]\\ 4&23.821&[15,55]\\ 5&6.927&[5,21]\\ 6&1.341&[1,7]\\ 7&0.042&[0,1]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.245&[1,3]\\ 4&23.171&[15,52]\\ 5&1.948&[1,8]\\ 6&0.086&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.201&[1,3]\\ 4&21.975&[15,48]\\ 5&0.345&[0,5]\\ 6&0.006&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.164&[1,3]\\ 4&20.902&[14,47]\\ 5&0.067&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$d^{\perp}\!=\!4$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&6.178&[1,14]\\ 5&6.514&[5,20]\\ 6&1.263&[1,7]\\ 7&0.035&[0,1]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.963&[1,14]\\ 5&1.882&[1,8]\\ 6&0.090&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.525&[1,12]\\ 5&0.357&[0,5]\\ 6&0.010&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&4.885&[1,11]\\ 5&0.053&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$d^{\perp}\!=\!5$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&5.847&[5,15]\\ 6&1.153&[1,6]\\ 7&0.020&[0,1]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&1.485&[1,6]\\ 6&0.055&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.197&[0,2]\\ 6&0.002&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.033&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
+Figure 14: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $56,16]_{2}-$codes
+Consequently, we postulate the following relaxation of the minimal resolution conjecture:
 Heuristic 1
-Fix a field cardinality $q$, assume $n$ is not too close to $k$ in order to stay away from the counterexamples to the minimal resolution conjecture, and $n<\binom{k+1}{2}$ in order to ensure regularity $2$.
-Then for random $[n,k]_{q}$-codes, with high probability:
+Fix a field cardinality $q$, assume $n$ is not too close to $k$ in order to stay away from the counterexamples to the minimal resolution conjecture (e.g.
+$n>k+\sqrt{2k}+2$ should suffice), and $n<\binom{k+1}{2}$ in order to ensure regularity $2$.
+Then for random $n,k]_{q}-$codes, with high probability:
   1. 1.
-    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$
+    if $d>k+1-r_{\operatorname{crit}}$ we expect $\beta_{r-1,r}=0$ for $r>r_{\operatorname{crit}}$
   2. 2.
-    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$.
+    if $d^{\perp}>r_{\operatorname{crit}}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<r_{\operatorname{crit}}$
+where $r_{\operatorname{crit}}=\frac{k(k+1)}{n}$.
 Remark 3
@@
 Using syzygies of higher degree $r$ allows to make the term $\frac{k(k+1)}{r}$ smaller and reach a broader range of parameters.
-We compared our new distinguisher (in this basic version) with the one from [7], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$.
-Results can be found at the end of the Supplementary material.
+In the following paragraphs we compare our new distinguisher (in this basic version) with the one from [10], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$.
 In both cases, we improve the distinguishability threshold.
-Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (86) (with $r=r_{\max}^{*}=4$), i.e.
+Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (111) (with $r=r_{\max}=4$), i.e.
 the conditions in Heuristic 1 are always met.
@@
 asymptotically all parameters will become distinguishable.
+Distinguishing Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
+For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{97}$, while the distinguisher of [10] works down to $n_{\text{CMT}}=\mathbf{80}$.
+We experimentally find that dual Goppa codes $\mathsf{C}$} with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, matching the nonbinary case at the end of Remark 1 ($a=\lceil\log_{q}(t/(q-1))\rceil=1$, $r^{*}=(a+1)t-q^{a}=4$).
+Moreover, their Betti numbers are:
+$\beta_{1,2}^{*}=40,\qquad\beta_{2,3}^{*}=80,\qquad\beta_{3,4}^{*}=12.$ (112)
+We then observe that the $\beta_{2,3}$-distinguisher works down to $n_{\beta_{2,3}}=\mathbf{86}$, and the $\beta_{3,4}$-distinguisher works down to $n_{\beta_{3,4}}=\mathbf{68}$, both of which coincide with (111).
+More precisely, computing the Betti numbers of dual Goppa and random codes around these values of $n$ consistently yields:
+[…]
+We see that they stick to their “predicted” values: $\beta_{r-1,r}=\max(\beta_{r-1,r}^{*},\operatorname{ind}(\varphi_{r})^{+})$ for dual Goppa, and $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})^{+}$ for random codes.
+Distinguishing Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
+For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{62}$, while the distinguisher of [10] works down to $n_{\text{CMT}}=\mathbf{59}$.
+We experimentally find that dual Goppa codes $\mathsf{C}$} with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, matching the value given in Remark 1 ($a=\lceil\log_{2}(t))\rceil=2$, $r^{*}=(a+2)t-2^{a}=8$).
+Moreover, their top Betti numbers are:
+$\beta_{5,6}^{*}=1020,\qquad\beta_{6,7}^{*}=288,\qquad\beta_{7,8}^{*}=42.$ (113)
+From (111) we expect to distinguish at $\beta_{5,6}$ for $n\geq 57$, at $\beta_{6,7}$ for $n\geq 49$, and at $\beta_{7,8}$ for $n\geq 43$.
+And indeed at $n_{\beta_{5,6}}=\mathbf{57}$ we consistently find $\beta_{5,6}\geq 1020$ for dual Goppa codes, while $\beta_{5,6}<500$ for random codes with quite high probability.
+For smaller $n$ we have to pass to $\beta_{6,7}$.
+The $\beta_{6,7}$-distinguisher works well for $n=56$, but the quality gradually falls (distinguishing errors occur more frequently) as $n$ becomes smaller.
+It is difficult to point a precise threshold where the distinguisher ceases to work.
+Arguably we still have a positive advantage at $n_{\beta_{6,7}}=\mathbf{50}$, but not anymore at $n=49$.
+We could then try with $\beta_{7,8}$, but this fails too.
+What happens?
+It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$.
+Indeed, in order to have $\beta_{r-1,r}=0$, we need r>k+1dr>k+1-d, where $d=\operatorname{d}_{\min}({\mathsf{C}})$.
+For $r=7$ and $k=mt=18$ this gives $d\geq 13$.
+Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of $\mathsf{C}$}.
+Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity.
+Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably.
+Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
+Remark 4
+The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher.
+Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$.
+Let us thus consider a distinguisher that decides that a given [n,n18][n,n-18]-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise.
+Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$.
+This means that this distinguisher has more than $99\%$ success rate!
+And it remains remarkably effective even for smaller $n$.
+For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.
+In these experiments, random codes were defined by taking a uniformly random $(n-18)\times n$ generator matrix, which might quite often produce degenerate codes as $n$ approaches $18$.
+Other probability distributions preventing this problem could be used, which would slightly alter the success rate, but not the qualitative behaviour of this distance-based distinguisher.
 Full version (with shortening)
 We can also use shortened dual codes as input in our distinguisher.
-In general, shortening $s$ times a $[n,k]$-code ${\mathsf{C}}$ produces a $[n_{s},k_{s}]$-code ${\mathsf{C}}_{s}$ with $n_{s}=n-s$, $k_{s}=k-s$.
+In general, shortening $s$ times a $[n,k]$-code $\mathsf{C}$} produces a $n_{s},k_{s}]-$code ${\mathsf{C}}_{s}$ with $n_{s}=n-s$, $k_{s}=k-s$.
 Having smaller parameters will decrease the complexity.
@@
 Without shortening, these codes are out of reach of our distinguisher: we would have to run Algorithm 1 up to $r=10$ and do linear algebra in dimension roughly $10^{20}$, which is infeasible.
-On the other hand, we can effectively distinguish after shortening $40$ times: we experimentally find that $40$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ have $\beta_{3,4}=30$, while $\beta_{3,4}=0$ for random codes with the same parameters.
-Computational cost (several days each, on a workstation with several hundreds of GB of RAM) prevented the author to run more than a few tests, but no deviation was observed.
+On the other hand, we can effectively distinguish after shortening $40$ times: using the code in [34] we experimentally find that $40$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ have $\beta_{3,4}=30$ (Figure 8), while $\beta_{3,4}=0$ for random codes with the same parameters.
+Computational cost (a couple of weeks each, on a workstation with several hundreds of GB of RAM) prevented the author to run more than a few tests, but no deviation was observed.
 In particular $\beta_{3,4}=0$ consistently in the random case confirms Heuristic 1 for these parameters.
@@
 for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of $\mathsf{C}$}.
-Set $k=mt$, and let ${\mathsf{C}}$ be either a random $[n,k]_{q}$-code, or an element of $\mathcal{T}_{q,m,n,t}$.
+Set $k=mt$, and let $\mathsf{C}$} be either a random $n,k]_{q}-$code, or an element of $\mathcal{T}_{q,m,n,t}$.
 Choose an integer $s$ such that
-$\frac{k_{s}(k_{s}+1)}{n_{s}}<r^{*}-s$ (88)
-where $n_{s}=n-s$ and $k_{s}=k-s$.
+$r_{\operatorname{crit}}({\mathsf{C}}_{s})=\frac{k_{s}(k_{s}+1)}{n_{s}}<r^{*}-s$ (115)
+where $n_{s}=n-s$ and $k_{s}=k-s$. Then set
+$r=r^{*}-s.$ (116)
   1. 1.
-    Let $d_{s}\approx\operatorname{d}_{GV}(q,n_{s},k_{s})$ be the typical minimum distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$.
-Then, under part 1.
-of Heuristic 1, computing
-    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s}),$ (89)
-    where ${\mathsf{C}}_{s}$ is any $s$-shortened subcode of ${\mathsf{C}}$, allows to distinguish whether ${\mathsf{C}}$ is a random code (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ w.h.p.) or an element of $\mathcal{T}_{q,m,n,t}$ (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$).
+    Let $d_{s}\approx\operatorname{d}_{GV}(q,n_{s},k_{s})$ be the typical minimum distance of random $n_{s},k_{s}]_{q}-$codes, and assume $d_{s}>k_{s}+1-r_{\operatorname{crit}}({\mathsf{C}}_{s})$.
+Then, under part 1 of Heuristic 1, computing
+    $\beta_{r-1,r}({\mathsf{C}}_{s})$ (117)
+    where ${\mathsf{C}}_{s}$ is any $s$-shortened subcode of $\mathsf{C}$}, allows to distinguish whether $\mathsf{C}$} is a random code (in which case $\beta_{r-1,r}({\mathsf{C}}_{s})=0$ w.h.p.) or an element of $\mathcal{T}_{q,m,n,t}$ (in which case $\beta_{r-1,r}({\mathsf{C}}_{s})>0$).
   2. 2.
-    Moreover, let $d_{s}^{\perp}\approx\operatorname{d}_{GV}(q,n_{s},n_{s}-k_{s})$ be the typical dual distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
-Then, under Heuristic 1, for large enough parameters the complexity of this distinguisher is dominated by the sum
-    $\kappa=\sum_{i\leq\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1}\max\left(k_{s}\operatorname{ind}(\varphi_{i-1}),\binom{k_{s}+1}{2}\operatorname{ind}(\varphi_{i-2})\right)^{\omega}$ (90)
+    Moreover, let $d_{s}^{\perp}\approx\operatorname{d}_{GV}(q,n_{s},n_{s}-k_{s})$ be the typical dual distance of random $n_{s},k_{s}]_{q}-$codes, and assume $d_{s}^{\perp}>r_{\operatorname{crit}}({\mathsf{C}}_{s})$.
+Then, under Heuristic 1, for large enough parameters, the complexity of this distinguisher when Betti numbers are computed with the iterative Algorithm 1 is dominated by the sum
+    $\kappa_{\operatorname{iter}}=\sum_{i\leq\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1}\max\left(k_{s}\operatorname{ind}(\varphi_{i-1}),\binom{k_{s}+1}{2}\operatorname{ind}(\varphi_{i-2})\right)^{\omega}$ (118)
     where $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$ and $\omega\approx 2.372$ is the exponent of linear algebra.
@@
 If $\mathsf{C}$} is random, so is ${\mathsf{C}}_{s}$.
-Then, as $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 1.
-of Heuristic 1 we indeed expect $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ for $r^{*}-s>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
-On the other hand, for ${\mathsf{C}}$ in $\mathcal{T}_{q,m,n,t}$, the lower bound $r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ precisely means $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$.
-Now using Algorithm 1 we obtain $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $2\leq i\leq r^{*}-s$.
+Then, under part 1 of Heuristic 1, condition (115) and $d_{s}>k_{s}+1-r_{\operatorname{crit}}({\mathsf{C}}_{s})$ indeed imply $\beta_{r-1,r}({\mathsf{C}}_{s})=0$ for $r=r^{*}-s$.
+On the other hand, for $\mathsf{C}$} in $\mathcal{T}_{q,m,n,t}$, hypothesis (114) precisely means $\beta_{r-1,r}({\mathsf{C}}_{s})>0$.
+Now using Algorithm 1 we obtain $\beta_{r-1,r}({\mathsf{C}}_{s})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $2\leq i\leq r$.
 The contributions for $i<4$ are negligible.
-For $i\geq 4$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}({\mathsf{C}}_{s})\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{s})$ and the contributions for $i>\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1$ are negligible because of part 1.
-Then, as $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 2.
-of Heuristic 1 we expect $\beta_{i-1,i}({\mathsf{C}}_{s})=\operatorname{ind}(\varphi_{i})$ for $2\leq i\leq\frac{k_{s}(k_{s}+1)}{n_{s}}$ for random ${\mathsf{C}}$; and we can assume this holds also for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
+For $i\geq 4$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}({\mathsf{C}}_{s})\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{s})$ and the contributions for $i>\lfloor r_{\operatorname{crit}}({\mathsf{C}}_{s})\rfloor+1$ are negligible because of part 1.
+Then, as $d_{s}^{\perp}>r_{\operatorname{crit}}({\mathsf{C}}_{s})$, under part 2 of Heuristic 1 we expect $\beta_{i-1,i}({\mathsf{C}}_{s})=\operatorname{ind}(\varphi_{i})$ for $2\leq i\leq r_{\operatorname{crit}}({\mathsf{C}}_{s})$ for random $\mathsf{C}$}; and we can assume this holds also for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, otherwise computing this dimension readily provides a distinguisher with lower complexity.
 From this we conclude.
-Example 2
-The following table shows how the distinguisher from Proposition 8 applies to Classic McEliece parameters.
-We take $r^{*}=\widehat{f}$ as given by Corollary 4, and then choose the maximum possible shortening order $s$ satisfying (88).
-$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr(n,m,t)&(3488,12,64)&(4608,13,96)&(6688,13,128)&(6960,13,119)&(8192,13,128)\\ \hline\cr r^{*}&427&683&939&867&939\\ \hline\cr s&377&568&816&769&848\\ \hline\cr r^{*}-s&50&115&123&98&91\\ \hline\cr[n_{s},k_{s}]&[3111,391]&[4040,680]&[5872,848]&[6191,778]&[7344,816]\\ \hline\cr\frac{k_{s}(k_{s}+1)}{{n_{s}}_{\phantom{0\!\!}}}&49.27&114.62&122.61&97.89&90.78\\ \hline\cr\operatorname{d}_{GV}&921&1069&1650&1828&2256\\ \hline\cr\operatorname{d}_{GV}^{\perp}&55&62&122&108&110\\ \hline\cr\kappa&{2^{528}}^{\phantom{i\!}}&(2^{1080})&(2^{1224})&2^{1030}&2^{997}\\ \hline\cr\end{array}$
-We observe that the condition $\operatorname{d}_{GV}(n_{s},k_{s})>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$ is always satisfied with a large margin, so we could choose a smaller $s$ and still be able to distinguish.
-Concerning the complexity estimate $\kappa$, it relies on the condition $\operatorname{d}_{GV}(n_{s},n_{s}-k_{s})>\frac{k_{s}(k_{s}+1)}{n_{s}}$.
-We see this condition is satisfied for the parameter sets $(3488,12,64)$, $(6960,13,119)$, $(8192,13,128)$, but for $(4608,13,96)$ and $(6688,13,128)$ it is not (then, it would still be possible to give a complexity estimate, using Lemma 5 instead of Heuristic 1).
-These complexities improve those from [7], although they remain practically unreacheable and well beyond security levels.
+Algorithm 1, that iteratively computes the linear strand of the minimal resolution, can be optimized using methods from sparse linear algebra and structured linear algebra, as is done in [1].
+This might procure practical improvements, both on memory usage and on time complexity, but the theoretical analysis of these improvements remains to be done.
+A tentative implementation of the method of [1] can be found in [27].
+It would be very nice if this project could be continued, and maybe eventually incorporated in some computer algebra systems.
+An alternative approach through Koszul cohomology and the (block) Wiedemann algorithm
+Koszul cohomology [18] allows to directly compute one syzygy space, with no need to iteratively compute all the intermediate steps in the resolution.
+Let $\mathsf{C}$} be a $[n,k]$-code, $S=\mathbb{F}[X_{1},\dots,X_{k}]$ the polynomial ring in $k$ indeterminates, and $\operatorname{ev}:S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$ the graded evaluation map associated to some basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of $\mathsf{C}$}.
+Proposition 14
+For any $r$, the syzygy space $M_{r-1,r}({\mathsf{C}})$ can be computed as the cohomology of the sequence
+$\bigwedge\nolimits^{r}S_{1}\overset{\partial}{\longrightarrow}\left(\bigwedge\nolimits^{r-1}S_{1}\right)\otimes{\mathsf{C}}\overset{\partial^{\prime}}{\longrightarrow}\left(\bigwedge\nolimits^{r-2}S_{1}\right)\otimes{\mathsf{C}}^{\langle 2\rangle}$ (119)
+i.e. as
+$M_{r-1,r}({\mathsf{C}})=\ker(\partial^{\prime})/\operatorname{im}(\partial)$ (120)
+where the linear maps $\partial,\partial^{\prime}$ are defined on the standard basis elements as
+$\begin{array}[]{cccc}\partial:&\bigwedge^{r}S_{1}&\longrightarrow&(\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}}\\ &X_{i_{1}}\wedge\cdots\wedge X_{i_{r}}&\mapsto&\sum_{j=1}^{r}(-1)^{j}(X_{i_{1}}\wedge\cdots\widehat{X}_{i_{j}}\cdots\wedge X_{i_{r}})\otimes\mathbf{c}_{i_{j}}\end{array}$ (121)
+and
+$\begin{array}[]{cccc}\partial^{\prime}:&(\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}}&\longrightarrow&(\bigwedge^{r-2}S_{1})\otimes{\mathsf{C}}^{\langle 2\rangle}\\ &(X_{i_{1}}\wedge\cdots\wedge X_{i_{r-1}})\otimes\mathbf{c}_{l}&\mapsto&\sum_{j=1}^{r-1}(-1)^{j}(X_{i_{1}}\wedge\cdots\widehat{X}_{i_{j}}\cdots\wedge X_{i_{r-1}})\otimes(\mathbf{c}_{i_{j}}\mathbf{c}_{l})\end{array}$ (122)
+where notation $\widehat{X}_{i_{j}}$ means that $X_{i_{j}}$ is omitted.
+Proof
+Apply [18, Th.(1.b.4)] with $B={\mathsf{C}}^{\langle\cdot\rangle}$, p=r1p=r-1, $q=1$ (or [13, Prop. 2.7] with $M={\mathsf{C}}^{\langle\cdot\rangle}$, i=r1i=r-1, $j=r$).
+By construction $\partial,\partial^{\prime}$ satisfy $\operatorname{im}(\partial)\subseteq\ker(\partial^{\prime})$, and elementary linear algebra allow to rephrase (120) as
+$M_{r-1,r}({\mathsf{C}})\simeq\ker(\partial^{\prime}|_{T})$ (123)
+where $T$ is any complementary subspace to $\operatorname{im}(\partial)$ in $(\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}}$.
+Let us denote the standard basis of $(\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}}$ by
+$\mathbf{e}_{i_{1},\dots,i_{r-1};l}=(X_{i_{1}}\wedge\cdots\wedge X_{i_{r-1}})\otimes\mathbf{c}_{l}$ (124)
+for $1\leq i_{1}<\cdots<i_{r-1}\leq k$ and $1\leq l\leq k$.
+Lemma 14
+The subfamily of the basis (124) made of those $\mathbf{e}_{i_{1},\dots,i_{r-1};l}$ with $l\leq i_{r-1}$ is a basis of a complementary subspace $T$ to $\operatorname{im}(\partial)$ in $(\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}}$.
+Proof
+First, a subfamily of a basis is linearly independent, hence a basis of its linear span $T$.
+Next we show $\operatorname{im}(\partial)\cap T=0$. An element of $\operatorname{im}(\partial)$ is of the form
+$\sum_{1\leq i_{1}<\cdots<i_{r}\leq k}\lambda_{i_{1},\dots,i_{r}}\sum_{1\leq j\leq r}(-1)^{j}\mathbf{e}_{i_{1},\dots\widehat{i_{j}}\dots,i_{r};i_{j}}.$ (125)
+If such an element is in $T$, then the coefficient of each $\mathbf{e}_{i_{1},\dots,i_{r-1};i_{r}}$ is 0 (because $i_{r}>i_{r-1}$).
+However the coefficient of $\mathbf{e}_{i_{1},\dots,i_{r-1};i_{r}}$ is $(-1)^{r}\lambda_{i_{1},\dots,i_{r}}$.
+Hence all $\lambda_{i_{1},\dots,i_{r}}=0$.
+To conclude, we show that $\operatorname{im}(\partial)+T$ contains all the $\mathbf{e}_{i_{1},\dots,i_{r-1};l}$.
+If $l\leq i_{r-1}$ we already have $\mathbf{e}_{i_{1},\dots,i_{r-1};l}\in T$.
+On the other hand, if $l>i_{r-1}$, then we can set $i_{r}=l$ and we find $\mathbf{e}_{i_{1},\dots,i_{r-1};l}=\partial(\mathbf{s})-\mathbf{t}$ where $\mathbf{s}=(-1)^{r}X_{i_{1}}\wedge\cdots\wedge X_{i_{r}}$ and where $\mathbf{t}=\sum_{j=1}^{r-1}(-1)^{r-j}\mathbf{e}_{i_{1},\dots\widehat{i_{j}}\dots,i_{r};i_{j}}\in T$ (because $i_{j}\leq i_{r}$ for $j\leq r-1$).
+Lemma 15
+We have
+$\dim(T)=(r-1)\binom{k+1}{r}=\frac{k(k+1)}{r}\binom{k-1}{r-2}$. (126)
+Proof
+We give two proofs.
+For the first proof, we count the number of indices $i_{1},\dots,i_{r-1};l$ in Lemma 14.
+Conditioning on $i_{r-1}=u$, the number of $i_{1},\dots,i_{r-2}$ with $1\leq i_{1}<\cdots<i_{r-2}<u$ is $\binom{u-1}{r-2}$ and the number of $l$ with $1\leq l\leq u$ is $u$.
+It follows $\dim(T)=\sum_{u\leq k}u\binom{u-1}{r-2}$, and an easy manipulation with binomal coefficients gives (126).
+For the second proof, we observe that $\partial$ is injective, so that $\dim(T)=\dim((\bigwedge^{r-1}S_{1})\otimes{\mathsf{C}})-\dim(\bigwedge^{r}S_{1})=k\binom{k}{r-1}-\binom{k}{r}$, and we conclude likewise.
+Remark 5
+The sequence (119) continues with
+$\left(\bigwedge\nolimits^{r-2}S_{1}\right)\otimes{\mathsf{C}}^{\langle 2\rangle}\overset{\partial^{\prime\prime}}{\longrightarrow}\left(\bigwedge\nolimits^{r-3}S_{1}\right)\otimes{\mathsf{C}}^{\langle 3\rangle}\longrightarrow\cdots\longrightarrow C^{\langle r\rangle}\longrightarrow 0$ (127)
+and we have $\operatorname{im}(\partial^{\prime})\subseteq U=\ker(\partial^{\prime\prime})$. Also identifying $T$ with $\operatorname{coker}(\partial)$, the map $\partial^{\prime}$ then induces a map
+$\overline{\partial^{\prime}}:T\longrightarrow U$ (128)
+with $\dim\ker(\overline{\partial^{\prime}})=\beta_{r-1,r}({\mathsf{C}})$ and $\dim\operatorname{coker}(\overline{\partial^{\prime}})=\beta_{r-2,r}({\mathsf{C}})$.
+Moreover, if $\mathsf{C}$} has regularity $2$, then:
+  • in (127) we can replace ${\mathsf{C}}^{\langle 2\rangle},{\mathsf{C}}^{\langle 3\rangle},\dots,{\mathsf{C}}^{\langle r\rangle}$ with $\mathbb{F}^{n}$
+  • this sequence (127) is acyclic.
+This allows to compute
+$\dim(U)=n\sum_{i=2}^{r}(-1)^{i}\binom{k}{r-i}=n\binom{k-1}{r-2}$. (129)
+We then find $\operatorname{ind}(\overline{\partial^{\prime}})=\dim(T)-\dim(U)=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ which, as expected, is the same as $\operatorname{ind}(\varphi_{r})$ given in Theorem 4.1.
+Remark 6
+On the other hand, if we compute the index of $\partial^{\prime}|_{T}$ without corestriction of the target space, we find, for $\mathsf{C}$} of regularity $2$,
+$\begin{split}\operatorname{ind}(\partial^{\prime}|_{T})&=\dim(T)-\dim\left(\left(\bigwedge\nolimits^{r-2}S_{1}\right)\otimes\mathbb{F}^{n}\right)\\ &=\frac{k(k+1)}{r}\binom{k-1}{r-2}-n\binom{k}{r-2}\;<\;\operatorname{ind}(\varphi_{r}).\end{split}$ (130)
+In most applications we will have $\operatorname{ind}(\varphi_{r})<0$, hence also $\operatorname{ind}(\partial^{\prime}|_{T})<0$.
+In the simplified version of our distinguisher, we do not need to actually compute the syzygy space $M_{r-1,r}=\ker(\partial^{\prime}|_{T})$, nor its dimension $\beta_{r-1,r}$: we only want to decide whether it is 0 or not.
+Wiedemann’s algorithm [40], in its basic version, probabilistically finds a preimage of a nonsingular linear map $A$, with only black-box queries to $A$.
+This means that the overall cost of the algorithm does not involve the cost of general linear algebra, but only the cost $c(A)$ of evaluation of $A$.
+Several reductions provide variants of the algorithm that find a nonzero element in the kernel of a singular $A$, at essentially the same cost.
+This allows us to propose a “Koszul-Wiedemann” (KW) alternative implementation of the distinguisher.
+We keep the setting of part 1 of Proposition 13, in particular the integer $s$ should satisfy (115), and the typical minimum distance $d_{s}$ should satisfy $d_{s}>k_{s}+1-r_{\operatorname{crit}}({\mathsf{C}}_{s})$ where $r_{\operatorname{crit}}({\mathsf{C}}_{s})=\frac{k_{s}(k_{s}+1)}{n_{s}}$.
+Then we apply Wiedemann’s algorithm to find a nonzero element in the kernel of the linear map $\partial^{\prime}|_{T}$ associated with ${\mathsf{C}}_{s}$ and $r=r^{*}-s$:
+  1. (a)
+    if Wiedemann’s algorithm succeeds, we declare $\mathsf{C}$} is in $\mathcal{T}_{q,m,n,t}$
+  2. (b)
+    otherwise, we declare $\mathsf{C}$} is probably random.
+Proposition 15
+Under part 1 of Heuristic 1, this procedure is correct in the following sense:
+  • in case (a), the code $\mathsf{C}$} indeed is in $\mathcal{T}_{q,m,n,t}$ with probability very close to 11
+  • in case (b), the code $\mathsf{C}$} is random with nonnegligible probability (and we can then iterate the procedure in order to make this probability become very close to $1$).
+Each application of this procedure has complexity
+$\kappa_{KW}=O\left(n_{s}^{2}\binom{k_{s}}{r-2}\left((r-1)^{2}\binom{k_{s}+1}{r}+\binom{k_{s}}{r-2}\right)\right)$. (131)
+Proof
+Correctness follows from the discussion just above, and part 1 of Proposition 13.
+Now consider the map $\partial^{\prime}|_{T}$ associated with ${\mathsf{C}}_{s}$ and $r=r^{*}-s$:
+$\partial^{\prime}|_{T}:T\longrightarrow(\bigwedge\nolimits^{r-2}S_{1})\otimes\mathbb{F}^{n_{s}}$ (132)
+where $S$ now is the polynomial ring in $k_{s}$ indeterminates. By (126) the source space of this map has dimension
+$M=\dim(T)=(r-1)\binom{k_{s}+1}{r}$ (133)
+while on the other end the target space has dimension
+$N=n_{s}\binom{k_{s}}{r-2}$. (134)
+Moreover, by formula (122), this map sends each basis vector of $T$ to a vector of weight at most $(r-1)n_{s}$ in the standard basis of the target space.
+Thus the cost of evaluation of $\partial^{\prime}|_{T}$ satisfies
+$c(\partial^{\prime}|_{T})\leq M(r-1)n_{s}.$ (135)
+Now the estimate for the complexity of Wiedemann’s algorithm is
+$O(N(c(\partial^{\prime}|_{T})+N))$ (136)
+and we conclude.
+It is also possible to use the block version of Wiedemann’s algorithm [8].
+This definitely improves the complexity, but does not qualitatively change the theoretical analysis (131) as this improvement can be incorporated in the constant in the big O notation.
+Two tentative implementations of the KW version of the distinguisher can be found in [28], one with the original Wiedemann algorithm, one with the block Wiedemann algorithm.
+It was especially tempting to leverage the very optimized implementation of block Wiedemann that can be extracted from [39].
+However several technical difficulties appear, such as the fact that this implementation of block Wiedemann was not designed to make black-box queries to the linear map, or also the fact that it uses heuristics adapted to the specifics of the linear systems appearing in the number field sieve, while these heuristics can lead to an unexpected behaviour in our use case.
+Application to the Classic McEliece parameters
+The following two tables show how the distinguishers from Propositions 13 and 15 apply to Classic McEliece parameters.
+In Figure 15 we use for $r^{*}$ the lower bound $\widehat{f}$ proven in Corollary 5, and then choose the maximum possible shortening order $s$ satisfying (115).
+We observe that the KW approach provides a noticeable complexity improvement over the iterative approach.
+In Figure 16 we do the same thing but using the tight lower bound on $r^{*}$ stated in Remark 1. This provides a further improvement, for both approaches.
+The reader should keep in mind that the expression used for $\kappa_{\operatorname{iter}}$ assumes that linear algebra in dimension $n$ can be done in complexity $n^{\omega}$, which only holds asymptotically.
+Likewise the value used for $\kappa_{KW}$ in the tables is the inside of the big O in formula (131) so that some constant factor is discarded.
+More fundamentally, we avoided defining a precise model of computation.
+As a consequence, the exact numbers given here should not be taken too literally.
+But still, they give a reasonably faithful idea of what is happening.
+In particular we see that these complexity estimates significantly improve those from [10], although they remain practically unreacheable and well beyond security levels.
+[…]
+Figure 15: complexity estimates using the lower bound $\widehat{f}$ for $r^{*}$
+[…]
+Figure 16: complexity estimates using the tight $r^{*}$
+In all instances we observe that the condition $\operatorname{d}_{GV}(n_{s},k_{s})>k_{s}+1-r_{\operatorname{crit}}({\mathsf{C}}_{s})$ is satisfied with a large margin.
+This means that if the reader is reluctant to use part 1 of Heuristic 1 with $r$ so close to $r_{\operatorname{crit}}$, then it would actually be possible to choose a smaller $s$ and still distinguish (at the cost of a slightly worse complexity).
+Concerning part 2 of Heuristic 1, we see that the condition $\operatorname{d}_{GV}(n_{s},n_{s}-k_{s})>r_{\operatorname{crit}}({\mathsf{C}}_{s})$ is satisfied in both tables for the parameter sets $(3488,12,64)$, $(6960,13,119)$ and $(8192,13,128)$.
+For $(6688,13,128)$ the condition holds when we use the tight estimate for $r^{*}$.
+For $(4608,13,96)$ it fails, so there might be some doubt on the validity of $\kappa_{\operatorname{iter}}$ in this particular case.
+On the other end we recall that Proposition 15 does not rely on this part of the heuristic, hence $\kappa_{KW}$ is unaffected.
 Asymptotics
@@
   • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$.
-In the case of binary Goppa codes, the $\widehat{f}$ from Corollary 4 improves the $f$ from Theorem 3.1.
-However one could show that both have the same asymptotics.
+In the case of binary Goppa codes, the $\widehat{f}$ from Corollary 5 and the $r^{*}$ from Remark 1 improve the $f$ from Theorem 3.1.
+However one could show that they all have the same asymptotics.
 So, in order to distinguish these codes from random codes, we will not care about the extra structure: treating Goppa codes as alternant codes will suffice.
@@
 Asymptotically, $q$-ary alternant (including Goppa) codes of dual rate $R$ can be distinguished from random codes, with complexity at most
-$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}$ (92)
-where $\omega\approx 2.372$ is the exponent of linear algebra.
+$\kappa=q^{\left(\frac{2R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}.$ (138)
 Proof
-We use Proposition 8 with $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, and $r^{*}=f$ as provided by Theorem 3.1.
+We use part 1 of Proposition 13 with $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, and $r^{*}=f$ as provided by Theorem 3.1.
 The analysis will be easier if rephrased in terms of $k_{r^{*}}=k-r^{*}$ and $r=r^{*}-s$.
@@
 Moreover the shortened code ${\mathsf{C}}_{s}$ has rate $\frac{k_{s}}{n_{s}}=\frac{k_{r^{*}}+r}{n-k+k_{r^{*}}+r}=o(1)$, so by Remark 3 both conditions in Heuristic 1 are satisfied.
-Thus Proposition 8 applies.
-[…]
-In [5, §3.4], the security level $2^{b}$ — or less formally, the number $b$ of targeted “bits of security” — of the Classic McEliece system, essentially defined as the complexity of the currently best attack (namely, by information set decoding), is shown asymptotically to be linear in $t\propto n/\log(n)$.
+Thus the hypotheses of Proposition 13 are satisfied, and both the iterative and the KW version of the distinguisher work with these parameters.
+It turns out the KW approach gives a slightly better asymptotic complexity.
+Now in (131) we see that, up to factors of lesser order, this complexity is essentially $\binom{k_{s}}{r}^{2}=\binom{k_{r^{*}}+r}{r}^{2}$.
+Using $r\ll k_{r^{*}}$ and Stirling’s formula we find $\log_{q}\binom{k_{r^{*}}+r}{r}\sim\log_{q}\binom{k_{r^{*}}}{r}\sim r\log_{q}\left(\frac{k_{r^{*}}}{r}\right)\sim r\log_{q}\log_{q}(n)$, and we conclude.
+In [6, §3.4], the security level $2^{b}$ — or less formally, the number $b$ of targeted “bits of security” — of the Classic McEliece system, essentially defined as the complexity of the currently best attack (namely, by information set decoding), is shown asymptotically to be linear in $t\propto n/\log(n)$.
 Then in (138) we have
-$\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$
+$\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$ (143)
 which means that the complexity of our distinguisher is subexponential in $b$ (although, admittedly, only very slightly so).
-6 Conclusion and open problems
-We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms; more precisely, subexponential in the error-correcting capability of the code.
-However this is only an asymptotic result: for concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the best attacks by a non-negligible factor.
-Problem 1
-Improve the implementation of this distinguisher.333addendum: actually, the improvements discussed below can already be found in works of Kreuzer-Lusteck and Albano-La Scala
-For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be (i) somewhat sparse and (ii) already partially reduced.
-This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster.
-Alternatively, instead of iteratively computing a whole range of the linear strand, one could directly compute a targeted Betti number, e.g.
-using Koszul cohomology (the author would like to thank W.
-Castryck for this suggestion).
-It is then likely that this Koszul cohomology computation could also benefit from similar optimizations (i) and (ii) exploiting sparsity and structure.
-Problem 2
-Can our distinguisher, and in particular the Betti number computation, benefit from a quantum speedup?
-Problem 3
-Our lower bounds on the Betti numbers, and in particular the $r_{\max}$, of dual alternant or Goppa codes, are not tight in general.
-Can they be improved?
-Better, can one give a complete, explicit description of the minimal resolution of these codes?
-(See at the beginning of the Supplementary material for elements in this direction.)
-We initiated the study of higher syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
-Problem 4
-Pursue this study in a systematic way.
-This should cover basic operations on codes (some of which have a nice geometric interpretation, some of which less so), and also crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
-Problem 5
-Can higher syzygies serve in decoding algorithms?
-Problem 6
-Betti numbers actually are invariants of the monomial equivalence class of a code (Lemma 3). Can they find other applications, for instance, to problems of code equivalence?
-Problem 7
-Can our distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack?
-For instance, one would like to combine our approach with:
-  • filtration arguments
-  • extraction of short relations, e.g. via MinRank techniques.
-(Observe that some of the explicit syzygies described in Proposition 3 are short, while Proposition 4 shows that short relations pass to shortened subcodes.)
-It is very common that new mathematical tools (Euclidean lattices, elliptic curves, pairings, isogenies…) are introduced in the cryptographic realm first for cryptanalytic purposes, i.e.
-to break systems.
-Once digested by the community, they are then used in a more constructive way, to build new cryptosystems.
-Problem 8
-Can one devise syzygy-based cryptography?
+As already mentioned, the existence of our subexponential distinguisher does not logically result in a downgrade of the security of the McEliece cryptosystem and its variants.
+All we can conclude is that the assumption that the original McEliece cryptosystem is OW-CPA is an ad hoc assumption, certainly backed by decades of practical experience, but that cannot easily be reduced to more fundamental theoretical justifications.
+It also leaves open the question of whether some of the ideas introduced here, possibly combined with other arguments, could lead to a subexponential key recovery algorithm.
 {credits}
@@
 The author is part of the ANR-21-CE39-0009 BARRACUDA and ANR-21-CE39-0008 KLEPTOMANIAC research projects. Some of the computations were done on the lame servers in Télécom Paris.
-Supplementary material
-More on syzygies of shortened dual alternant and Goppa codes
-It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$ in Theorem 3.1, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$.
-Let us first consider the simpler case $s=0$.
-The Eagon-Northcott complex of each such conjugate gives a subcomplex of the minimal resolution of ${\mathsf{C}}$, however one should take care of the fact that in general these subcomplexes are not in direct sum.
-Proposition 3 gives a basis $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ for each syzygy space $M_{r-1,r}(\mathbf{\Phi}^{(u)})$.
-The indices $i_{1},\dots,i_{r}$ correspond to columns of $\mathbf{\Phi}^{(u)}=\left(\begin{array}[]{c|c|c|c}\mathbf{B}^{(u+e)}_{0}&\mathbf{B}^{(u+e-1)}_{1}&\cdots&\mathbf{B}^{(u)}_{e}\end{array}\right)$.
-It is then easily seen:
-Proposition 9
-The $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ with column $i_{1}$ not in $\mathbf{B}^{(e+u)}_{0}$ also belong to $M_{r-1,r}(\mathbf{\Phi}^{(u-1)})$.
-Conversely, for $e<m/2$, those with $i_{1}$ in $\mathbf{B}^{(e+u)}_{0}$ are linearly independent, hence form a basis of the subspace $V_{r}=M_{r-1,r}(\mathbf{\Phi}^{(0)})+\cdots+M_{r-1,r}(\mathbf{\Phi}^{(r)})$ of $M_{r-1,r}({\mathsf{C}})$.
-This gives the improved lower bound:
-Corollary 6
-Set $e=\lfloor\log_{q}(t-1)\rfloor$ and $f=(e+1)t-\frac{q^{e+1}-1}{q-1}$. Then for any $r\geq 2$ we have
-$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{r})=m(r-1)\left(\binom{f}{r}-\binom{f-(t-1)}{r}\right).$ (97)
-For $r=2$ this specializes to
-$\beta_{1,2}({\mathsf{C}})\;\geq\;\dim(V_{2})=\frac{m(t-1)}{2}\left((2e+1)t-2\frac{q^{e+1}-1}{q-1}\right),$ (98)
-a result already proved in [12] by (essentially) the same method. It is observed moreover than in many cases the lower bound (98) is an equality, so we actually have
-$I_{2}({\mathsf{C}})=V_{2}.$ (99)
-Likewise for $r>f-(t-1)$ this gives
-$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{f})=m(r-1)\binom{f}{r}$ (100)
-and we will see examples where this is an equality.
-However for arbitrary $r$, the improved lower bound (97) still isn’t tight, because the inclusion $V_{r}\subseteq M_{r-1,r}({\mathsf{C}})$ is strict in general.
-For instance, one can show that $M_{2,3}({\mathsf{C}})$ contains syzygies of the form
-$\begin{split}X_{a}(&X_{b+1}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}+1}^{(u)}X_{c}^{(u+v)})-X_{a+q^{u}}(X_{b}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}}^{(u)}X_{c}^{(u+v)})+\\ &+X_{c}^{(u+v)}(X_{a}X_{b+q^{v}+1}^{(u)}-X_{a+q^{u}}X_{b+q^{v}}^{(u)})-X_{c+1}^{(u+v)}(X_{a}X_{b+1}^{(u)}-X_{a+q^{u}}X_{b}^{(u)})=0\end{split}$ (101)
-that do not belong to $V_{3}$ in general.
-In principle, under (99) it should be possible to use the explicit description of $V_{2}$ together with Proposition 2 to compute (the linear strand of) the minimal resolution of ${\mathsf{C}}$.
-However the combinatorics appears quite complicated.
-Now for general $s$, (68) and (73) can be extended to:
-Experimental fact 3
-Let ${\mathsf{C}}$ be a dual alternant code. Then for all $s$ we have
-$r_{\max}({\mathsf{C}}_{s})\geq r_{\max}({\mathsf{C}})-s.$ (102)
-It would be tempting to conjecture that (102) holds for all codes, but it turns out that one can find counterexamples.
-However, these counterexamples are quite rare.
-So maybe an interesting problem instead should be to give criteria for (102) to hold.
-Observe that shortening is dual to puncturing.
-As the behaviour of minimal resolutions of codes under puncturing is quite well understood (Corollary 3), maybe the study of shortening should go hand-in-hand with that of code duality.
-Experiments suggest a loose link between $\beta_{r-2,r}({\mathsf{C}}^{\perp})$ and $\beta_{k-r,k+1-r}(C)$, at least for some regimes of parameters.
-Back to dual alternant or Goppa codes, we actually observe a strong regular pattern (for Goppa codes the author only tested the irreducible case, but the result is likely to generalize):
-Experimental fact 4
-Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$ or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes, namely, either dual alternant codes, or dual Goppa codes with irreducible Goppa polynomial.
-Let $q$ be a field cardinality, and $t\geq 3$ an integer.
-  1. 1.
-    For all $m$ large enough,
-    $r_{\max}({\mathsf{C}})=r^{*}_{\mathcal{T},q,t}$ (103)
-    is the same for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,q^{m},t}$, i.e. it generically does not depend on $m$ nor on the choice of ${\mathsf{C}}$, but only on $\mathcal{T},q,t$.
-Now $\mathcal{T},q,t$ being fixed, we set $r^{*}=r^{*}_{\mathcal{T},q,t}$. Also, given $m,t$, we set $k=mt$.
-  1. 2.
-    For all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, and for all $s\leq r^{*}-2$, if $r^{*}-s>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
-    $r_{\max}({\mathsf{C}}_{s})=r^{*}-s.$ (104)
-  2. 3.
-    For $20\leq i\leq r^{*}-$ 2 there are functions $b_{i}(r)$ (actually depending on $\mathcal{T},q,t$, so $b_{i}(r)=b_{i,\mathcal{T},q,t}(r)$) such that, for all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, for all $s\leq r^{*}-2$, and for all $r$ in the interval $2\leq r\leq r^{*}-s$, if $r>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
-    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,b_{r^{*}-s-r}(r).$ (105)
-  3. 4.
-    We have $b_{0}(r)=r-1$ independently of $\mathcal{T},q,t$, so (105) reduces to
-    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=m\,(r^{*}-s-1).$ (106)
-    More generally for $i\geq 1$ we have $b_{i}(r)=(r-1)\binom{r+i}{i}$ for a certain number of (but now, not for all) values of $\mathcal{T},q,t$, so then (105) reduces to
-    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,(r-1)\binom{r^{*}-s}{r}.$ (107)
-We observe that (68) gives a lower bound on $r^{*}$.
-For alternant codes, experimentally, this lower bound is an equality.
-However for Goppa codes this lower bound seems to be always a strict inequality.
-And even for binary Goppa codes, the improved lower bound (73) still is a strict inequality.
-$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&251&1400&3230&2480&\mathbf{1400}&\mathbf{480}&\mathbf{70}\\ 1&202&880&1170&\mathbf{840}&\mathbf{350}&\mathbf{60}&-\\ 2&154&440&\mathbf{450}&\mathbf{240}&\mathbf{50}&-&-\\ 3&107&\mathbf{200}&\mathbf{150}&\mathbf{40}&-&-&-\\ 4&66&\mathbf{80}&\mathbf{30}&-&-&-&-\\ 5&31&\mathbf{20}&-&-&-&-&-\\ 6&\mathbf{10}&-&-&-&-&-&-\\ 7&-&-&-&-&-&-&-\end{array}$
-Figure 4: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{2,10,5}\;$ ($r^{*}=8$)
-$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ \end{array}$
-Figure 5: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{3,5,6}\;$ ($r^{*}=8$)
-$\begin{array}[]{c|ccccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}&\beta_{8,9}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 8&280&3224&7464&4272&\mathbf{3360}&\mathbf{1728}&\mathbf{504}&\mathbf{64}\\ 9&249&2510&1800&\mathbf{1792}&\mathbf{1120}&\mathbf{384}&\mathbf{56}&-\\ 10&219&1856&\mathbf{840}&\mathbf{672}&\mathbf{280}&\mathbf{48}&-&-\\ 11&190&1260&\mathbf{360}&\mathbf{192}&\mathbf{40}&-&-&-\\ 12&162&720&\mathbf{120}&\mathbf{32}&-&-&-&-\\ 13&135&234&\mathbf{24}&-&-&-&-&-\\ 14&109&\mathbf{16}&-&-&-&-&-&-\end{array}$
-Figure 6: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,8,5}\;$ ($r^{*}=17$)
-$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr 0&180&1293&3090&3144&960&\mathbf{36}\\ 1&151&810&1350&732&\mathbf{30}&-\\ 2&123&450&450&\mathbf{24}&-&-\\ 3&96&210&24&-&-&-\\ 4&70&30&-&-&-&-\\ 5&45&-&-&-&-&-\\ 6&21&-&-&-&-&-\\ 7&-&-&-&-&-&-\end{array}$
-Figure 7: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{3,6,5}\;$ ($r^{*}=7$)
-$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 32&719&8474&\mathbf{450}&\mathbf{240}&\mathbf{50}&-\\ 33&662&6216&\mathbf{150}&\mathbf{40}&-&-\\ 34&606&4070&\mathbf{30}&-&-&-\\ 35&551&2034&-&-&-&-\\ \end{array}$
-Figure 8: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}\;$ ($r^{*}=38$)
-Figures 8-8 illustrate the last Experimental fact.
-Boldface values match (107).
-This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
-Computing the Betti diagram of some families of codes
-Regularity $2$ helps in computing the full Betti diagram of codes, as one row can be deduced from the other.
-In order to illustrate this, let us first recall that a $[n,k]$-code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
-Lemma 12 ([15, Th. 1], reformulated)
-Let ${\mathsf{C}}$ be a $[n,k]$-code with $n\leq 2k-1$. Assume ${\mathsf{C}}$ is MDS. Then for all $r\leq 2k+1-n$ we have $\beta_{r-2,r}({\mathsf{C}})=0$.
-Proposition 10
-Let ${\mathsf{C}}$ be a $[k+1,k]$ MDS code, for instance a parity code or a $[k+1,k]$ $\operatorname{{\mathsf{GRS}}}$ code.
-Then the nonzero Betti numbers of ${\mathsf{C}}$ are $\beta_{0,0}=\beta_{k-1,k+1}=1$, and
-$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (108)
-for $12\leq r\leq k-$ 1. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
-Proof
-The parameters imply that ${\mathsf{C}}$ has regularity $2$.
-By Lemma 12 we have $\beta_{r-2,r}=0$ for $r\leq k$.
-Then $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-(k+1)\right)\binom{k-1}{r-2}$ and we conclude with a straightforward calculation.
-Proposition 11
-Let ${\mathsf{C}}$ be a $[2k-1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are:
-  • $\beta_{0,0}=1$
-  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
-  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
-In particular the ideal $I({\mathsf{C}})$ is generated by $\beta_{1,2}=\frac{(k-1)(k-2)}{2}$ quadratic forms and $\beta_{1,3}=k-1$ cubic forms.
-Proof
-Using the basis $\mathbf{y},\mathbf{y}\mathbf{x},\dots,\mathbf{y}\mathbf{x}^{k-1}$ of ${\mathsf{C}}$, we see $I_{2}({\mathsf{C}})$ contains the $2\times 2$ minors of $\mathbf{\Phi}=\left(\begin{array}[]{cccc}X_{0}&X_{1}&\dots&X_{k-2}\\ X_{1}&X_{2}&\dots&X_{k-1}\end{array}\right)$, i.e.
-$I_{2}({\mathsf{C}})\supseteq\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$ of dimension $\binom{k-1}{2}$.
-Now set $n=2k-1$.
-Because $n\geq 2k-1$, we have $\dim{\mathsf{C}}^{\langle 2\rangle}=2k-1$, hence $\dim I_{2}({\mathsf{C}})=\binom{k-1}{2}$, so $I_{2}({\mathsf{C}})=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi})$.
-Corollary 1 and the Eagon-Northcott complex of $\mathbf{\Phi}$ then give $\beta_{r-1,r}=(r-1)\binom{k-1}{r}$.
-Because $n\leq 2k-1$, we have ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$.
-Then $\beta_{r-2,r}=\beta_{r-1,r}-\operatorname{ind}(\varphi_{r})=(r-1)\binom{k-1}{r}-\left(\frac{k(k+1)}{r}-(2k-1)\right)\binom{k-1}{r-2}$, and a straightforward calculation allows to conclude.
-Figures 10 and 10 illustrate Propositions 10 and 11.
-$\begin{array}[]{c|cccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&27&105&189&189&105&27&-\\ 2&-&-&-&-&-&-&-&1\end{array}$
-Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
-$\begin{array}[]{c|ccccccccc}&0&1&2&3&4&5&6&7\\ \hline\cr 0&1&-&-&-&-&-&-&-\\ 1&-&21&70&105&84&35&6&-\\ 2&-&7&42&105&140&105&42&7\end{array}$
-Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
-Exercise 1
-Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for $k+1<n<2k-1$.
-Experimental data on defects for random codes
-Figures 13-13 compare Betti diagrams for various types of codes of length $n=23$ and dimension $k=12$. Observe that these parameters give $\frac{k(k+1)}{n}\approx 6.78$.
-First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e.
-with all defects equal to zero.
-It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram.
-Such codes are MDS, i.e.
-they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified.
-Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g.
-$q=5$ suffices (the author did not search for optimality).
-These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified.
-Moreover, the converse Experimental fact 2 is also verified as all defects are zero.
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&880&1353&990&-&-&-&-&-&-\\ 2&-&-&-&-&-&330&1617&1870&1221&485&110&11\\ \end{array}$
-Figure 11: an idealized $[23,12]$-code according to the minimal resolution conjecture
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1397&1224&490&121&18&1&-&-\\ 2&-&-&4&44&234&820&1738&1888&1222&485&110&11\\ \end{array}$
-Figure 12: a (pseudo)random $[23,12]_{2}$-code ($d=3$, $d^{\perp}=4$)
-$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
-Figure 13: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
-Second, Figure 13 presents the Betti diagram of the binary code whose generator matrix
-$G=\left(\begin{array}[]{ccccccccccccccccccccccc}1&1&0&0&1&0&0&1&0&0&0&0&1&1&1&1&1&1&0&1&1&0&1\\ 0&1&0&1&0&0&0&1&0&0&0&1&0&0&0&0&1&0&1&1&0&1&0\\ &&&&&\vdots&&&&&&&&&&&&\vdots&&&&&\\ 0&1&0&0&1&0&1&0&0&0&1&0&1&0&0&1&0&1&0&0&0&0&0\end{array}\right)$
-of size $12\times 23$, is constructed from the binary expansion of $\pi$.
-We consider this code as representative of random codes.
-One can check that it has $d=3$, $d^{\perp}=4$.
-Then Experimental fact 1 is verified, as the defect is positive for $4\leq r\leq 10$.
-Moreover at the extremities of this interval we have equality $\beta_{2,4}=A_{4}(C^{\perp})=4$ and $\beta_{9,10}=A_{3}(C)=1$, as predicted.
-And the defect is zero for $r<4$ and $r>10$ so Experimental fact 2 is verified also.
-Last, Figure 13 reproduces Figure 3, the binary Golay code.
-As $d,d^{\perp}\geq 7$, Experimental fact 1 is trivially verified.
-On the other hand, Experimental fact 2 is not verified, but this only illustrates the fact that the Golay code is certainly not representative of random codes.
-Indeed it has a lot of algebraic structure, which explains it admitting special syzygies.
-More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes.
-For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling).
-The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
-Here we have $\frac{k(k+1)}{n}\approx 4.86$.
-Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
-$d=11$
-$d=12$
-$d=13$
-$d=14$
-$d^{\perp}\!=\!3$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.269&[1,3]\\ 4&23.821&[15,55]\\ 5&6.927&[5,21]\\ 6&1.341&[1,7]\\ 7&0.042&[0,1]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.245&[1,3]\\ 4&23.171&[15,52]\\ 5&1.948&[1,8]\\ 6&0.086&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.201&[1,3]\\ 4&21.975&[15,48]\\ 5&0.345&[0,5]\\ 6&0.006&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&1.164&[1,3]\\ 4&20.902&[14,47]\\ 5&0.067&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$d^{\perp}\!=\!4$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&6.178&[1,14]\\ 5&6.514&[5,20]\\ 6&1.263&[1,7]\\ 7&0.035&[0,1]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.963&[1,14]\\ 5&1.882&[1,8]\\ 6&0.090&[0,1]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&5.525&[1,12]\\ 5&0.357&[0,5]\\ 6&0.010&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&4.885&[1,11]\\ 5&0.053&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$d^{\perp}\!=\!5$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&5.847&[5,15]\\ 6&1.153&[1,6]\\ 7&0.020&[0,1]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&1.485&[1,6]\\ 6&0.055&[0,1]\\ 7&0.001&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.197&[0,2]\\ 6&0.002&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-$\begin{array}[]{c|c|c}r&\;\,\textrm{mean}&\;\;99\%\\ \hline\cr 2&0.000&[0,0]\\ 3&0.000&[0,0]\\ 4&0.000&[0,0]\\ 5&0.033&[0,1]\\ 6&0.000&[0,0]\\ 7&0.000&[0,0]\\ 8&0.000&[0,0]\end{array}$
-Figure 14: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes
-Distinguishing Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
-For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{97}$, while the distinguisher of [7] works down to $n_{\text{CMT}}=\mathbf{80}$.
-We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, which is better than the lower bound in Theorem 3.1 ($e=0$, $r_{\max}\geq t-1=3$).
-Moreover, their Betti numbers are:
-$\beta_{1,2}^{*}=40,\qquad\beta_{2,3}^{*}=80,\qquad\beta_{3,4}^{*}=12.$ (109)
-We then observe that the $\beta_{2,3}$-distinguisher works down to $n_{\beta_{2,3}}=\mathbf{86}$, and the $\beta_{3,4}$-distinguisher works down to $n_{\beta_{3,4}}=\mathbf{68}$, both of which coincide with (86).
-More precisely, computing the Betti numbers of dual Goppa and random codes around these values of $n$ consistently yields:
-$\begin{array}[]{|c|ccccccccccccc|}\hline\cr n&\cdots&88&87&\mathbf{86}&85&84&\cdots&70&69&\mathbf{68}&67&66&\cdots\\ \hline\cr\beta_{2,3}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&80&80&\mathbf{80}&85&100&&310&325&340&355&370&\\ \beta_{2,3}^{\text{random}}&&40&55&\mathbf{70}&85&100&&310&325&340&355&370&\\ \hline\cr\beta_{3,4}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&12&12&12&12&12&&12&12&\mathbf{12}&105&210&\\ \beta_{3,4}^{\text{random}}&&0&0&0&0&0&&0&0&\mathbf{0}&105&210&\\ \hline\cr\end{array}$
-We see that they stick to their “predicted” values: $\beta_{r-1,r}=\max(\beta_{r-1,r}^{*},\operatorname{ind}(\varphi_{r})^{+})$ for dual Goppa, and $\beta_{r-1,r}=\operatorname{ind}(\varphi_{r})^{+}$ for random codes.
-Distinguishing Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
-For these parameters, the square code distinguisher works down to $n_{\text{square}}=\mathbf{62}$, while the distinguisher of [7] works down to $n_{\text{CMT}}=\mathbf{59}$.
-We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, which is better than the lower bound in Corollary 4 ($\widehat{e}=1$, $r_{\max}\geq 7$).
-Moreover, their top Betti numbers are:
-$\beta_{5,6}^{*}=1020,\qquad\beta_{6,7}^{*}=288,\qquad\beta_{7,8}^{*}=42.$ (110)
-From (86) we expect to distinguish at $\beta_{5,6}$ for $n\geq 57$, at $\beta_{6,7}$ for $n\geq 49$, and at $\beta_{7,8}$ for $n\geq 43$.
-And indeed at $n_{\beta_{5,6}}=\mathbf{57}$ we consistently find $\beta_{5,6}\geq 1020$ for dual Goppa codes, while $\beta_{5,6}<500$ for random codes with quite high probability.
-For smaller $n$ we have to pass to $\beta_{6,7}$.
-The $\beta_{6,7}$-distinguisher works well for $n=56$, but the quality gradually falls (distinguishing errors occur more frequently) as $n$ becomes smaller.
-It is difficult to point a precise threshold where the distinguisher ceases to work.
-Arguably we still have a positive advantage at $n_{\beta_{6,7}}=\mathbf{50}$, but not anymore at $n=49$.
-We could then try with $\beta_{7,8}$, but this fails too.
-What happens?
-It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$.
-Indeed, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$.
-For $r=7$ and $k=mt=18$ this gives $d\geq 13$.
-Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$.
-Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity.
-Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably.
-Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
-Remark 3
-The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher.
-Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$.
-Let us thus consider a distinguisher that decides that a given $[n,n-18]$-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise.
-Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$.
-This means that this distinguisher has more than $99\%$ success rate!
-And it remains remarkably effective even for smaller $n$.
-For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.
-In these experiments, random codes were defined by taking a uniformly random $(n-18)\times n$ generator matrix, which might quite often produce degenerate codes as $n$ approaches $18$.
-Other probability distributions preventing this problem could be used, which would slightly alter the success rate, but not the qualitative behaviour of this distance-based distinguisher.
+The use of Koszul cohomology in this updated version of the paper originates from a suggestion by Wouter Castryck.
+The author also would like to thank the Eurocrypt 2025 anonymous referees for their comments.